A Bibliography of Publications in Notices of the American Mathematical Society (2020–2029)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 1.02

Title word cross-reference

R [Roj25]. ∞ [Rie23]. K [Xu22]. G_{a} [Gup23]. N_{0} [HB22]. CAT(0) [AM20].

-actions [Gup23]. -category [Rie23]. -stability [Xu22].

1 [BCG22a]. 101st [ABB^{+}22]. 19 [WWL^{+}21]. 1940s [Mur21]. 1950s [Mur23].

2 [BPJ24, BCG22b]. 2017 [Rib21]. 2020 [Ano20a, Ano20k, Ano20b, Ano20c, Ano20l, Ano20d, Ano20m, Ano20e, Ano20g, Ano20f, Ano20j, Ano20h, Ano20i, DS21a, HLQ^{+}20, Ull22]. 2021 [Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21g, Ano21f, Ano21m, Ano21h, Ano21i, Ano21j, BNS23, DS22, HPLQ^{+}21a, Mas22, MWW^{+}23, Tuf22]. 2022 [Ano22b, Ano22c, Ano22a, Ano22d, DDV^{+}20, Keh22, Tie22]. 2023 [Hau24]. 2024 [vOL25]. 2025 [Ano25a]. 20th [Fer25]. 21st [Dev22].

3-dimensional [WY23]. 3-manifold [Pot21]. 3000 [Wes23]. 3D [LST20, Mer22d].

A.T.V. [McF22]. AAUP [Tie22]. Abel [Ano20m, Ano21m, Ano22c, Rib21, DS20b, DS21a, DS21b, DS22, DS23, DS24, DS25, Gam25]. Abelian [Ort20]. Ability [GM23a]. abstract [Gal23]. abstraction [Fer25]. academic [Hau24]. access [Har23]. Accessible [LV23]. Accessing [Wol24]. achievements [Her22a]. acrobatic [Seg21]. across [DNF24]. actions [Gup23, RR24]. acts [Poi22]. acute [Bae24b]. Ada [Ric24].

Aderemi [Hou23b]. ADJOINT [BG21a]. administration [HW25a]. administrators [BFH^{+}21]. adventures [Gam20, MMP20]. adversarial [TJ23]. advisor [DHD^{+}22]. advocate [And24, HWW21]. affine [Gup23]. afraid [HKL25]. Africa [JL21, KN22]. African [BG21a, IJS22, Mas22, Sax20]. after [Ryb22a]. again [GPS22]. AI [ABH^{+}25, GGT23, HKL25, SM24]. aided [PST23]. airport [EMS22]. AJ [Her21a]. Alan [BF23, O’S25]. Albert [Ano21a]. Aleksandrovich [Ano21n]. Alele [HN22]. Alele-Williams [HN22]. Alexandrovna [DDV^{+}20]. algebra [BGT24, DM21, Kas23a, Mar23, MMP20, MWW^{+}23, Ram22, RR24, Ano21d]. Algebraic [BE22, Dic20, GK22, EK20, HMR^{+}25, Jen21, RR21, ST20, Xu22]. algorithmic [Gam25, Roj25]. algorithms [Ber22, Pom25]. Alice [GW20]. alignment [Tad21]. allocating [Fie20]. Almost [Cal23a, Tor20, Cal21b, Gal23, Gal23]. Almost-orthogonality [Tor20]. Alon [Ano22b]. along [EST24]. Alperin [FS20]. alternative [PST23]. Altona [Gar24b]. Amer. [Hau24]. America [Mar24, Ros20a, Ros21, dLLOW23]. American [ACFH24, LP24, Mas22, Mur21, Mur23, Sax20, Ken20b, Spr25]. Americans [IJS22]. Ampére [Awa21, Maw23]. Ampère-type [Maw23]. AMS [Ano24a, Har23, HW25b, Her21a, Her22b, Her22f, TM23]. analog [Gna21]. Analysis [BCH^{+}20, Agu24, Che23, DNF24, FLOS24, Gre23, GW20, MMP20, Men20, Sch21a, Tor20, Wes23]. analyst [Kuz20]. analytic [GK22]. analytical [TJ23]. Anatomy [ACMT25]. Ancient [DS20a, HP25]. André [Tsi24]. Andrei [FM20]. Andrew [ESSV23]. anemia [Yak20]. Ángel [DS24]. Angenent [Cal20a]. angle [HM22]. animal [Sch21a]. anisotropic [De 24]. announced [Keh22]. annually [RRS22]. anomalous [Sti23]. anonymous [ACFH24]. Aperiodic [Tre23]. app [OWD20]. apparatus [Seg21]. Applications [Cio22, PCR22, Che23, DR23, Gup23, HL24, Has23, Kom20, Rub20, Sch20, SSU23, Agu24]. Applied [BCC^{+}22, Ano22g, E21]. approach [JLL^{+}21, Ros21, Cen22]. approximation [CCL^{+}22, MM21, Tre25a]. Archimedean [Ben22]. arithmetic [Esn23, McF22, Pas23]. Arkansas [Kas23b]. Ars [GPS22]. art [Cas21, Far21, GPS22, Har24, Rat23]. article [ABC^{+}21]. Artist [Mer22a]. arts [Nor20]. arXiv [Her22b]. Askey [CIW22]. aspects [BK25]. Aspiring [Hau24]. Assessing [Lav20]. assessment [BWW24]. assistant [TM23]. assistants [PQY^{+}25]. assisted [Tao25]. Assoc [Ros20a, Ros21]. associate [Her22d, TM23]. Association [Hou23a]. astronomy [Ste23]. atmospheric [Cam23]. auctions [Lop21]. August [BNS23]. Australia [Kas21]. authors [Bar24a, Her22b]. automorphic [Gal23]. Avi [DS22]. Award [Ano20a, Ano20f]. Awarded [Ano21m, Ano20k, Ano20l, Ano20m, Ano21k, Ano22b, Ano22c, Ano22a, Ano22d, RRS22]. Awards [Keh22]. AWM [BLW21]. Axiomatics [Fer25]. axions [Kir21]. Aziz [Bee24].

B [ACF^{+}24, KMMM21]. Bôcher

[Ano20b, Ano23a]. Babylonian [Ste23]. back [GPS22]. backwards [Ran23]. bad [Shi25]. baguenaudier [Cal24a]. baguenaudiest [Cal24a]. Baguenaudy [Cal24a]. baking [Her21a]. Ball [Bas24a]. Banach [SSU23]. band [Tre25b]. band-limited [Tre25b]. Barbara [FMM24]. Baron [Gam20]. barrier [Peh22]. barriers [HN22]. baseball [OWD20]. based [BPJ24, BLSY22, KFN23, MPDL25, Ros21, TSAY22, Ude24]. Basic [Ald22]. Bass [BCL25]. Bateman [Gar24a]. Battista [Par21]. Bayesian [TSAY22]. be [Kir21, Rie23]. beauty [BCD23]. becoming [Ful24]. before [Kes22]. beginners [CWA25]. behavior [Sch21a, Tad21]. behind [Gam25, GLY20]. Beilinson [Ano20k]. belonging [Ols23a]. benefited [Her22b]. Benkart [AHRZ23, HR22]. Berenstein [ABC^{+}20]. Berkeley [Kes22]. Bernoulli [De 23]. Bernoulli-type [De 23]. Bernsen [Ste23]. Bertrand [Ano21b, vOL25]. Bessenrodt [Ols23b]. Beth [ADG^{+}24]. between [BBB^{+}24a, Dip20, FLOS24, RR24, Tay24, Xu22]. beyond [Bar24a, BLW21, Har21, Mas22, MPDL25, Lor20]. Białynicki [Fra24]. bias [LDAST24]. Bicycle [Ros20a]. bifurcation [BP20]. Bilinear [Nai21]. bind [DDV^{+}20]. biology [Cio22, CKR21, Dic20]. bioluminescence [Yan24]. biomedical [Kom20]. birational [HM21]. Birkhoff [Ano21e]. Birman [Her22f]. birthday [ABB^{+}22]. Birula [Fra24]. Bismut [Ano21k]. black [BLOS23, Csi24, Gio23, HL22, Jet22, HWW21]. Blaschke [Weg20]. blend [HK21]. Blockchain [Zha21]. blogs [Tho20]. blow [Mus22]. blow-up [Mus22]. Board [Ano20f]. Bob [BGH^{+}23, MWW^{+}23]. bombed [VZ23]. book [Agu24, Ald22, All20, ARS23, Bie20, Bim23, Bon23, Can23, Cen22, Clo22, Dav25, DeL22, Des21, Fer25, Fou20, Gal23, Gar24b, Hau24, Hay21, Kas21, Kaw23, Ken20b, Law20, Lon23, Lor20, Mak20, Mar24, Mar20, Nat23, Nor20, Ols23a, Oma21, Pak22, Par21, Poi22, Pom25, Rib21, Roh23, Ros20a, Ros21, Row22, San21, Shl21, Sil22, Tur22, Ull21, Ull22, Vel22, Wes23, Wil20, Ull21, Kaw23]. Books [Ald22, Har21]. Boris [Cro21a]. Borwein [Bai23]. Bose [Kir21]. Boston [EW25]. boundaries [FGKS25]. boundary [BdMG23, De 23, Mav23, PL25]. bounds [Sea23]. Bourgain [Ano21o, Gam20]. Boyd [Hou24, Tur24]. brace [Ven24]. Brack [Ste23]. Brack-Bernsen [Ste23]. Bratteli [BDE^{+}20]. bread [Her21a]. Breaking [Peh22, HN22]. Brennie [Ano20d, Ano21c, Ano23b]. Brian [DFP^{+}22]. Bridging [BGH20]. brief [Ass22]. Brilliance [Dum22]. bringing [Sch24]. Brownian [Bae21a, DKR22]. Bryna [Bee25a]. Bubbling [Mus22]. Buchsbaum [EW22]. budget [Ful24]. bundles [ST20, Sch20]. business [AHJ^{+}22].

C [KR24]. C. [ABB^{+}22, BHL^{+}21].

CAARMS [Mas22]. CAARMS25 [Mas22]. Caffarelli [DS24]. Calabi [Che24, Sun22]. calculus [CMN22, HB22, Spo22]. Calibrating [Lan22]. California [BCF^{+}24]. call [AAD^{+}25]. Calvin [KR24]. Can [Zon20, Har24]. Cancer [Foo21]. Cantor [MM21]. capillarity [Val25]. cardinals [Sin22]. career [HW25a, KLW25a, KLW25b]. careers [Cla20, LIJM22]. Carlos [ABC^{+}20]. Carolina [Kas20]. Caroline [PT23, TM23]. Carter [Kas23b]. Cartier [CK25]. case [Pak25, Ude23]. catching [Ken20b]. categorical [Cen22]. Categories [BDS20]. categorification [LS22]. category [BCC^{+}22, BGT24, Rie23]. Catherine [BL24]. celebrates [ABB^{+}22]. Celebrating [DHD^{+}22, DR20]. celebration [MLW^{+}24, Mer20a]. cell [Cio22]. cellular [Kom20]. centuries [Bim23]. century [Dev22, Fer25]. chain [JLL^{+}21]. Chandler [ABB^{+}23]. Chang [GW20]. changed [MMS25]. characterizations [Tor20]. charms [Roh23]. Charney [Lam21b]. chart [dLLOW23]. Chayes [Bon22a]. Cheeger [Ano21k]. chemistry [BCC^{+}22]. Chevalley [Ano20c, He22]. children [Har21]. Chile [STHRVG22]. China [Zon23]. Chip [Jen21]. Christine [Ols23b]. Christoph [SW25]. Chudnovsky [Bon22c]. Chung [But20]. Chuu [Uhl24]. Chuu-Lian [Uhl24]. Cibercoloquio [CRS^{+}21]. Ciprian [BCT^{+}22]. Circles [Cal22a]. civil [MWW^{+}23]. class [Ano24a, Lic22]. classes [Nai21]. classic [Ang24]. Classical [Hol24, Cha21, Dip20]. classification [DH20, Wei20]. Climate [FLOS24]. closing [dLLOW23]. club [Her22e]. Clusters [GS24]. Coates [MLW^{+}24]. Codimension [Per24]. coding [BLSY22]. cohomology [Wit20]. cohort [Lam20, LIJM22]. coin [Mil22]. coin-flipping [Mil22]. Colding [BW24]. Cole [Ano20e, Ano21d, Ano23c]. Coleman [EO21]. collaboration [Gry22, Mer22a]. collected [Bon23]. collection [CL20, HPLQ^{+}21b, Ros20a]. College [BLOS23]. colleges [D’A23]. Coloring [vW22]. Combinatorial [EL25, MMP20, Sch25, Sin22]. combinatorics [Ber22, BHO^{+}24, Che22, Cli21, EK20, Mic23, RTY23, Pak22]. Combining [Ste23]. commentary [Ull22, Bon23]. commissions [Her22c]. communication [Her22g, Spr25]. Communications [Ano20f]. communicators [Clo25]. community [Ham21, JMM^{+}23, Ols23a, Wei20]. commutative [Lup21]. compactifications [GK22]. companion [Cad21]. comparison [Dip20]. Competition [Ull22]. complex [AF22, CWA25, Esn23, Hed20, Xu22]. complexity [BW24, Lan22]. Compressed [BPJ24]. Computational [Awa21, GLY20, Lan22, WGG22]. computers [Buz20]. computing [BCC^{+}22, RN23, Tur24]. Conant [Ano20j, Ano21g, Ano22f, Ano23h]. Concepts [Bim23]. condensate [Kir21]. conditionality [Lav20]. conditions [Mav23]. conferences [AAD^{+}25, BBB21, Mas22]. conformal [GLY20]. conformally [Wan24]. confront [LP24]. Congress [Row22, Sax22]. Congressional [Her21a, Sax20]. conics [BST20]. conjecture [Gar24a, Tsi24, BMP24]. conjectures [Hay21]. connectedness [Ham21]. Connections [Tay24]. consensus [LP20]. Conservation [BKK22]. Considerations [ÁC21, HKL25]. constant [BKK22]. Constraint [FGKS25]. constructions [BKK22]. consultations [Her22c]. Contact [Cas21, Lic22]. contagion [Ald22]. contexts [BBC^{+}22b]. continua [DNF24]. continuation [Ber21, DR23]. continuous [Rut24]. continuous-time [Rut24]. continuum [Gam20]. contributions [CGHP^{+}21, Ste23, Uhl24, ARS23]. contributors [Hou23a]. control [XBH^{+}22]. conventions [Lam25]. convergence [ABLL21, JLL^{+}21]. Conversation [Raj21, BF23, FMM24, vOL25]. Conversational [Oma21]. conversations [Rin24]. Convex [EK20]. Convolutional [Has23]. Conway [Ryb22a, Ryb22b]. cooking [Cas21]. cooperation [Par21]. Corbin [Kas23b]. coronavirus [Dun21c]. correct [GGT23]. correction [BdMG23]. cosmology [Mum21]. Could [Rie23]. count [Dun21a, Ols23a]. Counting [Har21, Wol21, Cla24]. countries [Dun21d]. coupling [JLL^{+}21]. coupling/minorization/drift [JLL^{+}21]. courses [LaR24, MPDL25, Rey24]. Courts [Her22c]. CoV [BPJ24]. covering [BL23, Rub20]. coverings [Ort20]. COVID [WWL^{+}21]. COVID-19 [WWL^{+}21]. Cox [GL22]. Craig [SW20]. creation [GLR23]. Creative [Mer20a]. critical [Mus22]. Crossed [Hue23]. Cryptocurrencies [LP20]. cryptography [KFN23]. cull [D’A23]. Curating [Dev22]. curious [Spr25]. current [DW22]. curvature [BKK22, Sea23]. Curve [Cla24, EST24]. Curve-counting [Cla24]. curves [Bar24b, Cap20, Cha21, Hed20, Jen21]. cuts [Ful24]. cycles [Müt24, Yak20].

D.C [Ros20a]. D’Ambrosio [Sco22]. dance [Sch24]. Dark [Kir21]. Data [JMM^{+}23, BGH20, DNF24, FLOS24, Wol24]. database [GLR23]. David [Ano21e, ACF^{+}24, EW22, GM21]. Davis [ABB^{+}23, CHO^{+}24]. dawning [E21]. deblurring [AEP22]. Deborah [Bas24a]. decades [Hou23a, RRS22]. declaration [Tur22]. decomposition [Fra24]. deep [BHB25, Rut24]. defects [Le 20]. Definability [Pas23, Lan22]. defined [Müt24]. definition [Ang24]. degenerate [CL23]. Demailly [CDX23]. Democracy [CSV21]. Demographic [Yak20]. Dennis [DS23, Ano22c]. department [AK25, Whi21]. departments [Rey24]. Derek [BNS23]. derivatives [BP25, Sti23, Tor20]. Descriptive [Ber22]. descriptors [WGG22]. design [Bab22]. Designing [Hei21, BF21]. Determinantal [DM21]. developing [Dun21d]. development [GLR23, Hou23a, PQY^{+}25]. developments [Bam21, Fis20, RR21]. diagrams [BDE^{+}20]. Diary [Tur22]. diaspora [BG21a]. Dick [CIW22]. diffeomorphism [Man21, Nar24]. difference [Her22a]. Differential [Rut24, BdMG23, Bec20, BG21b, Che23, Mel21, Nag21, Poi22, PT24]. differentials [BP25]. Differentiating [Jef23b]. diffusion [Ros20b, Tei20]. diffusions [Sti23]. digital [Rob23]. dimension [Ten21]. dimensional [WY23]. Dimensionality [Nel20]. Dionne [Inn24]. Diophantine [MM21, VA21]. Dirac [EGT21]. discipline [Far21]. discourse [SM24]. Discrete [Jos20, HB22, Par23, Roh23, Tor20]. discusses [SW25]. disparities [dLLOW23]. disposition [VA21]. disrupting [LIJM22]. dissipation [ZCIM24]. Distinguished [Ano20a]. distributed [Ber22]. Divergence [CT21, MMM20]. Divergence-measure [CT21]. Diversity [Bar24a, GM23a, Ham21]. DNA [BLSY22]. DNA-based [BLSY22]. do [MWW^{+}23, Pak25]. Documents [LV23, CL20]. Dolciani [Ano21h, Ano23i]. Donaldson [Ano20l, HWW21]. Don’t [Dun21a]. DONUT [GLR23]. Doob [Ano20g, Ano23e]. Dorothy [KMWW20].

Double [ACFH24]. Double-anonymous [ACFH24]. doughnuts [Cal20a]. Douglas [Cro21b]. Dowling [BMP24]. down [HN22]. downhill [EST24]. Dr. [IJS22, Inn24, KMMM21, ML24, Raj21, Woo20, dL21]. drama [Poi22]. drift [JLL^{+}21]. dual [OZ23]. duality [Bae23b]. Duren [BCF^{+}21]. Durfee [O’S25]. Dusa [Sch21b]. dynamical [BGH20, FLOS24, LST20, MM21, Mez21, WGG22]. dynamics [Bec20, Ben22, CGHP^{+}21, CM23, Pot21, Tad21]. Dyson [AFS21].

Earl [IMNR23]. Early [Tur24]. Eastern [Bee24]. EDGE [Mer20b]. Edinburgh [Mak20]. editing [ELW25]. editor [Ano20n, AAD^{+}25, TM23]. editors [Her22d]. Eduardo [dL21]. education [RRS22, Rey24]. educator [Jet22, ST22]. Edward [Jor20]. effects [Sti23, TGG^{+}25]. Effros [Jor20]. Egyptian [HP25]. eigenvalue [FHLS20, Zha24]. eigenvalues [CCH21]. Einstein [Kir21]. Eisenbud [Ano20h, Ano23f]. elections [Hei21]. Elements [Zon23]. elephants [Sti23]. Eli [PSAH24]. Elias [BCW23, Pie21]. Eliashberg [Ano20l]. elimination [Urs25]. Elizabeth [Dum22]. elliptic [Bar24b, CL23, De 23, Mav23, Mus22]. emergent [Tad21]. Emma [DAG^{+}23]. enchantress [Ric24]. end [Her22c, Kas20]. engagement [Ude23]. Enolski [Pre20]. entropy [BW24, CP22]. enumeration [CN23]. Enumerative [Cap20, Cli21, Mic23]. eponymous [Mer20c]. equation [BHW22, EGT21, Hyn21]. equations [Bec20, CL23, Che23, Jos20, Maw23, Nag21, Pin24, Ros20b, Rut24, VA21]. Equiangular [Zha24]. equity [Jet22, PQY^{+}25]. equity-oriented [PQY^{+}25]. equivariant [LN23]. era [E21, KFN23]. Erasable [Sil22]. Erdos [Roh23]. ergodic [Ass22]. ergodicity [Dya23]. Eric [Sax21]. errors [GGT23, Mer20a]. Ethical [HKL25]. ethics [CM24]. ethnic [ÁC21]. Euclid [Zon23]. Euclidean [Nel20, Oh22]. Eugenio [Che24]. Europe [Ken20b]. eve [Row22]. Evelyn [Hou24, Tur24]. everyday [ABH^{+}25]. evidence [Lav20]. evolution [Foo21, Kom20]. example [FLU23]. excellence [Ano23i, ST22]. Exceptional [Pol21]. Exciting [BLSY22]. executive [Ano20n, TM23]. Exercise [Cal24b]. expanded [Wei20]. experiences [STHRVG22]. Exploring [Gry22]. exponential [Rub20]. extraordinaire [ALM21, Ano20n]. extraordinary [Kuz20]. Extremal [BK25].

F [Lan22]. F. [GM21]. face [HB24]. faces [Nie22]. families [Cla20, Di 21]. Fan [But20]. far [KLW25a, KLW25b]. favorite [Mer20e]. features [Rui23]. Federal [MMS25]. feeling [Lob24]. Feigenbaum [KLSS21]. Fellow [Her21a]. Fellows [Ano24a]. Fellowship [Her22f, Mer20b]. fewnomial [Roj25]. fibers [Wil20]. fiction [Mak20]. field [CM24, Fou20, Mer20e]. Fields [Ano22a, CT21, vdD22]. figure [KMWW20]. Figuring [Wil20]. Finding [Lau20a]. Finite [GL22]. finiteness [RR21]. firing [Jen21]. first [Ber21, Hou23a, Mar24]. firsts [HN22]. five [Hou23a, Poi22]. fix [Fer21]. flexibility [HS24]. flip [Sch25]. flipping [Mil22]. Floer [Gre21, PCR22]. flow [Cam23]. flows [Bam21, Cam23, DS20a, Dip20, PST23, THSA23, ZCIM24]. fluid [Cam23, CGHP^{+}21]. fluids [Vas24]. Foias [BCT^{+}22]. Foliations [AF22, Nar24, Per24]. Fontaine [BIK^{+}20]. forcing [FHLS20, HB24]. Ford [ML24, Her22a]. forecasting [WWL^{+}21]. formal [ACMT25]. forms [Poi22, Pol21]. formula [Cal20b]. formulas [CT21, Spo22]. forwards [Ran23]. Foundation [Her22a]. foundations [TSAY22]. Four [BCK24, Kuz20]. Fourier [Sti23, Tor20]. fractal [Rui23, Tay24]. fractal-like [Rui23]. Fractional [Sti23, HB22, Tor20]. framework [BCF^{+}24]. Francoise [Bas24b]. Frank [Ano20d, Ano20e, Ano21c, Ano21d, Ano23b, Ano23c]. free [BdMG23, De 23, FGKS25]. Freeform [Maw23]. Freeman [And24, AFS21]. French [CL20]. Friedlander [CGHP^{+}21]. FriesenPress [Gar24b]. front [Cas21]. fruitful [Mer22a, RR24]. full [Her22f, San21]. fully [ILM^{+}22]. function [Mor24b, Tor20]. functional [DNF24]. functions [Gal23, Tre25b]. funding [MMS25]. funds [Mer20b]. Furstenberg [Ano20m, DS21b]. future [KN22, Wei20, Nat23].

G [Jor20]. game [RR22b, XBH^{+}22]. games [Fou20, PT24, Clo22, DeL22]. gap [FLU23]. gaps [dLLOW23]. Gauss [Bae21b, CT21]. Gaussian [Urs25]. Gelfand [Far22]. Gems [HR22]. gender [FLU23]. general [Gna21]. generation [Dun23, GL22]. generations [BBB^{+}24a]. generative [HKL25, SM24]. generosity [Dum22]. genius [Kuz20]. gentle [McF22]. gentleman [HWW21]. geodesics [Wol21]. Geography [Dun21b]. Geometric [Cla22, Awa21, DS20a, Dip20, GW20, McF22, TJ23, VA21]. geometries [Gan22]. Geometry [Sun22, AM20, BHB25, BdMG23, Ber20, Cap20, Dic20, EK20, GGT23, GLY20, HMR^{+}25, HM21, Lic22, Lup21, Mar23, Mel21, Mez21, Mic23, Nie22, PT23, Poi22, Ran23, Rit23, RR24, ST20, Sch21b, Sut21, Tay24, Uhl24, Wan24, Wer24, Xu22]. George [Ano21e, Ano22d]. Georgia [AHRZ23, HR22]. gerrymandering [Her22c]. Gift [Mer20b]. Gilmer [ML24]. Giovanni [Par21]. give [GM23b, Shi25]. given [EST24]. Gladys [KMMM21]. Gloria [ML24]. gluing [BKK22]. Goble [EO21]. Gold [Sax20]. Goldstein [BL24]. gonality [DV21]. good [Mer21, RR21]. Goodman [PSAH24]. Goro [BBH^{+}20]. got [MWW^{+}23]. Göttingen [Lor20]. government [AHJ^{+}22]. Grabiner [DR20]. Grace [GW22, HN22]. gradient [THSA23]. graduate [Can23, DW22, IJS22, Lam20, PQY^{+}25]. Graham [BBS21]. Granville [Hou24, Tur24]. Graph [CM23, HK21, FHLS20, Gry22, Ram22, TSAY22, Mar24]. graph-based [TSAY22]. graphical [Bab22]. graphics [Mer22d]. graphs [DV21, Müt24, Dav25]. gravity [Gwy20]. Gray [GM23b]. great [Cal22a]. Green [CT21]. Gregory [DS21a]. Grenander [Ano21j]. Group [KFN23, Cla22, Gna21, Lic22, OZ23, RR24, ST20, Weg20, Wei23b]. Group-based [KFN23]. groups [ÁC21, Man21, Mel21, Nar24, Pol21, RR21]. GROWing [Lam20]. growth [Hou23a, Tre25b]. Guccia [Par21]. guide [CM24]. Guiding [PQY^{+}25]. Guido [HW23]. Guy [GP22].

H [O’S25]. Haken [CKL^{+}23]. hallmarks [Kuz20]. Hamilton [Hyn21, Müt24]. Hanlon [HW25a]. Hardy [vdD22]. Harington [HW25b]. Harold [BBC^{+}22a]. Harry [GL20]. Hartwell [Raj21]. Hasselblatt [Cro21a]. head [Whi21]. heart [Bim23, Dum22, Her22d]. Heather [JSS23]. Heegaard [Gre21]. held [Bee24]. help [GGT23, dLLOW23]. Henry [BBE^{+}25]. her [KLW25a, KLW25b]. Hermann [CGH^{+}21]. Hex [San21]. hexagonal [Bae24a]. hey [Clo25]. hidden [KMWW20]. hierarchical [Sch21a]. Higgs [Sch20]. higher [RRS22]. Hilbert [Shl21]. Hillel [DS21b]. Hispanic [HPLQ^{+}21b]. Hispanics [HLQ^{+}20, HPLQ^{+}21a]. historical [Edw24]. histories [BL24]. history [Ros21, Ste23]. Hochschild [Wit20]. Hodge [EL25]. holes [Gio23, HL22]. homeomorphism [Man21]. homology [Gre21, Lob24, PCR22]. Homotopical [BHO^{+}24]. homotopy [DR23, Lev20]. honeycomb [Bae24a]. honored [Sax20]. Honoring [Bee24]. Hoover [KMWW20]. Hörmander [Nai21]. Horn [Gar24a]. horse [RR22b]. Horton [Ryb22b]. Hrabowski [And24]. Hrushovski [Ano22b]. Hughes [Hug24]. human [Jef23a]. hundred [Mar24]. hybrid [AAD^{+}25]. Hyman [BCL25]. hyper [AHJ^{+}22, BKK22]. hyperbolic [Oh22, PT23, Pot21, Wol21]. Hypergeometric [RR22a]. hypermatrix [Gna21]. hyperpolygon [RS21].

ice [GBC^{+}20]. ICM [DDV^{+}20].

icosahedron [Bar24b]. icosidodecahedron [Bae23a]. ideal [ABLL21]. ideals [Nor20]. identical [Zon20]. identities [All20]. identity [BL24]. Igor [BEO^{+}24]. IHES [FLU23]. II [KLW25b, Fou20]. Ilie [BCT^{+}22]. illustration [CDHH^{+}24]. image [AEP22]. imaging [Has23]. immersions [HL24]. impact [Her22f, Tur22]. impacts [KN22]. implementation [ACFH24]. implosion [PV20]. importance [CDHH^{+}24]. IMU [Keh22, Ken20a, SW25]. inclusive [ILM^{+}22]. incompressible [ZCIM24]. increase [Ham21]. increasing [ÁC21]. index [CM21]. Indigenising [Tuf22]. indomitable [Lup21]. inductive [Nie23]. industry [AHJ^{+}22]. inequality [BE24]. infectious [Yak20]. inference [PST23]. Infinite [Sin22, Agu24]. information [Nie22]. inimitable [GM21]. inspiration [Inn24]. inspiring [Hau24]. Instability [BP20]. instanton [PCR22]. Institute [CSV21]. institutional [FLU23]. institutions [RRS22]. instructors [HKL25, Her22b]. integrable [Ros20b]. integral [Row22]. integrals [Skr20, Tor20]. integrity [SM24]. interaction [RR24, Xu22]. interactions [Cam23]. Interdisciplinary [Edw24]. interface [FLOS24]. Interfacing [Ude23]. international [Par21, Ken20a, RN23]. internationally [Hou23b]. interplay [ZCIM24]. intersecting [Müt24]. Interview [BCL25, Bee25b, Bon22a, Bon22b, Bon22c, Che22, DS20b, DS21a, DS21b, DS22, DS23, Bee25a, Cro21a, Cro21b, DS24, DS25, Her21b, Lam21a, Lam21b, Sax22]. intriguing [Ros20a]. Introducing [HPLQ^{+}21b]. introduction [DNF24, McF22, Roj25, Yan24, Pom25, Roh23]. invaluable [Her22d]. invariant [BSV23, Wan24]. inverse [FHLS20]. invisibilized [LDAST24]. invitation [All20, CKR21, LS22]. involvement [Tur24]. IRS [Fer21]. Isadore [BCG22a, BCG22b]. Isbell [Bae23b]. Isometric [HL24]. isoperimetric [BE24]. Israel [BGH^{+}23]. iteration [Fer21]. IUM1 [Tuf22].

J [Ano24b, DFP^{+}22, ESSV23]. Jacobi [Hyn21]. Jacques [Wei23a]. James [HWW21, Kas20]. Jane [Ano20n]. January [DFP^{+}22, Zab25, MWW^{+}23]. Jay [IMNR23]. Jean [Ano21o, BIK^{+}20, CDX23, CK20]. Jean-Marc [BIK^{+}20]. Jean-Pierre [CDX23, CK20]. Jeffrey [Bon22b]. Jenda [ALM21]. Jennifer [Bon22a]. jeopardy [SM24]. Jerry [Sax22, BKM23]. Jewett [BGH^{+}23]. Jill [Lam21a]. Jim [Ano25b]. JMMs [Wei20]. John [ABC^{+}21, MLW^{+}24, Nat23, Ryb22b]. Johnson [EO21, IJS22]. Joint [Ano20f, BG21a]. joke [Mer20d]. Jones [BEKP21]. Joseph [Ano20g, Ano23e, Ano24b, Kas23b]. Josephine [RN23]. journals [ACFH24]. journey [Des21, Her22e, IJS22, Kas23a, Zon23]. Journeys [Csi24]. Józef [Kuz20]. Jr. [BBE^{+}25]. Judith [DR20, Zab25]. Jue [RN23]. Julia [Lan22]. July [MWW^{+}23]. June [BNS23]. just [Clo25]. justice [Her21a, JMM^{+}23, LOW23, Law20].

Kähler [Wan24]. Kardar [Gan22]. Karen [DHD^{+}22, DS20b, Mer20b]. Katherine [EO21]. Kazhdan [Ano20k]. keep [Lau20b]. Ken [DHMV22]. kernels [Ros20b]. Kesten [GL20]. Keyfitz [FMM24]. Khovanov [Lob24]. kinematic [Cal20b]. Kister [Ano20n]. Klawe [KLW25a, KLW25b]. Kloefkorn [Her21b]. Knight [Lan22]. Knot [Hed20]. knots [Pin24]. know [Can23, Ele24]. Kohn [Ano24b]. Kolmogorov [Peh22]. Konstantinov [ST22]. Koopman [Mez21]. Kra [Bee25a]. Krichever [BEO^{+}24]. Krylov [Gre23]. Kuku [Hou23b]. Kummer [Dol20]. Kunen [DHMV22]. Kuramoto [BHW22]. Kuranishi [PS22].

L [Ano20g, Ano20j, Ano21g, Ano22f, Ano23e, Ano23h, BL21, CFKS22, HW23, IJS22, Inn24, Kas20]. Label [LDAST24]. labyrinth [Gam20]. Ladie [Tur22]. lady [Bie20]. Ladyzhenskaya [DDV^{+}20]. Lagrange [MM21]. Lagrangian [WGG22]. Landau [vOL25]. Lander [Sax21]. lands [Mer20d]. language [BGT24]. large [Cal23a, LaR24, TGG^{+}25]. last [MMS25, RRS22]. László [DS22]. later [ILM^{+}22]. Lathisms [HLQ^{+}20, HPLQ^{+}21a]. Latinoamericano [CRS^{+}21]. Latinx [HPLQ^{+}21b]. Latinxs [HLQ^{+}20, HPLQ^{+}21a]. lattice [CN23, Cli21, Far22]. Laureate [DS21b, DS23, DS20b, DS21a]. laws [BKK22]. leaders [Hou23a]. leaders/contributors [Hou23a]. leadership [Hau24, Inn24]. learning [BHB25, Bas24a, BSV23, Fok20, Mez21, Rut24, TSAY22, XBH^{+}22]. leaving [Inn24]. lecture [Her22b]. Lectures [Vel22]. Lee [FMM24, Raj21]. Lefschetz [HM21]. Legacy [Liu22, ACD^{+}21, Che24, CIW22, Hou24, Inn24, KMMM21, Kuz20, Tur22]. Leibniz [Tor20]. lens [RR24]. Leon [Ano21a]. Leonard [Ano20h, Ano23f].

Leroy [Ano20i, Ano21f, Ano22e, Ano23g]. letter [Wei20, AAD^{+}25]. Levi [Ano20j, Ano21g, Ano22f, Ano23h]. Lewis [BBS21]. LGBTQ [BBB21]. Lian [Uhl24]. Lie [Ano20c, He22, Ang24]. life [ACD^{+}21, But20, DMW22, KMMM21, MLW^{+}24, Sch24]. lifelong [Inn24]. Liggett [ACD^{+}21]. light [KMWW20, Yan24]. like [Kas21, Rui23]. likelihood [Lav20]. limited [Tre25b, vW22]. limits [SSU23]. linear [BGT24, DM21, Fie20, Gna21, RR21]. lines [Zha24]. Liouville [Gwy20]. Lis [Ste23]. Lisa [Bon22b]. literature [SM24]. lives [Cal20b]. local [Esn23, Her22e]. Loewenberg [Bas24a]. logic [Ans24]. Loh [Mer21]. long [BL24, Val25]. long-range [Val25]. long-term [BL24]. longer [Mer22b]. look [DW22]. Lorenz [Pin24]. Louis [Ano21l]. Lovász [Ano21m, DS22]. loved [GP22]. Lovelace [Ric24]. loves [Far22]. Low [CCL^{+}22]. Low-rank [CCL^{+}22]. Lowell [Ull22]. lower [Sea23]. Lucas [Bae21b]. Lucien [EGMD^{+}21]. Luis [DS24]. Luna [dL21]. Lusztig [Ano22d]. Lynne [JSS23]. Lyttle [Ano21i, Ano23j].

M [ACD^{+}21, BCW23, BCG22a, BCG22b, Inn24, Pie21]. M. [BGPH^{+}22]. MAA [Ull22, DeL22, Ols23a]. Machine [BSV23, Tao25, Fok20, XBH^{+}22]. Machine-assisted [Tao25]. Machines [Kas21]. MAD [Di 21, Goi21]. magical [HM22]. Majda [ESSV23]. major [Wol24]. makes [Her22a, Her22d, Mer20d, Mer21]. Making [Csi24, LV23]. man [GP22, Nat23]. managing [LaR24]. Mandelbrot [Mer22d]. manifold [Pot21]. manifolds [AF22, Czu24, Per24]. many [BBC^{+}22b, Far22, HB24, HN22, Nie22]. map [Bae21a, Ort20]. mapping [Lic22]. maps [FGKS25]. Marc [BIK^{+}20]. March [Zab25]. Marcinkiewicz [Kuz20]. Margulis [Ano20m, DS21a]. Maria [Bon22c, KLW25a, KLW25b]. Marie [Bas24b]. Marie-Francoise [Bas24b]. Markov [JLL^{+}21, MM21]. Martin [CHO^{+}24]. Mary [Ano21h, ADG^{+}24, Ano23i, GM23b]. Masamu [JL21]. Masatake [PS22]. mass [HL22]. mastering [Cas21]. Mastery [BWW24]. Matemáticas [CRS^{+}21]. materials [Le 20, Sti23]. Math [Dun21c, Hau24, Sch24, BCF^{+}24, BCK24, Can23, D’A23, GPS22, Har24, Her22b, Her22e, Her22g, Mer20d, Mer22d, MWW^{+}23, Shi25, Tho20, Wol24]. Math. [Ros20a, Ros21]. math/art [Har24]. mathemalchemica [GPS22]. mathematica [ELW25]. Mathematical [BBB^{+}24a, Far21, Ken20a, LP24, LLM21, TSAY22, Ull22, ACD^{+}21, BL24, But20, CL20, CGHP^{+}21, Cla20, CKR21, CDHH^{+}24, Gry22, HLQ^{+}20, HPLQ^{+}21a, Har21, Her22a, IJS22, JMW24, Kha25, Lam25, Mas22, PA24, Poi22, Ric24, Roh23, Ros20a, ST22, Wei20, dLLOW23, ACFH24, Dun21b, Her22d]. Mathematician [Mer22b, Mer22c, Bai23, BFH^{+}21, Bie20, CK25, Hou23b, HN22, HWW21, Kas23b, Mer20a, Sax22, Gar24b]. Mathematicians [Hou23a, LP24, Mer20c, Row22, ARS23, HPLQ^{+}21b, Has23, Her22c, Her22d, LIJM22, Csi24]. Mathematics [Ano20h, Ano20o, Ano21c, CSV21, DH20, EMS22, Kom20, Le 20, Lop21, Rey24, Tuf22, ACFH24, AK25, ÁC21, And24, Ano22g, Ano23f, Ano23j, AAD^{+}25, BG21a, Bas24a, Bim23, BBB21, BLOS23, CM24, Clo25, DeL22, Dev22, DW22, Dum21, DKR22, E21, Edw24, ELW25, Ful24, Gia22, Gio23, Gry22, Ham21, HKL25, HS24, Her21a, HP25, Hug24, ILM^{+}22, Jet22, JMM^{+}23, Ken20b, LaR24, LOW23, Lor20, MMS25, Mil22, Mur21, Mur23, Nor20, PQY^{+}25, RRS22, Rob23, Ros22, Ros21, Ryb22a, STHRVG22, Tad21, Ude23, Yan24, Zha21, Ano20d, Ano20f, Law20, Mar20, Des21, Hau24, Mak20, Ols23a, Par21, Vel22]. Mathématiques [CL20]. Mathematizing [Jef23a]. MathSciNet [Dun21d]. MathSciNet(R) [Dun21b, Her22d]. MathSciNet(R)/Mathematical [Dun21b]. matrices [AEP22, CN23]. matter [Kir21]. may [GGT23, Kir21]. MB [Gar24b]. McBay [Inn24, Woo20]. McDuff [Sch21b]. McKean [BBE^{+}25]. MCMC [PST23]. McNerney [Sax22]. me [Ols23a, Kas21]. mean [BKK22, Cam23, Fou20, Nie23]. meaning [BGT24]. measure [CT21]. measuring [Rat23]. Meckes [Dum22]. Medals [Ano22a, Sax20]. medical [Has23]. Meeting [Bee24, Row22]. meets [BCW23, Mic23, Ram22]. melodist [Mer22c]. memoir [Kha25]. memoirs [Bie20]. Memorial [ABC^{+}21, Ano20b, Ano21a, Ano23a, CGH^{+}21, ACF^{+}24, Bee24, BBB^{+}24b, EO21, DAG^{+}23, FS20, Pie21]. Memoriam [AFS21, BL21, BCG22a, BCG22b, CFKS22, FM20, Kuz20, O’S25, Sco22]. Memories [BEKP21, BBH^{+}20, KR24, SKT^{+}24]. memory [ADG^{+}24, Ano23d, BEO^{+}24, BCF^{+}21, CHO^{+}24, DHMV22, ESSV23, IMNR23, Sti23]. Mendel [Ros22]. mentor [ALM21]. metaplectic [Wei23b]. method [Maw23, Rub20]. methods [AHJ^{+}22, AEP22, Gre23]. metric [BCH^{+}20, DV21, Gan22, Rit23]. metrics [Sun22]. Mexican [ARS23]. Michael [BBB^{+}24b]. Michel [DS25]. mid [Fer25]. mid-20th [Fer25]. Mikhail [Ano21n]. Mikio [SKT^{+}24]. mind [Clo22]. Minicozzi [BW24]. minimal [CM21, De 24, Dip20, Mor24a]. Minkowski [Maw23]. minor [Ram22]. minoritized [ÁC21]. minorization [JLL^{+}21]. minuscule [Gre20]. mirror [Cla24, WY23]. misinformation [Gia22]. missed [Can23]. Mitchell [KLSS21]. mixed [Che23]. Mixing [Mar23, ZCIM24]. MLBDecideR [OWD20]. Model [Fok20, Nag21, BLOS23, Mor24a, Peh22, vdD22]. Modeling [GBC^{+}20, LST20, WWL^{+}21, dLLOW23, Cio22, Le 20, LLM21, Vas24]. Models [AHJ^{+}22, Sch21a, Yak20, dLLOW23]. modern [GLY20, Tho20]. Modernism [Mak20]. modular [Pol21]. modules [Hue23]. Moduli [Cha21, Bae24b, GK22, RW22]. Monge [Awa21, Maw23]. monodromy [Weg20]. Monotonicity [Spo22]. Montgomery [BL21, Kas23a]. Moore [KR24]. Morgan [Ano20d, Ano21c, Ano23b]. Morse [CM21]. Moses [MWW^{+}23]. motion [DKR22]. motives [RR22a]. motivic [Lev20]. movement [LOW23, MWW^{+}23, Sch21a]. Moving [MPDL25]. MR3642967 [Nor20]. MR3675486 [Ken20b]. MR3753659 [Lor20]. MR3967051 [Law20]. MSRI [BG21a, Ham21]. MSRI-UP [Ham21]. multilevel [AEP22]. multiplicities [Zha24]. Multiscale [Vas24, CCL^{+}22]. museums [BCK24]. music [Ude23]. my [Spr25].

N [GM21]. NAM [Hou23a, Hou23a]. Nan [Kas21]. Narasimhan [BGPH^{+}22]. narratives [BGT24]. national [KN22, Hou23a]. nature [Spr25]. Navigating [AK25, KMMM21]. nearly [Ham25]. need [Can23]. Nelson [Ano20e, Ano21d, Ano23c]. network [AHJ^{+}22, CM23, LN23]. networks [Agu24, BCC^{+}22, BCJ^{+}25, DKR22, Has23]. Neugebauer [Dun20]. Neumann [Nat23]. neural [CM23, Has23, LN23]. never [GM23b]. newest [BCK24]. next [Dun23]. Nigel [EW25]. Nigerian [HN22]. Nikolaevna [APS22]. Nikolay [ST22]. Nina [APS22]. Nirenberg [Ano21l]. no [Jun24, Mer22b]. Nobel [Gam25]. Non [Lup21, Ben22]. non-Archimedean [Ben22]. Non-commutative [Lup21]. noncommutative [Kas23a, Rey25]. noncommutativity [Skr20]. nonimaging [Awa21]. nonlinear [Mav23, Peh22, Tei20]. Nonlocal [Ros20b, Dip20]. nontangential [MMM20]. Norbert [Ano22g]. normal [CT21]. Normalizing [PST23]. Norton [Nat23]. Notes [Kha25, Her22b]. Notices [Bar24a]. nowadays [Ang24]. Number [Ano20e, Ano23c, HK21, Ric24]. numbers [Jef23b, Lon23]. Numerical [Urs25, DM21, HMR^{+}25]. NY [Ald22, Kas21, Nat23].

O [Hou23b]. oand [Wil20]. Obamacare [Fer21]. object [PL25]. obstacle [Dan20]. October [DFP^{+}22, Goi21]. of4205950 [Vel22]. off [McF22]. off-roading [McF22]. offer [Rey24]. Ola [BDE^{+}20]. oldest [BCK24]. Olga [DDV^{+}20, Goo20]. One [Her22e, ILM^{+}22, Per24]. Oort [Tsi24]. Open [Gre23, Har23, Her22b]. opening [HN22]. Operads [BF21]. operator [Czu24, Mez21, Skr20, Agu24]. operators [Nai21]. opportunities [GLR23]. optical [Wer24]. optics [Awa21, Maw23]. Optimal [RR22b, Fok20, Maw23]. optimism [ML24]. optimization [BL23, HK21, THSA23]. Or/oand [Wil20]. order [Ber21, Tre23]. oriented [PQY^{+}25]. orthogonality [Tor20]. Oslo [Row22]. Otto [Dun20]. Outstanding [Ano20d, Ano21c]. Overtoun [ALM21]. overview [Dan20].

P [Ano20i, Ano21f, Ano21h, Ano22e, Ano23g, Ano23i, BBE^{+}25, Bon23, Che22]. pages [Goi21]. Painlevé [Jos20]. paintbox [vW22]. Painting [WGG22]. Pandemic [XBH^{+}22]. papers [Pak25]. Parabolic [PV20, Mus22]. paradox [GN22]. parallel [BGT24]. parameters [FHLS20]. parametric [Sch21a]. Parisi [Gan22]. parking [Mor24b]. Parshall [DHD^{+}22]. part [KLW25a, BCG22a, BCG22b, KLW25b]. Partial [Che23, Bec20]. Partially [Pot21]. participation [ÁC21]. partnership [JL21]. party [Jun24]. passing [Hou23a]. past [KN22]. path [Cal23b, CN23, KMMM21]. paths [HN22]. pathways [dLLOW23]. patterns [Pud20, RRS22]. Paul [Roh23]. pave [Zon20]. PDEs [BdMG23, Ber21, CCL^{+}22, Cio22, FMM24]. pebbling [HK21]. peer [ACFH24]. pentagonorum [Cal21a]. Pentagonum [Cal21a]. people [Ano20o]. perception [Jef23a]. perfect [Ham25]. periodic [Gal23, Pud20]. permutation [Pud20]. Perseverance [EO21]. Persistence [KM23]. personal [GL20, BCG22b]. perspective [CL23, FLU23, OZ23, PA24, TJ23]. pervasive [LDAST24]. Peter [Bai23, BL21, BCF^{+}21]. Petersburg [DDV^{+}20]. phase [BE22, Dip20, WGG22]. PhD [Rey24]. PhDs [Mur21, Mur23, RRS22]. phenomena [Par23, Rui23]. Philip [HW25a]. philosophy [BG21b, Vel22]. photonics [PA24]. Physical [Pak25]. physicists [Pak25]. physics [Ano23f, Ano20h]. Picador [Kas21]. picture [Lor20]. pictures [Dya23]. Pierre [CDX23, CK20, CK25]. pioneer [Par21]. Pipher [Lam21a]. piranha [TGG^{+}25]. place [Tho20]. plague [BFH^{+}21]. plane [Cap20, Gan22, Zon20]. planning [BDS20]. plasticity [Ude24]. Pless [DMW22]. Po-Shen [Mer21]. point [DM21]. Pointless [Cal24b]. Poisson [Fie20]. policing [JMW24]. Policy [Ano20f]. political [LP24]. Pollack [PSAH24]. polymers [Cli21]. Polynomial [DR23]. polytopes [EK20, Dav25]. pond [TGG^{+}25]. pool [BPJ24]. population [PT24, Yak20]. portrait [WGG22]. posets [KM23]. positive [CP22, HL22]. possible [Her22d, Hol24]. post [Lau20b]. post-quantum [Lau20b]. postdocs [Her22a]. potential [Her22f]. power [Ans24]. practice [Lop21, Nor20]. practices [Tie22]. prediction [Fok20, Ham25]. Predictive [JMW24]. Presbyopia [BdMG23]. presence [Sut21]. Press [DeL22, Mak20, Ols23a, Ull22]. Previato [DAG^{+}23]. Price [Inn24]. primary [Edw24]. prime [Jef23b]. primer [Har23, Has23, JMW24]. Princ [ELW25]. Princ-wiki-a [ELW25]. principle [HM21]. principles [Lav20, PQY^{+}25, Sin22]. printing [LST20]. privacy [BG21b]. Prize [Ano20c, Ano20d, Ano20e, Ano20h, Ano21c, Ano21d, Ano23c, He22, Rib21, Ano20k, Ano20b, Ano20l, Ano20m, Ano20g, Ano20j, Ano21a, Ano21k, Ano21b, Ano21e, Ano21g, Ano21m, Ano21h, Ano21i, Ano21j, Ano22b, Ano22c, Ano22d, Ano22f, Ano22g, Ano23a, Ano23b, Ano23e, Ano23h, Ano23f, Ano23i, Ano23j, vOL25]. prizes [Gam25, Ano20i, Ano21f, Ano22e, Ano23g, Ano25a]. Probabilistic [Fou20]. problem [AEP22, Dan20, FHLS20, LDAST24, MN24, Oma21, TGG^{+}25, Shl21]. problems [BLSY22, BL23, De 23, GP22, Gre23, Mav23, Mus22, Ull22]. procedures [Lav20]. process [AK25, Sch24]. processes [DM21, Fie20, Tei20]. producing [BLOS23]. product [Weg20]. profession [ILM^{+}22]. professional [Her22e, PQY^{+}25]. Professor [Hou23b]. profound [Kas23b]. program [Dun21d, Fis20, JL21, Mor24a, Wol24]. programs [D’A23, DW22, Fie20, IJS22]. projections [Tay24]. projective [BSSV21, Per24]. Promoting [STHRVG22]. proof [ACMT25, BBB^{+}24a, MPDL25, Tao25]. proof-based [MPDL25]. Proofs [Cal22b, Rin24]. properties [Esn23, HM22, RR21, RW22]. property [Di 21]. Prospering [Dum21]. protein [Cio22]. protocols [LP20]. Providence [DeL22, Hau24, Ols23a, Ros21, Ull22]. Proving [Buz20]. Prym [Ort20]. pseudodifferential [Nai21]. pseudonymous [Mer22b]. Public [Ano20a]. Publication [CL20, Her22e]. publish [Pak25]. publishing [HW25b]. Putnam [Ull22, Ull22]. puzzles [Ros20a]. quantification [CWA25]. Quantum [Dya23, Gwy20, KFN23, Lau20b, Mil22, Pom25]. queen [Des21]. Queer [Cla20]. question [BL24]. questions [Nat23]. quintic [Bar24b]. quixotic [Mer22d].

R [ABB^{+}22]. Racial

[Jet22, ÁC21, LOW23, dLLOW23]. Ramanujan [All20]. Ramia [dL21]. Ramsey [Di 21]. Random [Gan22, Gwy20, Rui23, Par23, Wan24, Wol21]. randomized [DM21]. randomness [Tre23]. range [Val25]. rank [CCL^{+}22]. Rao [ABB^{+}22]. Rate [Ude24]. Rate-based [Ude24]. rates [JLL^{+}21]. rating [KN22]. Rational [Tre25a]. Ravi [Bee25b]. Real [HMR^{+}25, BSSV21]. really [Shi25]. Reassessing [Ric24]. recognized [Hou23b]. Recollections [BDF^{+}20, BCG22b]. recommendations [Jet22]. recurrent [CM23]. Red [LP24]. reduction [Nel20, Peh22, RR21]. reflecting [DKR22]. reflection [Hou23a]. Reflections [BFH^{+}21, IJS22, Jet22]. reflects [HW25a, KLW25a, KLW25b]. Regularity [Tei20, CL23]. related [FHLS20]. Reliability [Men20]. Remembering [AHRZ23, ABC^{+}20, Ano21o, BGPH^{+}22, CHJ21, EW22, Hou23b]. Remembrances [ABB^{+}23, BNS23, BBC^{+}22a, BCT^{+}22, Jor20, SW20]. renaissance [GM21]. René [CL20]. representation [BK25, EO21, Gre20, Sam20]. Representative [Sax22]. reprint [DS21b]. reputation [Ric24]. required [Rey24]. Research [Ano20d, Ano21c, BCJ^{+}25, ABH^{+}25, Ano23i, CDHH^{+}24, Her22d, KN22, MMS25]. researchers [Mas22]. residency [Mer22a]. results [Lam25]. retrieval [BE22]. Review [Pak25, ACFH24, Agu24, Ald22, All20, ARS23, Bie20, Bim23, Bon23, Can23, Cen22, Clo22, Dav25, DeL22, Des21, Fer25, Fou20, Gal23, Gar24b, Hau24, Hay21, Kas21, Kaw23, Ken20b, Law20, Lon23, Lor20, Mak20, Mar24, Mar20, Nat23, Nor20, Ols23a, Oma21, Pak22, Par21, Poi22, Pom25, Rib21, Roh23, Ros20a, Ros21, Row22, San21, Shl21, Sil22, Tur22, Ull21, Ull22, Vel22, Wes23, Wil20]. Reviews [Dun21b, Her22d]. Rhonda [Hug24]. RI [Hau24, Ull22, DeL22, Ols23a, Ros21]. Ricci [Bam21]. Richard [CFKS22, Che22, GP22]. richer [Lor20]. Ricky [PSAH24]. Riemannian [Czu24]. right [Far22]. rights [MWW^{+}23]. Rigidity [Mel21]. rings [GL22]. Riordan [CN23]. rise [BBB21, CWA25]. roading [McF22]. Robert [BGH^{+}23, CGH^{+}21, HW25b, ORT22]. Robin [Liu22]. Robinson [BNS23]. robots [AM20]. robustness [TJ23]. Roger [FS20]. Rogers [All20]. role [LOW23]. Rolling [Seg21]. Ronald [BBS21]. Root [Ang24]. roots [BCD23, Cal23b]. Roy [Bas24b]. Rudin [BCW23]. rules [Ald22, CM23, RR22b, Tor20, Ude24]. Ruminations [Mum21]. Running [Mor24a]. Ruskai [ADG^{+}24]. Russell [Ano21b, vOL25]. Ruth [Ano21i, Ano23j, DKR22, Lam21b].

S [BHL^{+}21, BGPH^{+}22, BDF^{+}20, HB24]. SACNAS [Mil23]. saints [Cal20b]. Sally [Zab25]. salmon [Yak20]. sampling [THSA23]. SARS [BPJ24]. SARS-CoV-2 [BPJ24]. Sartorial [Mer20a]. Sato [SKT^{+}24]. Satter [Ano21i, Ano23j]. say [Har24]. Scare [LP24]. scheduling [BDS20]. scholar [Hou23b]. scholarly [SM24]. scholarship [Gry22]. school [Can23, GW22]. science [AHJ^{+}22, BGH20, FLOS24, JMM^{+}23, Spr25, Ste23]. sciences [HLQ^{+}20, HPLQ^{+}21a, Her22a, IJS22, Mas22]. scientific [BCG22a, GL20, SM24, Spr25]. sea [GBC^{+}20]. search [Czu24, Mer22d]. searches [SM24]. Second [Row22, BST20]. secrets [Lau20b]. Sectional [Bee24].

Seeing [Weg20]. Seely [TM23]. segregation [Kas20, Whi21]. Selected [Lam25]. selection [Fok20]. semi [TSAY22]. semi-supervised [TSAY22]. sensing [BPJ24]. sensing-based [BPJ24]. separation [GGT23]. September [Tuf22]. series [Ans24, PT23]. servant [Des21]. Service [Ano20a]. Seshadri [BHL^{+}21]. session [Bee24]. sessions [Wei20]. set [Müt24]. sets [MM21]. shadows [Gam25]. Shape [BL23]. shaping [ABH^{+}25]. share [Her22b, Her22d]. shared [Fie20]. sharp [MMM20]. Shaw [Ano20k, Ano22b, Ano21k]. Shen [Mer21]. Shimura [BBH^{+}20]. Shining [KMWW20]. SHINY [OWD20]. Shirley [Inn24, Woo20]. Shock [FMM24]. show [MPDL25]. shrinking [Cal20a]. Shubin [Ano21n]. sign [Row22]. signatures [Rob23]. Simons [Ano25b]. simple [Hol24, Roh23]. simpler [Hol24]. Simplicity [Nor20]. Singer [BCG22a, BCG22b]. Singular [Gam20, Tor20]. Sivashinky [BHW22]. Six [Cal21b, MMS25]. size [PT24, Wol24, Kas21]. skew [Ven24]. skills [GM23a]. Smale [Raj21]. small [Cal22a, TGG^{+}25]. Soc [Hau24]. social [BCC^{+}22, Her21a, JMM^{+}23, Law20]. society [AM20, ACFH24, CL20, LP24]. solitons [Lau20a]. Solomon [Kas20]. solutions [DS20a, Ull22]. solving [Oma21]. Some [D’A23, Esn23, Gup23, Lev20, RRS22, Har21, IJS22, Kom20, Rub20]. Sorger [SW25]. source [Ros21]. source-based [Ros21]. sources [Edw24]. South [Kas20, KN22]. Space [RN23, Bae24b, Nel20, Tur24, Wan24, Wol21]. spaceflight [RN23]. Spaces [BBB21, BCH^{+}20, Cha21, GK22, Rit23, RW22, Sea23, Tor20, Gal23]. sparks [Mer22a]. sparse [AHJ^{+}22]. spatial [Bec20]. spatially [Foo21]. spectra [MM21]. Spectral [Bec20, Sut21, Rey25]. spectrum [Mav23]. Spelman [BLOS23]. sphere [Bae25]. Spivak [BBB^{+}24b]. spline [GW22]. spread [Ald22]. Spring [Bee24]. squares [BSSV21]. St. [DDV^{+}20]. Stability [RW22, Bec20, Sam20, Urs25, Xu22]. stable [Gio23, Hei21]. stages [Tur24]. stand [Mer20d]. stand-up [Mer20d]. Stanley [Che22]. States [RRS22]. Station [RN23]. Statistical [Sch21a, Fok20]. statistician [GM21]. statistics [Mic23]. status [Hug24]. Steele [Ano20i, Ano21f, Ano22e, Ano23g]. Stein [BCW23, Pie21]. Steklov [Mav23]. STEM [ALM21, And24]. Stepen [DMW22]. stereotypes [LIJM22]. Steve [Raj21, Ano23d, CHJ21]. Stewart [Her21a]. Stochastic [DKR22, GGT23, Cio22, CKR21, PT24]. stop [Ald22]. storage [BLSY22]. stories [HPLQ^{+}21b]. story [BSSV21, Gam25, HB22, San21]. strange [Bie20]. strangers [Jun24]. strategies [RR22b]. Strategy [HS24]. strength [Rat23]. Strichartz [ORT22]. structure [BGT24, Man21]. structured [Foo21]. Structures [Sam20, Par23]. student [Ano20d, Mer20a, Ano21c]. students [D’A23]. study [Dol20]. subject [DH20]. submanifolds [BW24]. subspace [Gre23]. success [BLOS23, DR20]. sufficiency [Lav20]. sufficient [Kas21]. sufficiently [Cal23a]. Sullivan [Ano22c, DS23]. summer [Spr25]. Sums [BSSV21, Rub20]. Sun [GW20]. Sun-Yung [GW20]. Supercentenarian [GN22]. superpower [Mer20e]. supervised [TSAY22]. sure [EMS22, Ele24]. surely [Cal21b]. Surface [ST20, CM21]. surfaces [BKK22, Cal21b, De 24, Dip20, Dol20, Gwy20, HL22, Mar23, Wol21]. survey [Ben22, Lam25, Tie22]. Susan [CGHP^{+}21, Kas23a, vOL25]. Suslin [FM20]. sustainability [Mar20]. swarming [Tad21]. swimming [TGG^{+}25]. Sydney [Kas21]. symmetric [OZ23]. Symmetries [Sea23]. symmetry [Cla24, Sut21, WY23]. symplectic [Cas21, Sch21b]. Symposium [Tuf22]. synaptic [Ude24]. system [KN22, LLM21, Wei20, WGG22]. systems [Ang24, BF21, BLSY22, Ber21, BGH20, Dic20, DR23, Esn23, FLOS24, LST20, Mez21, Müt24, Wan24]. syzygies [Ber20]. Szpiro [EGMD^{+}21].

t [HB24]. table [Pud20]. Tackling [FLU23]. Taft [IMNR23]. Talagrand [DS25]. Talese [Kas21]. talk [Shi25]. talks [Wei20]. Taming [CWA25]. Tate [ABC^{+}21]. taught [Rie23]. Taussky [Goo20]. Taussky-Todd [Goo20]. Teaching [Jun24, LaR24, Bas24a, Edw24, PQY^{+}25, DeL22]. teaching/learning [Bas24a]. technologies [GLY20]. teen [Her22e]. Teichmüller [Wan24, Wol21]. telecommunications [Men20]. tell [MPDL25]. template [DR20]. tenfold [Bae20]. tenth [Shl21]. tenure [AK25, Hau24, Tie22]. Tenured [Kes22]. term [BL24]. Terng [Uhl24]. Testimonios [HPLQ^{+}21b]. testing [BPJ24]. tests [LP24]. their [Has23, Her22d, Her22f, LIJM22, Pol21, SSU23, ZCIM24]. theorem [Ass22, Bae21b, Clo25, Ham25, HL22, MMM20, Ram22]. theorems [Buz20, Dya23, GGT23]. theories [Rat23, Rey25]. theory [Agu24, Ang24, BCC^{+}22, BK25, BE22, BGH20, BSV23, BGT24, CL23, Cla22, De 24, EL25, FLOS24, Far22, Fou20, Hed20, HK21, Lev20, Lop21, Mar24, Nag21, RR21, Rie23, Roj25, ST20, Sch25, Tei20, Val25, XBH^{+}22, vdD22, Ano20c, Ano20e, Ano23c, He22]. things [Ald22]. Thom [CL20]. Thomas [ACD^{+}21, Liu22]. threat [Ful24]. Three [BCF^{+}24, Nat23]. Threshold [Par23]. Thurston [Bon23]. ties [DDV^{+}20]. TikTok [Clo25]. tiles [Zon20]. tiling [Bae24a, Kaw23]. tilings [Tre23]. till [RN23]. time [Dun21c, Mum21, Rut24]. tissues [Foo21]. Tits [Wei23a]. today [LOW23]. Todd [Goo20]. tomography [Yan24]. tool [Her22d, WGG22]. tools [Dic20]. Topological [PA24, FLOS24]. topology [Cas21, HL22, MMP20, Mic23, Pot21, RR24, ST20, Cen22, Lon23]. torch [Hou23a]. toric [Ber20, McF22]. tour [Rey25]. traces [CT21, MMM20]. Tracing [Yan24]. tradition [Lor20]. transformation [Mel21]. transforms [Mer20a]. transitions [Dip20]. Transmission [CCH21]. transport [Cio22, Maw23, ZCIM24]. Transseries [vdD22]. Trapped [HL22]. traveller [Oh22]. Trees [BBC^{+}22b]. treewidth [Epp25]. trends [ARS23, HW25b, Lev20]. triangles [Bae24b]. Triangulating [HP25]. Triangulations [Bae25]. tribute [ACF^{+}24, BBE^{+}25, DAG^{+}23, FS20, GL20, Pie21]. Tricky [Mer22d]. trippy} [Mer22d]. Tropical [Ran23, RTY23, Cha21]. Tumbling [EST24]. Tunnell [BKM23]. Turing [Gam25]. turning [Mil23]. turns [Mer22c]. tutorial [CWA25]. two [Gam25, RRS22]. Tyler [Her21b]. type [Che23, De 23, Maw23].

Ubiratan [Sco22]. Uhlenbeck [DS20b, Mer20b]. Ukraine [VZ23]. Ulf [Ano21j]. Ulmer [Cro21b]. Unbounded [Tre25b]. uncertainty [CWA25]. uncharted [KMMM21]. Uncovering [DNF24]. Undergraduate [Ano21c, Ano20d, LaR24]. undergraduates [Rie23]. understanding [Sch21a]. unfailing [ML24]. unicycle [Ros20a]. Uniformization [MN24]. Union [Ken20a]. Unique [Ber21]. United [RRS22]. unity [Pak22]. Univ [Mak20]. universal [Wan24]. universality [Gan22]. universe [Bie20, Kas21]. universities [VZ23]. University [Tuf22, Ful24, HW25a, HS24, Kas20]. unknown [Ans24]. unlikely [LIJM22]. Untangling [Skr20]. unveiled [Goi21]. Updated [Goi21]. Uraltseva [APS22]. useful [PST23]. Using [LV23, SM24, Ham21, Her21a].

V. [BDF^{+}20]. Vakil [Bee25b]. values

[Hol24]. Varadarajan [BDF^{+}20]. variation [Spo22]. variational [PST23]. varieties [Gup23, McF22, Ort20, RR24]. vascular [LLM21]. Vaughan [BEKP21]. VC [Ten21]. VC-dimension [Ten21]. vehicles [Fie20]. Vera [DMW22]. via [Lan22]. Victor [Pre20]. Victorian [Des21]. viewed [Zha21]. views [BCF^{+}24]. viscoelastic [Sti23, Vas24]. viscosity [Czu24]. visionary [And24, Bai23, CK25]. Visual [Poi22]. Voices [VZ23]. Volumes [Fou20, Ros21]. vs [Lav20]. vulnerable [D’A23].

Wahba [GW22]. Wales [ACF^{+}24]. Walling [JSS23]. Walter [BCW23, SW20]. War [Row22]. Was [Far22]. Washington [Ros20a]. Wave [Cam23]. Wave-mean-flow [Cam23]. wavelets [AEP22]. waves [FMM24]. way [Bae20]. weak [HM21]. wealth [dLLOW23]. weaves [GS24]. website [Goi21]. Weil [Hay21]. Weingarten [CMN22]. Weinstein [BF23]. Weiss [HW23]. West [KMMM21]. Wheat [GM23b]. Wheeden [CFKS22]. Whiteman [Ano21a]. Who [HKL25, GP22]. Widom [BBC^{+}22a]. Wiener [Ano22g, Ass22]. Wigderson [Ano21m, DS22]. wiki [ELW25]. Wikipedia [ELW25]. William [BNS23, Bon23, Ull22]. Williams [DKR22, HN22]. Wilson [BMP24]. winner [vOL25]. Wintenberger [CK20]. Wintner [Ass22]. Wisconsin [GW22]. within [Dum21]. without [Cal22b]. Wolf [Ano20l, Ano22d]. Wolfgang [CKL^{+}23]. women [BCJ^{+}25, BLOS23, Hug24, Kes22, Mur21, Mur23, STHRVG22, Sax20, Tur22]. wonderful [Che22]. word [Hug24, Yon22]. words [Cal22b]. work [DMW22, DR20, HR22, Her22d, Lan22]. Working [Her22f, Her22c]. works [Bon23, BCG22a, CL20]. workshop [BG21a]. World [Row22, ARS23, BCK24, Che22, CWA25, Her22g, Lau20b, Tho20]. worlds [Dip20, Oh22]. writing [Kha25, SM24].

Yamabe [MN24]. Yau [Sun22]. year [BFH^{+}21, Her22a, ILM^{+}22]. years [Dol20, MMS25, Wes23, Mar24]. York [Ald22, Kas21, Nat23]. YouTube [Clo25, Her22g]. Yung [GW20].

Zelditch [Ano23d]. zero [CP22, FHLS20, HB24]. zeta [Hol24]. Zhang [Gan22]. Zimmer [Fis20]. Zucker [CHJ21].

References

Athreya:2025:CHC

[AAD^{+}25] Jayadev Athreya, Padi Fuster Aguilera, Matthew Gentry Durham, Seppo Niemi Colvin, Florencia Orosz Hunziker, Daniel Reinholz, Keri Ann Sather-Wagstaff, Noelle Sawyer, Chad M. Topaz, and Ila Varma. A call for hybrid conferences in mathematics [Letter to the editor]. Notices of the American Mathematical Society, 72(2): 132–133, February 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With a response by John Meier and Michael Pearson.

Andrews:2022:CRR

[ABB^{+}22] George E. Andrews, G. J.

Babu, Zhidong Bai, David Banks, Herman Chernoff, Roy Frieden, Yasunori Fujikoshi, Kanti V. Mardia, Basilio de Bragança Pereira, Urmila Pingle, M. B. Rao, Arni S. R. Srinivasa Rao, James Rosenberger, S. R. S. Varadhan, and Yuehua Wu. C. R. Rao (1920– ) celebrates his 101st birthday. Notices of the American Mathematical Society, 69(6):982–997, June/July 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Axler:2023:RCD

[ABB^{+}23] Sheldon Axler, John J. Benedetto, Rajendra Bhatia, Man-Duen Choi, Aaron Davis, Natalie Zemon Davis, Simone Weil Davis, John Friedlander, Mary Gray, Stephen Halperin, Peter Rosenthal, Marjorie Senechal, and Hannah Davis Taïeb. Remembrances of Chandler Davis (1926–2022). Notices of the American Mathematical Society, 70(7):1107–1119, August 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Agranovsky:2020:RCB

[ABC^{+}20] Mark Agranovsky, John Baras, Der-Chen Chang, Roger Gay, Peter Kuchment, Bao Qin Li, Domenico Napoletani, Irene Sabadini, Ahmed Sebbar, Daniele C. Struppa, and Alain Yger. Remembering Carlos A. Berenstein (1944–2019). Notices of the American Mathematical Society, 67(8):1146– 1157, September 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Acquista:2021:MAJ

[ABC^{+}21] Karen Acquista, Joe Buhler, Dustin Clausen, John H. Coates, Benedict H. Gross, Stephen Lichtenbaum, Jonathan Lubin, Barry Mazur, V. Kumar Murty, Bernadette Perrin-Riou, Cristian D. Popescu, Kenneth A. Ribet, Joseph H. Silverman, Karen Uhlenbeck, and José Felipe Voloch. Memorial article for John Tate. Notices of the American Mathematical Society, 68(5):768–781, May 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anderson:2025:HAS

[ABH^{+}25] Lara B. Anderson, Steve Bradlow, Yang-Hui He, Laura P. Schaposnik, and Maximilian J. Telford. How AI is shaping everyday research. Notices of the American Mathematical Society, 72(5):529–535, May 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Aguilar:2021:IC

[ABLL21] Konrad Aguilar, Samantha Brooker, Frédéric Latrémolière, and Alejandra López. An ideal convergence. Notices of the American Mathematical Society, 68(8):1269–1281, September 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Alvarez:2021:CIP

[ÁC21] James A. Mendoza Álvarez and Minerva Cordero. Considerations for increasing participation of minoritized ethnic and racial groups in mathematics. Notices of the American Mathematical Society, 68(2):235–239, February 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Aldous:2021:LML

[ACD^{+}21] David Aldous, Pietro Caputo, Rick Durrett, Alexander E. Holroyd, Paul Jung, and Amber L. Puha. The life and mathematical legacy of Thomas M. Liggett. Notices of the American Mathematical Society, 68 (1):67–79, January 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Aschbacher:2024:MTD

[ACF^{+}24] Michael Aschbacher, Arjeh M. Cohen, Richard Foote, Robert Guralnick, Philip Hanlon, and Claire Levaillant. A memorial tribute to David B. Wales. Notices of the American Mathematical Society, 71(7):909–915, August 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Abramovich:2024:DAP

[ACFH24] Dan Abramovich, Henry Cohn, David Futer, and Robert Harington. Double-anonymous peer review in mathematics: implementation for American Mathematical Society journals. Notices of the American Mathematical Society, 71(8):1079– 1081, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Avigad:2025:AFP

[ACMT25] Jeremy Avigad, Johan Commelin, Heather Macbeth, and Adam Topaz. Anatomy of a formal proof. Notices of the American Mathematical Society, 72(2):188–195, February 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Aizenman:2024:MMB

[ADG^{+}24] Michael Aizenman, Ingrid Daubechies, Mary Gray, Cathy Kessel, Harriet Pollatsek, Graeme Smith, Elisabeth Werner, and Bei Zeng. In memory of Mary Beth Ruskai. Notices of the American Mathematical Society, 71(10):1362–1371, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Austin:2022:IDP

[AEP22] David Austin, Malena I.

Español, and Mirjeta Pasha. The image deblurring problem: matrices, wavelets, and multilevel methods. Notices of the American Mathematical Society, 69(8):1284–1295, September 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Araujo:2022:FCM

[AF22] Carolina Araujo and João Paulo Figueredo. Foliations on complex manifolds. Notices of the American Mathematical Society, 69(7):1137–1144, August 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Andrews:2021:MFD

[AFS21] George E. Andrews, Jürg Fröhlich, and Andrew V. Sills. In memoriam: Freeman Dyson (1923–2020). Notices of the American Mathematical Society, 68(7):1140–1155, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Aguilar:2024:OTA

[Agu24] Konrad Aguilar. Operator theory and analysis of infinite networks: theory and applications [book review of 4641241]. Notices of the American Mathematical Society, 71(10):1381– 1383, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Aksoy:2022:MMS

[AHJ^{+}22] Sinan G. Aksoy, Aric Hagberg, Cliff A. Joslyn, Bill Kay, Emilie Purvine, and Stephen J. Young. Models and methods for sparse (hyper)network science in business, industry, and government. Notices of the American Mathematical Society, 69(2):287–291, February 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Adem:2023:RGB

[AHRZ23] Alejandro Adem, Tom Halverson, Arun Ram, and Efim Zelmanov. Remembering Georgia Benkart. Notices of the American Mathematical Society, 70 (3):420–433, March 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Alm:2025:NTP

[AK25] Jeremy F. Alm and Kevin P.

Knudson. Navigating the tenure process in a mathematics department. Notices of the American Mathematical Society, 72 (5):507–513, May 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Aldous:2022:RCW

[Ald22] David Aldous. The rules of contagion: why things spread — and why they stop [book review, Basic Books, New York, NY, 2020]. Notices of the American Mathematical Society, 69 (6):1008–1011, June/July 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Alladi:2020:IRR

[All20]  Krishnaswami Alladi.  An

invitation to the Rogers– Ramanujan identities [book review of 3752624]. Notices of the American Mathematical Society, 67(1):68–71, January 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Abebe:2021:OJS

[ALM21] Ash Abebe, Suzanne Lenhart, and Brittany McCullough. Overtoun Jenda: a STEM mentor extraordinaire. Notices of the American Mathematical Society, 68(2):258–261, February 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ardila-Mantilla:2020:GRS

[AM20] Federico Ardila-Mantilla. CAT(0) geometry, robots, and society. Notices of the American Mathematical Society, 67(7):977–987, August 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anderson:2024:FHA

[And24] Christian Anderson. Freeman Hrabowski: advocate for mathematics and STEM visionary. Notices of the American Mathematical Society, 71(8):1056– 1061, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Angiono:2024:RSL

[Ang24] Iván Angiono. Root systems in Lie theory: from the classic definition to nowadays. Notices of the American Mathematical Society, 71(5):605–612, May 2024. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:ADP

[Ano20a] Anonymous. 2020 Award for Distinguished Public Service. Notices of the American Mathematical Society, 67(4):575–576, April 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:BMP

[Ano20b]  Anonymous.  2020 Bôcher

Memorial Prize. Notices of the American Mathematical Society, 67(4):546–549, April 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:CPL

[Ano20c] Anonymous. 2020 Chevalley Prize in Lie Theory. Notices of the American Mathematical Society, 67(4):554–556, April 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:FBM

[Ano20d] Anonymous. 2020 Frank and Brennie Morgan Prize for Outstanding Research in Mathematics by an undergraduate student. Notices of the American Mathematical Society, 67 (4):569–572, April 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:FNC

[Ano20e] Anonymous. 2020 Frank Nelson Cole Prize in Number Theory. Notices of the American Mathematical Society, 67 (4):550–551, April 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:JPB

[Ano20f] Anonymous. 2020 Joint Policy Board for Mathematics Communications Award. Notices of the American Mathematical Society, 67(4):573–574, April 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:JDP

[Ano20g] Anonymous. 2020 Joseph L. Doob Prize. Notices of the American Mathematical Society, 67(4):557–559, April 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:LEP

[Ano20h] Anonymous. 2020 Leonard Eisenbud Prize for Mathematics and Physics. Notices of the American Mathematical Society, 67(4):560–562, April 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:LPS

[Ano20i] Anonymous. 2020 Leroy P. Steele Prizes. Notices of the American Mathematical Society, 67(4):563–567, April 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:LCP

[Ano20j] Anonymous. 2020 Levi L. Conant Prize. Notices of the American Mathematical Society, 67(4):552–553, April 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:BKA

[Ano20k] Anonymous. Beilinson and Kazhdan awarded 2020 Shaw Prize. Notices of the American Mathematical Society, 67 (8):1252–1253, September 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:DEA

[Ano20l] Anonymous. Donaldson and Eliashberg awarded 2020 Wolf Prize. Notices of the American Mathematical Society, 67(6): 912–913, June/July 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:FMA

[Ano20m] Anonymous. Furstenberg and Margulis awarded 2020 Abel Prize. Notices of the American Mathematical Society, 67(6): 908–911, June/July 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:JKE

[Ano20n] Anonymous. Jane Kister: executive editor extraordinaire. Notices of the American Mathematical Society, 67(6):877–881, June/July 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2020:MP

[Ano20o] Anonymous. Mathematics people. Notices of the American Mathematical Society, 67 (4):580–583, April 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:ALW

[Ano21a] Anonymous. 2021 Albert Leon Whiteman Memorial Prize. Notices of the American Mathematical Society, 68(4):634–635, April 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:BRP

[Ano21b] Anonymous. 2021 Bertrand Russell Prize. Notices of the American Mathematical Society, 68(4):624–625, April 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:FBM

[Ano21c] Anonymous. 2021 Frank and Brennie Morgan Prize for Outstanding Research in Mathematics by an Undergraduate Student. Notices of the American Mathematical Society, 68 (4):621–623, April 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:FNC

[Ano21d] Anonymous. 2021 Frank Nelson Cole Prize in Algebra. Notices of the American Mathematical Society, 68(4):612–613, April 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:GDB

[Ano21e]  Anonymous.  2021 George

David Birkhoff Prize. Notices of the American Mathematical Society, 68(4):610–611, April 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:LPS

[Ano21f] Anonymous. 2021 Leroy P.

Steele Prizes. Notices of the American Mathematical Society, 68(4):629–633, April 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:LCP

[Ano21g] Anonymous. 2021 Levi L. Conant Prize. Notices of the American Mathematical Society, 68(4):614–615, April 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:MPD

[Ano21h] Anonymous. 2021 Mary P. Dolciani Prize. Notices of the American Mathematical Society, 68(4):616–617, April 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:RLS

[Ano21i] Anonymous. 2021 Ruth Lyttle Satter Prize. Notices of the American Mathematical Society, 68(4):626–627, April 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:UGP

[Ano21j] Anonymous. 2021 Ulf Grenander Prize. Notices of the American Mathematical Society, 68 (4):618–619, April 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:BCA

[Ano21k]  Anonymous.  Bismut and

Cheeger awarded Shaw Prize. Notices of the American Mathematical Society, 68(8):1402– 1403, September 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:LN

[Ano21l] Anonymous. Louis Nirenberg (1925–2020). Notices of the American Mathematical Society, 68(6):959–979, June/July 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Coordinated by Robert V. Kohn and Yanyan Li.

Anonymous:2021:LWA

[Ano21m]  Anonymous.  Lovász and

Wigderson awarded 2021 Abel Prize. Notices of the American Mathematical Society, 68 (6):1034–1037, June/July 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:MAS

[Ano21n] Anonymous. Mikhail Aleksandrovich Shubin (1944–2020). Notices of the American Mathematical Society, 68(1):81–93, January 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2021:RJB

[Ano21o]  Anonymous.  Remembering

Jean Bourgain (1954–2018). Notices of the American Mathematical Society, 68(6):942–957, June/July 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2022:FMA

[Ano22a]  Anonymous.  2022 Fields

Medals awarded. Notices of the American Mathematical Society, 69(9):1601–1603, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2022:AHA

[Ano22b] Anonymous. Alon and Hrushovski awarded 2022 Shaw Prize. Notices of the American Mathematical Society, 69(8):1432– 1433, September 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2022:DSA

[Ano22c] Anonymous. Dennis Sullivan awarded 2022 Abel Prize. Notices of the American Mathematical Society, 69(6):1065– 1067, June/July 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2022:GLA

[Ano22d] Anonymous. George Lusztig awarded 2022 Wolf Prize. Notices of the American Mathematical Society, 69(6):1068– 1069, June/July 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2022:LPS

[Ano22e] Anonymous. Leroy P. Steele Prizes. Notices of the American Mathematical Society, 69 (4):663–667, April 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2022:LCP

[Ano22f] Anonymous. Levi L. Conant Prize. Notices of the American Mathematical Society, 69 (4):656–657, April 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2022:NWP

[Ano22g] Anonymous. Norbert Wiener Prize in applied mathematics. Notices of the American Mathematical Society, 69(4):637–638, April 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2023:BMP

[Ano23a] Anonymous. Bôcher Memorial Prize. Notices of the American Mathematical Society, 70 (4):664–666, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2023:FBM

[Ano23b] Anonymous. Frank and Brennie Morgan Prize. Notices of the American Mathematical Society, 70(4):659–661, April 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2023:FNC

[Ano23c] Anonymous. Frank Nelson Cole Prize in Number Theory. Notices of the American Mathematical Society, 70(4):667–669, April 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2023:MSZ

[Ano23d] Anonymous. In memory of Steve Zelditch. Notices of the American Mathematical Society, 70(10):1667–1683, November 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Coordinated by Bernard Shiffman and Jared Wunsch.

Anonymous:2023:JDP

[Ano23e] Anonymous. Joseph L. Doob Prize. Notices of the American Mathematical Society, 70 (4):674–675, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2023:LEP

[Ano23f] Anonymous. Leonard Eisenbud Prize for mathematics and physics. Notices of the American Mathematical Society, 70 (4):676–678, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2023:LPS

[Ano23g] Anonymous. Leroy P. Steele Prizes. Notices of the American Mathematical Society, 70 (4):687–689, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2023:LCP

[Ano23h] Anonymous. Levi L. Conant Prize. Notices of the American Mathematical Society, 70 (4):670–671, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2023:MPD

[Ano23i] Anonymous. Mary P. Dolciani Prize for excellence in research. Notices of the American Mathematical Society, 70 (4):672–673, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2023:RLS

[Ano23j] Anonymous. Ruth Lyttle Satter Prize in mathematics. Notices of the American Mathematical Society, 70(4):685–686, April 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2024:CFA

[Ano24a] Anonymous. 2024 class of Fellows of the AMS. Notices of the American Mathematical Society, 71(2):258–259, February 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2024:JJK

[Ano24b] Anonymous. Joseph J. Kohn (1932–2023). Notices of the American Mathematical Society, 71(6):773–785, June/July 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2025:P

[Ano25a] Anonymous. 2025 Prizes. Notices of the American Mathematical Society, 72(4):437–458, April 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2025:JS

[Ano25b] Anonymous. Jim Simons (1938– 2024). Notices of the American Mathematical Society, 72(1):32– 47, January 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions.

Anscombe:2024:ULP

[Ans24] Sylvy Anscombe. The unknown logic of power series. Notices of the American Mathematical Society, 71(11):1448– 1455, 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Apushkinskaya:2022:NNU

[APS22] Darya Apushkinskaya, Arshak Petrosyan, and Henrik Shahgholian. Nina Nikolaevna Uraltseva. Notices of the American Mathematical Society, 69 (3):385–395, March 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arroyo-Rabasa:2023:MMW

[ARS23] Adolfo Arroyo-Rabasa and José Simental. Mexican mathematicians in the world: trends and recent contributions [book review of 4344303]. Notices of the American Mathematical Society, 70(8):1274–1277, September 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Assani:2022:WWE

[Ass22] Idris Assani. Wiener–Wintner ergodic theorem, in brief. Notices of the American Mathematical Society, 69(2):198–209, February 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Awanou:2021:CNG

[Awa21] Gerard Awanou. Computational nonimaging geometric optics: Monge–Ampére. Notices of the American Mathematical Society, 68(2):186–193, February 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Babecki:2022:WLG

[Bab22] Catherine Babecki. What is … a graphical design? Notices of the American Mathematical Society, 69(9):1571–1573, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2020:TW

[Bae20] John C. Baez. The tenfold way.

Notices of the American Mathematical Society, 67(10):1599– 1601, November 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2021:BM

[Bae21a] John C. Baez. The Brownian map. Notices of the American Mathematical Society, 68 (5):801–803, May 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2021:GLT

[Bae21b] John C. Baez. The Gauss– Lucas theorem. Notices of the American Mathematical Society, 68(11):1988–1989, December 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2023:I

[Bae23a] John C. Baez. The icosidodecahedron. Notices of the American Mathematical Society, 70 (5):821–823, May 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2023:ID

[Bae23b] John C. Baez. Isbell duality. Notices of the American Mathematical Society, 70(1):140–141, January 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2024:HTH

[Bae24a] John C. Baez. The hexagonal tiling honeycomb. Notices of the American Mathematical Society, 71(10):1390–1391, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2024:MSA

[Bae24b] John C. Baez. The moduli space of acute triangles. Notices of the American Mathematical Society, 71(5):664–665, May 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2025:TS

[Bae25] John C. Baez. Triangulations of the sphere. Notices of the American Mathematical Society, 72 (4):430–431, April 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bailey:2023:PBV

[Bai23] David H. Bailey. Peter Borwein:

a visionary mathematician. Notices of the American Mathematical Society, 70(4):610–613, April 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

https://community.ams.org/
journals/notices/202304/rnoti-
p610.pdf.

Bamler:2021:RDR

[Bam21] Richard H. Bamler. Recent developments in Ricci flows. Notices of the American Mathematical Society, 68(9):1486– 1498, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barber:2024:DAN

[Bar24a] Janet Barber. Diversity of authors in Notices and beyond.

Notices of the American Mathematical Society, 71(9):1203– 1209, September 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bartlett:2024:QIE

[Bar24b] Bruce Bartlett. The quintic, the icosahedron, and elliptic curves. Notices of the American Mathematical Society, 71 (4):444–453, April 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bass:2024:DLB

[Bas24a] Hyman Bass. Deborah Loewenberg Ball: teaching/learning mathematics teaching. Notices of the American Mathematical Society, 71(3):302–311, March 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Basu:2024:MFR

[Bas24b] Saugata Basu. Marie-Françoise Roy. Notices of the American Mathematical Society, 71 (3):313–325, March 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bonato:2021:SAR

[BBB21] Anthony Bonato, Juliette Bruce, and Ron Buckmire. Spaces for all: the rise of LGBTQ+ mathematics conferences. Notices of the American Mathematical Society, 68(6):

998–1003, June/July 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bayer:2024:MPB

[BBB^{+}24a] Jonas Bayer, Christoph Benzmüller, Kevin Buzzard, Marco David, Leslie Lamport, Yuri Matiyasevich, Lawrence Paulson, Dierk Schleicher, Benedikt Stock, and Efim Zelmanov. Mathematical proof between generations. Notices of the American Mathematical Society, 71(1):79–92, January 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Buckmire:2024:MSM

[BBB^{+}24b] Ron Buckmire, Barbara Beeton, Robert Bryant, Fernando Q. Gouvêa, Anthony V. Phillips, Dennis Sullivan, and Michael Wolf. Michael Spivak: a memorial. Notices of the American Mathematical Society, 71(6): 786–795, June/July 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Basor:2022:RHW

[BBC^{+}22a] Estelle Basor, Albrecht Böttcher, Ivan Corwin, Persi Diaconis, Torsten Ehrhardt, Al Kelley, Barry Simon, Craig A. Tracy, and Tony Tromba. Remembrances of Harold Widom. Notices of the American Mathematical Society, 69(4):586–598, April 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blake:2022:TMC

[BBC^{+}22b] Heather Smith Blake, Miklós Bóna, Éva Czabarka, Stephan Wagner, and Hua Wang. Trees in many contexts. Notices of the American Mathematical Society, 69(2):284–286, February 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berger:2025:THP

[BBE^{+}25] Marsha Berger, Björn Birnir, Nick Ercolani, Jonathan Goodman, F. Alberto Grünbaum, EnriKo Loubet, Victor H. Moll, David Mumford, Brian Rider, Jalal Shatah, Dan Stroock, Eugene Trubowitz, and Pierre van Moerbeke. A tribute to Henry P. McKean Jr. Notices of the American Mathematical Society, 72(3):275–283, March 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blasius:2020:MGS

[BBH^{+}20] Don Blasius, Toni Bluher, Haruzo Hida, Kamal Khuri-Makdisi, Kenneth Ribet, Alice Silverberg, and Hiroyuki Yoshida. Memories of Goro Shimura. Notices of the American Mathematical Society, 67 (5):677–689, May 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Buhler:2021:RLG

[BBS21] Joe Buhler, Steve Butler, and Joel Spencer. Ronald Lewis Graham (1935–2020). Notices of the American Mathematical Society, 68(11):1931–1950, December 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2022:ACT

[BCC^{+}22] John Baez, Simon Cho, Daniel Cicala, Nina Otter, and Valeria de Paiva. Applied category theory in chemistry, computing, and social networks. Notices of the American Mathematical Society, 69(2):292–297, February 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2023:BR

[BCD23] John C. Baez, J. Daniel Christensen, and Sam Derbyshire. The beauty of roots. Notices of the American Mathematical Society, 70(9):1495–1497, October 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brown:2021:MPD

[BCF^{+}21] Johnny Brown, Martin Chuaqui, Tim Ferguson, Dima Khavinson, Steven G. Krantz, Brad Osgood, Alex Schuster, and Rachel Weir. In memory of Peter Duren. Notices of the American Mathematical Society, 68(8):1324–1335, September 2021. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Boaler:2024:TVC

[BCF^{+}24] Jo Boaler, Brian Conrad, Ben Ford, Rafe Mazzeo, and Jelani Nelson. Three views on the California math framework. Notices of the American Mathematical Society, 71(6):797–805, June/July 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bryant:2022:IMSa

[BCG22a] Robert Bryant, Jeff Cheeger, and Phillip Griffiths. Isadore M. Singer (1924–2021) in memoriam Part 1: scientific works. Notices of the American Mathematical Society, 69(9):1547– 1563, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bryant:2022:IMSb

[BCG22b] Robert Bryant, Jeff Cheeger, and Phillip Griffiths. Isadore M. Singer (1924–2021) in memoriam Part 2: Personal recollections. Notices of the American Mathematical Society, 69(10):1770–1781, November 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bonk:2020:AMS

[BCH^{+}20] Mario Bonk, Luca Capogna, Piotr Hajłasz, Nageswari Shanmugalingam, and Jeremy Tyson. Analysis in metric spaces. Notices of the American Mathematical Society, 67(2):253–256, February 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bauer:2025:RNW

[BCJ^{+}25] Kristine Bauer, Erin Chambers, Brenda Johnson, Kristin Lauter, Kathryn Leonard, and Michelle Snider. Research networks for women. Notices of the American Mathematical Society, 72(4):393–403, April 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bowman:2024:FMM

[BCK24] Joshua Bowman, Katie Chicot, and Mustafa Kayamaz. Four math museums around the world: from oldest to newest. Notices of the American Mathematical Society, 71(9):1210– 1217, September 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bass:2025:IHB

[BCL25] Hyman Bass, Lisa Carbone, and Yvonne Lai. Interview with Hyman Bass. Notices of the American Mathematical Society, 72(6):653–665, June/ July 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bercovici:2022:RCI

[BCT^{+}22] Hari Bercovici, Peter Constantin, Allen Tannenbaum, Roger Temam, and Edriss S.

Titi. Remembrances of Ciprian Ilie Foias. Notices of the American Mathematical Society, 69 (9):1529–1545, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bakas:2023:WRM

[BCW23] Odysseas Bakas, Valentina Ciccone, and James Wright. Walter Rudin meets Elias M. Stein. Notices of the American Mathematical Society, 70(10):1697– 1703, November 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bratteli:2020:OBH

[BDE^{+}20] Tone Bratteli, Trond Digernes, George A. Elliott, David E. Evans, Palle E. T. Jorgensen, Aki Kishimoto, Magnus B. Landstad, Derek W. Robinson, and Erling Størmer. Ola Bratteli and his diagrams. Notices of the American Mathematical Society, 67(5):665–675, May 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blasius:2020:RVV

[BDF^{+}20] Don Blasius, Trond Digernes, Rita Fioresi, Ramesh Gangolli, Michael Rapoport, and S. R. S. Varadhan. Recollections of V. S. Varadarajan. Notices of the American Mathematical Society, 67(9):1365–1373, October 2020. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Barbero:2023:PCD

[BdMG23] Sergio Barbero and María del Mar González. Presbyopia correction, differential geometry, and free boundary PDEs. Notices of the American Mathematical Society, 70(11):1763– 1771, December 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Breiner:2020:CPS

[BDS20] Spencer Breiner, Peter Denno, and Eswaran Subrahmanian. Categories for planning and scheduling. Notices of the American Mathematical Society, 67(11):1666–1677, December 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bendory:2022:ATP

[BE22] Tamir Bendory and Dan Edidin. Algebraic theory of phase retrieval. Notices of the American Mathematical Society, 69(9):1487–1495, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brendle:2024:II

[BE24] Simon Brendle and Michael Eichmair. The isoperimetric inequality. Notices of the American Mathematical Society, 71(6):708–713, June/July 2024. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Beck:2020:SSS

[Bec20] Margaret Beck. Spectral stability and spatial dynamics in partial differential equations. Notices of the American Mathematical Society, 67(4):500–507, April 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beebe:2024:HAM

[Bee24] Elaine Beebe. Honoring Aziz: memorial session held at Spring Eastern Sectional Meeting. Notices of the American Mathematical Society, 71(7):964–965, August 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beebe:2025:IBK

[Bee25a] Elaine Beebe. An interview with Bryna Kra. Notices of the American Mathematical Society, 72(4):459–460, April 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beebe:2025:IRV

[Bee25b] Elaine Beebe. Interview with Ravi Vakil. Notices of the American Mathematical Society, 72(2):214–215, February 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bisch:2021:MVJ

[BEKP21] Dietmar Bisch, David E. Evans, Robion Kirby, and Sorin Popa. Memories of Vaughan Jones. Notices of the American Mathematical Society, 68(9):1540– 1563, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Benedetto:2022:SNA

[Ben22] Robert L. Benedetto. A survey of non-Archimedean dynamics. Notices of the American Mathematical Society, 69(5):715–723, May 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Braverman:2024:MIK

[BEO^{+}24] Alexander Braverman, Pavel Etingof, Andrei Okounkov, Duong Phong, and Paul Wiegmann. In memory of Igor Krichever. Notices of the American Mathematical Society, 71 (4):483–499, April 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berkesch:2020:GTS

[Ber20] Christine Berkesch. The geometry of toric syzygies. Notices of the American Mathematical Society, 67(4):537, April 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berhanu:2021:UCF

[Ber21] Shiferaw Berhanu. Unique continuation for first order systems of PDEs. Notices of the American Mathematical Society, 68(9):1479–1485, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bernshteyn:2022:DCD

[Ber22] Anton Bernshteyn. Descriptive combinatorics and distributed algorithms. Notices of the American Mathematical Society, 69(9):1496–1507, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2021:ODS

[BF21] John C. Baez and John D. Foley. Operads for designing systems of systems. Notices of the American Mathematical Society, 68(6):1005–1007, June/ July 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bursztyn:2023:CAW

[BF23] Henrique Bursztyn and Rui Loja Fernandes. A conversation with Alan Weinstein. Notices of the American Mathematical Society, 70(1):125–133, January 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bennett:2021:RMA

[BFH^{+}21] Curtis Bennett, Robin Forman, Phil Hanlon, Herbert A. Medina, Sastry G. Pantula, Wayne Raskind, and Rachel Vincent-Finley. Reflections of mathematician administrators in the year of the plague. Notices of the American Mathematical Society, 68(8):1357–1366, September 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barcelo:2021:MAA

[BG21a] Hélène Barcelo and Edray Herber Goins. MSRI’s ADJOINT: African diaspora joint mathematics workshop. Notices of the American Mathematical Society, 68(2):246–253, February 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bowen:2021:PDP

[BG21b] Claire McKay Bowen and Simson Garfinkel. The philosophy of differential privacy. Notices of the American Mathematical Society, 68(10):1727– 1739, November 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berry:2020:BDS

[BGH20] Tyrus Berry, Dimitrios Giannakis, and John Harlim. Bridging data science and dynamical systems theory. Notices of the American Mathematical Society, 67(9):1336–1349, October 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bloom:2023:RIB

[BGH^{+}23] Walter R. Bloom, Richard J. Gardner, Al Hales, Joel Spencer, Terence Tao, and Benjamin Weiss. Robert Israel “Bob” Jewett (1937–2022). Notices of the American Mathematical Society, 70(5):772–781, May 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Beauville:2022:RMN

[BGPH^{+}22] Arnaud Beauville, Oscar García-Prada, Nigel Hitchin, Chandrashekhar Khare, Shrawan Kumar, Herbert Lange, Nitin Nitsure, M. S. Raghunathan, T. R. Ramadas, and S. Ramanan. Remembering M. S. Narasimhan (1932–2021). Notices of the American Mathematical Society, 69(7):1189– 1203, August 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bradley:2024:SML

[BGT24] Tai-Danae Bradley, Juan Luis Gastaldi, and John Terilla. The structure of meaning in language: parallel narratives in linear algebra and category theory. Notices of the American Mathematical Society, 71(2):174–185, February 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Balestriero:2025:GDL

[BHB25] Randall Balestriero, Ahmed Imtiaz Humayun, and Richard G.

Baraniuk. On the geometry of deep learning. Notices of the American Mathematical Society, 72(4):374–385, April 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Balaji:2021:CS

[BHL^{+}21] Vikraman Balaji, William Haboush, Venkatramani Lakshmibai, Peter Littelmann, and David Mumford. C. S. Seshadri (1932–2020). Notices of the American Mathematical Society, 68(11):1951–1967, December 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blumberg:2024:HC

[BHO^{+}24] Andrew J. Blumberg, Michael A.

Hill, Kyle Ormsby, Angélica M. Osorno, and Constanze Roitzheim. Homotopical combinatorics. Notices of the American Mathematical Society, 71(2):260–266, February 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Baez:2022:KSE

[BHW22] John C. Baez, Steve Huntsman, and Cheyne Weis. The Kuramoto–Sivashinky equation. Notices of the American Mathematical Society, 69(9): 1581–1583, October 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bieri:2020:LMS

[Bie20] Lydia Bieri. A lady mathematician in this strange universe: memoirs [book review of 3753699]. Notices of the American Mathematical Society, 67 (3):384–389, March 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berthelot:2020:JMF

[BIK^{+}20] Pierre Berthelot, Luc Illusie, Nicholas M. Katz, William Messing, and Peter Scholze. Jean-Marc Fontaine (1944– 2019). Notices of the American Mathematical Society, 67(7): 1010–1027, August 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bimbo:2023:CHM

[Bim23] Katalin Bimbó. Concepts at the heart of mathematics — through the centuries [book review of 4529888]. Notices of the American Mathematical Society, 70(10):1690–1693, November 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Belkale:2025:EAR

[BK25] Prakash Belkale and Joshua Kiers. Extremal aspects of representation theory. Notices of the American Mathematical Society, 72(4):365–373, April 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Breiner:2022:CLG

[BKK22] Christine Breiner, Nikolaos Kapouleas, and Stephen Kleene. Conservation laws and gluing constructions for constant mean curvature (hyper)surfaces. Notices of the American Mathematical Society, 69(5):762–773, May 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Buhler:2023:JT

[BKM23] Joe Buhler, Alex Kontorovich, and Stephen D. Miller. Jerry Tunnell (1950–2022). Notices of the American Mathematical Society, 70(6):945–952, June/ July 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bos:2021:MPM

[BL21] Joppe W. Bos and Kristin E.

Lauter. In memoriam: Peter L. Montgomery (1947–2020). Notices of the American Mathematical Society, 68(4):538–545, April 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Birgin:2023:SOC

[BL23] Ernesto G. Birgin and Antoine Laurain. Shape optimization for covering problems. Notices of the American Mathematical Society, 70(10):1584– 1591, November 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Boucard:2024:CGQ

[BL24] Jenny Boucard and Jemma Lorenat. Catherine Goldstein: the question of long-term histories and mathematical identity. Notices of the American Mathematical Society, 71(3):338–345, March 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bozeman:2023:SCM

[BLOS23] Sylvia T. Bozeman, Emille Davie Lawrence, Yewande Olubummo, and Monica Stephens. Spelman College: a model of success for producing black women in mathematics. Notices of the American Mathematical Society, 70(2):301–308, February 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bar-Lev:2022:ECP

[BLSY22] Daniella Bar-Lev, Omer Sabary, and Eitan Yaakobi. Exciting coding problems for DNAbased storage systems. Notices of the American Mathematical Society, 69(11):1987–1991, December 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Benkart:2021:AB

[BLW21] Georgia Benkart, Kristin Lauter, and Sylvia Wiegand. AWM at 50 and beyond. Notices of the American Mathematical Society, 68(3):387–397, March 2021. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Braden:2024:DWC

[BMP24] Tom Braden, Jacob P. Matherne, and Nicholas Proudfoot. the Dowling–Wilson Conjecture? Notices of the American Mathematical Society, 71 (9):1197–1201, September 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barnsley:2023:RDW

[BNS23] Michael Barnsley, Bruno Nachtergaele, and Barry Simon. Remembrances of Derek William Robinson, June 25, 1935– August 31, 2021. Notices of the American Mathematical Society, 70(8):1252–1267, September 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions by Alain Connes, David Evans, Giovanni Gallavotti, Sheldon Glashow, Arthur Jaffe, Palle Jorgensen, Aki Kishimoto, Elliott Lieb, Heide Narnhofer, David Ruelle, Mary Beth Ruskai, Adam Sikora, and A. F. M. ter Elst.

Bonato:2022:IJC

[Bon22a] Anthony Bonato. Interview with Jennifer Chayes. Notices of the American Mathematical Society, 69(3):403–407, March 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bonato:2022:ILJ

[Bon22b] Anthony Bonato. Interview with Lisa Jeffrey. Notices of the American Mathematical Society, 69(3):410–412, March 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bonato:2022:IMC

[Bon22c] Anthony Bonato. Interview with Maria Chudnovsky. Notices of the American Mathematical Society, 69(3):407–410, March 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bonahon:2023:CWW

[Bon23] Francis Bonahon. The collected works of William P. Thurston with commentary [book review of 4554427; 4556463; 4556600; 4554426]. Notices of the American Mathematical Society, 70 (4):626–629, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bettiol:2020:IB

[BP20] Renato G. Bettiol and Paolo Piccione. Instability and bifurcation. Notices of the American Mathematical Society, 67(11):1679–1691, December 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bradley:2025:DD

[BP25] Robert E. Bradley and Salvatore J. Petrilli, Jr. From differentials to derivatives. Notices of the American Mathematical Society, 72(6):624–635, June/July 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bah:2024:CSB

[BPJ24] Bubacarr Bah, Hendrik Bernd Petersen, and Peter Jung. Compressed sensing-based SARS-CoV-2 pool testing. Notices of the American Mathematical Society, 71(2):158–165, February 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blekherman:2021:SSR

[BSSV21] Grigoriy Blekherman, Rainer Sinn, Gregory G. Smith, and Mauricio Velasco. Sums of squares: a real projective story. Notices of the American Mathematical Society, 68(5):734–747, May 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Breiding:2020:CS

[BST20] Paul Breiding, Bernd Sturmfels, and Sascha Timme. 3264 conics in a second. Notices of the American Mathematical Society, 67(1):30–37, January 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blum-Smith:2023:MLI

[BSV23] Ben Blum-Smith and Soledad Villar. Machine learning and invariant theory. Notices of the American Mathematical Society, 70(8):1205–1213, September 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Butler:2020:MLF

[But20] Steve Butler. The mathematical life of Fan Chung. Notices of the American Mathematical Society, 67(3):327–335, March 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Buzzard:2020:PTC

[Buz20] Kevin Buzzard. Proving theorems with computers. Notices of the American Mathematical Society, 67(11):1791–1799, December 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bernstein:2024:CME

[BW24] Jacob Bernstein and Lu Wang. Colding–Minicozzi entropy and complexity of submanifolds. Notices of the American Mathematical Society, 71(7):861–867, August 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bennett:2024:MAW

[BWW24] Hanna Bennett, Nina White, and Beth Wolf. Mastery assessment: what, why, and how? Notices of the American Mathematical Society, 71(10):1344– 1347, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cadoret:2021:WC

[Cad21] Anna Cadoret. What is … a companion? Notices of the American Mathematical Society, 68(7):1179–1182, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2020:ASD

[Cal20a] Danny Calegari. Angenent’s shrinking doughnuts. Notices of the American Mathematical Society, 67(3):412–414, March 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2020:KFL

[Cal20b] Danny Calegari. On the kinematic formula in the lives of the saints. Notices of the American Mathematical Society, 67(7):1042–1044, August 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2021:PP

[Cal21a] Danny Calegari. Pentagonum pentagonorum. Notices of the American Mathematical Society, 68(8):1367–1369, September 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2021:SSA

[Cal21b] Danny Calegari. Six surfaces (almost) surely. Notices of the American Mathematical Society, 68(2):267–269, February 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2022:CGS

[Cal22a] Danny Calegari. Circles great and small. Notices of the American Mathematical Society, 69 (3):465–467, March 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2022:PW

[Cal22b] Danny Calegari. Proofs without words. Notices of the American Mathematical Society, 69(10):1807–1809, November 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2023:ASL

[Cal23a] Danny Calegari. Almost sufficiently large. Notices of the American Mathematical Society, 70(9):1428–1437, October 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2023:PTR

[Cal23b]  Danny Calegari.  A path

through the roots. Notices of the American Mathematical Society, 70(7):1143–1145, August 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2024:BBB

[Cal24a] Danny Calegari. Baguenaudy, baguenaudier, baguenaudiest.

Notices of the American Mathematical Society, 71(1):117–119, January 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2024:PE

[Cal24b] Danny Calegari. A pointless exercise. Notices of the American Mathematical Society, 71 (7):955–957, August 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Campbell:2023:WMF

[Cam23] Lucy J. Campbell. Wavemean-flow interactions in atmospheric fluid flows. Notices of the American Mathematical Society, 70(2):200–213, February 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cannon:2023:AMY

[Can23] Sarah Cannon. All the math you missed (but need to know for graduate school) [book review of 4273409]. Notices of the American Mathematical Society, 70(6):973–975, June/July 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Caporaso:2020:EGP

[Cap20] Lucia Caporaso. Enumerative geometry of plane curves. Notices of the American Mathematical Society, 67(6):771–779, June/July 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Casals:2021:CST

[Cas21] Roger Casals. Contact and symplectic topology: mastering the art of front cooking. Notices of the American Mathematical Society, 68(5):697–706, May 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cakoni:2021:TE

[CCH21] Fioralba Cakoni, David Colton, and Houssem Haddar. Transmission eigenvalues. Notices of the American Mathematical Society, 68(9):1499–1510, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2022:LRA

[CCL^{+}22] Ke Chen, Shi Chen, Qin Li, Jianfeng Lu, and Stephen J. Wright. Low-rank approximation for multiscale PDEs. Notices of the American Mathematical Society, 69(6):901–913, June/July 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Coulon:2024:IIM

[CDHH^{+}24] Rémi Coulon, Gabriel Dorfsman-Hopkins, Edmund Harriss, Martin Skrodzki, Katherine E. Stange, and Glen Whitney. On the importance of illustration for mathematical research. Notices of the American Mathematical Society, 71(1):105–115, January 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cao:2023:JPD

[CDX23] Junyan Cao, Ya Deng, and Jian Xiao. Jean-Pierre Demailly (1957–2022). Notices of the American Mathematical Society, 70(5):782–795, May 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ceniceros:2022:TCA

[Cen22] Jose Ceniceros. Topology: a categorical approach [book review of 4232168]. Notices of the American Mathematical Society, 69(8):1383–1385, September 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chanillo:2022:RWM

[CFKS22] Sagun Chanillo, Bruno Franchi, Carlos E. Kenig, and Eric T. Sawyer. Richard L. Wheeden — in memoriam 1940–2020. Notices of the American Mathematical Society, 69(2):222–231, February 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Caines:2021:MRH

[CGH^{+}21] Peter E. Caines, Phillip Griffiths, Gabrielle Hermann, P. Robert Kotiuga, Arthur J. Krener, Emma Previato, Alan Weinstein, and Peter Woit. Memorial for Robert Hermann. Notices of the American Mathematical Society, 68(9):1530–1539, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cheskidov:2021:SFC

[CGHP^{+}21] Alexey Cheskidov, Nathan Glatt-Holtz, Natasa Pavlovic, Roman Shvydkoy, and Vlad Vicol. Susan Friedlander’s contributions in mathematical fluid dynamics. Notices of the American Mathematical Society, 68 (3):331–343, March 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chan:2021:MSC

[Cha21] Melody Chan. Moduli spaces of curves: classical and tropical. Notices of the American Mathematical Society, 68(10):1700– 1713, November 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2022:IRP

[Che22] Shaoshi Chen. Interview with Richard P. Stanley: the wonderful world of combinatorics. Notices of the American Mathematical Society, 69(7):1231– 1235, August 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2023:PDE

[Che23] Gui-Qiang G. Chen. Partial differential equations of mixed type — analysis and applications. Notices of the American Mathematical Society, 70(1):8– 23, January 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2024:LEC

[Che24] Xiuxiong Chen. The legacy of Eugenio Calabi (1923–2023). Notices of the American Mathematical Society, 71(11):1502– 1519, 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cox:2021:RSZ

[CHJ21] David A. Cox, Michael Harris, and Lizhen Ji. Remembering Steve Zucker. Notices of the American Mathematical Society, 68(7):1156–1172, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calver:2024:MMD

[CHO^{+}24] Wesley Calver, Valentina Harizanov, Eugenio G. Omodeo, Alberto Policriti, and Alexandra Shlapentokh. In memory of Martin Davis. Notices of the American Mathematical Society, 71(7):898–907, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ciocanel:2022:APS

[Cio22] Maria-Veronica Ciocanel. Applications of PDEs and stochastic modeling to protein transport in cell biology. Notices of the American Mathematical Society, 69(11):1899–1907, December 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cohl:2022:LDA

[CIW22] Howard S. Cohl, Mourad E. H. Ismail, and Hung-Hsi Wu. The legacy of Dick Askey (1933– 2019). Notices of the American Mathematical Society, 69(1):59– 75, January 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Colmez:2020:JPW

[CK20] Pierre Colmez and Chandrashekhar Khare. Jean-Pierre Wintenberger. Notices of the American Mathematical Society, 67(1):59–63, January 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Connes:2025:PCV

[CK25] Alain Connes and Joseph Kouneiher. Pierre Cartier: a visionary mathematician. Notices of the American Mathematical Society, 72(5):514–522, May 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Callahan:2023:WH

[CKL^{+}23] Patrick Callahan, Ilya Kapovich, Marc Lackenby, Peter Shalen, and Robin Wilson. Wolfgang Haken, 1928–2022. Notices of the American Mathematical Society, 70(9):1452–1467, October 2023. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Constable:2021:ISM

[CKR21] George Constable, Yvonne Krumbeck, and Tim Rogers. An invitation to stochastic mathematical biology. Notices of the American Mathematical Society, 68(11):1882–1889, December 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chaperon:2020:PMW

[CL20] Marc Chaperon and François Laudenbach. Publication of the mathematical works of René Thom in the collection Documents Mathématiques of the French Mathematical Society. Notices of the American Mathematical Society, 67(7):1034– 1041, August 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chang-Lara:2023:PRT

[CL23] Héctor A. Chang-Lara. A perspective on the regularity theory of degenerate elliptic equations. Notices of the American Mathematical Society, 70(10):1604–1614, November 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Clarkson:2020:QFM

[Cla20] Corrin Clarkson. Queer families and mathematical careers. Notices of the American Mathematical Society, 67(6):862–865, June/July 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Clay:2022:GGT

[Cla22] Matt Clay. Geometric group theory. Notices of the American Mathematical Society, 69(10): 1689–1699, November 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Clader:2024:CCM

[Cla24] Emily Clader. Curve-counting and mirror symmetry. Notices of the American Mathematical Society, 71(9):1140– 1149, September 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Clisby:2021:ECL

[Cli21] Nathan Clisby. Enumerative combinatorics of lattice polymers. Notices of the American Mathematical Society, 68 (4):504–515, April 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Clontz:2022:GYM

[Clo22] Steven Clontz. Games for your mind [book review of 4298720]. Notices of the American Mathematical Society, 69(5):821–825, May 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Clontz:2025:HJT

[Clo25] Steven Clontz. But hey, it’s just a theorem — mathematics communicators on YouTube and TikTok. Notices of the American Mathematical Society, 72 (3):300–303, March 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chodosh:2021:MIM

[CM21] Otis Chodosh and Davi Maximo. The Morse index of a minimal surface. Notices of the American Mathematical Society, 68(6):892–898, June/ July 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Curto:2023:GRR

[CM23] Carina Curto and Katherine Morrison. Graph rules for recurrent neural network dynamics. Notices of the American Mathematical Society, 70 (4):536–551, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chiodo:2024:FGE

[CM24] Maurice Chiodo and Dennis Müller. A field guide to ethics in mathematics. Notices of the American Mathematical Society, 71(7):939–947, August 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Collins:2022:WC

[CMN22] Benoît Collins, Sho Matsumoto, and Jonathan Novak. The Weingarten calculus. Notices of the American Mathematical Society, 69(5):734–745, May 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cameron:2023:RML

[CN23] Naiomi Cameron and Asamoah Nkwanta. Riordan matrices and lattice path enumeration. Notices of the American Mathematical Society, 70(2):231–243, February 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Crovisier:2022:ZPE

[CP22] Sylvain Crovisier and Enrique Pujals. From zero to positive entropy. Notices of the American Mathematical Society, 69 (5):748–761, May 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Crowell:2021:IBH

[Cro21a] Rachel Crowell. An interview with Boris Hasselblatt. Notices of the American Mathematical Society, 68(3):426–429, March 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Crowell:2021:IDU

[Cro21b] Rachel Crowell. An interview with Douglas Ulmer. Notices of the American Mathematical Society, 68(3):430–433, March 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Campos:2021:CLM

[CRS^{+}21] Daniel Campos, Manuel Rivera, María Amelia Salazar, José Alejandro Samper, José Simental, and Soledad Villar. Cibercoloquio Latinoamericano de Matemáticas. Notices of the American Mathematical Society, 68(5):793–797, May 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Csicsery:2024:MJB

[Csi24] George Csicsery. Making journeys of black mathematicians. Notices of the American Mathematical Society, 71(2):240–244, February 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chang:2021:IMD

[CSV21] Stanley Chang, Andy Schultz, and Ismar Volić. Institute for Mathematics and Democracy. Notices of the American Mathematical Society, 68(11):1990– 1993, December 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2021:DMF

[CT21] Gui-Qiang G. Chen and Monica Torres. Divergence-measure fields: Gauss–Green formulas and normal traces. Notices of the American Mathematical Society, 68(8):1282–1290, September 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chen:2025:TUC

[CWA25] Nan Chen, Stephen Wiggins, and Marios Andreou. Taming uncertainty in a complex world: the rise of uncertainty quantification — a tutorial for beginners. Notices of the American Mathematical Society, 72 (3):250–260, March 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Czubak:2024:SVO

[Czu24] Magdalena Czubak. In search of the viscosity operator on Riemannian manifolds. Notices of the American Mathematical Society, 71(1):8–16, January 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DAgostino:2023:SCV

[D’A23] Susan D’Agostino. Some colleges with vulnerable students cull math programs. Notices of the American Mathematical Society, 70(10):1711–1715, November 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Donagi:2023:EPM

[DAG^{+}23] Ron Donagi, Malcolm R. Adams, James Glazebrook, Lisa Jeffrey, Shigeki Matsutani, David Mumford, Volodya Roubtsov, and Ben Thompson. Emma Previato [Memorial tribute]. Notices of the American Mathematical Society, 70(8):1243–1251, September 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Danielli:2020:OOP

[Dan20] Donatella Danielli. An overview of the obstacle problem. Notices of the American Mathematical Society, 67(10):1487– 1497, November 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Davis:2025:PGB

[Dav25] Robert Davis. Polytopes and graphs [book review of 4748520]. Notices of the American Mathematical Society, 72(6):637–641, June/July 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2020:TBO

[DDV^{+}20] Della Dumbaugh, Panagiota Daskalopoulos, Anatoly Vershik, Lev Kapitanski, Nicolai Reshetikhin, Darya Apushkinskaya, and Alexander Nazarov. The ties that bind: Olga Alexandrovna Ladyzhenskaya and the 2022 ICM in St. Petersburg. Notices of the American Mathematical Society, 67 (3):373–381, March 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DeSilva:2023:BTE

[De 23] Daniela De Silva. On Bernoullitype elliptic free boundary problems. Notices of the American Mathematical Society, 70(11):1782–1791, December 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DeRosa:2024:TAM

[De 24] Antonio De Rosa. On the theory of anisotropic minimal surfaces. Notices of the American Mathematical Society, 71 (7):853–859, August 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DeLong:2022:TMT

[DeL22] Matt DeLong. Teaching mathematics through games [book review, MAA Press, Providence, RI, 2021]. Notices of the American Mathematical Society, 69(10):1783–1785, November 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Despeaux:2021:SQJ

[Des21] Sloan Evans Despeaux. From servant to queen: a journey through Victorian mathematics [book review of 3929740]. Notices of the American Mathematical Society, 68(6):987–989, June/July 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devadoss:2022:CMC

[Dev22] Satyan L. Devadoss. Curating mathematics for the 21st century. Notices of the American Mathematical Society, 69 (6):1004–1007, June/July 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Du:2022:BJP

[DFP^{+}22] Jie Du, Eric M. Friedlander, Karen H. Parshall, Leonard L. Scott, and Jianpan Wang. Brian J. Parshall (October 28, 1945–January 17, 2022). Notices of the American Mathematical Society, 69(8):1353– 1357, September 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2020:MSC

[DH20] Edward Dunne and Klaus Hulek. Mathematics subject classification 2020. Notices of the American Mathematical Society, 67(3):410–411, March 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2022:CKP

[DHD^{+}22] Della Dumbaugh, Patti Hunter, Sloan Despeaux, Deborah Kent, and Laura Martini. Celebrating Karen Parshall as an advisor. Notices of the American Mathematical Society, 69 (3):400–403, March 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dzamonja:2022:MKK

[DHMV22] Mirna Džamonja, Joan Hart, Andrea Medini, and Andrés Villaveces. In memory of Ken Kunen. Notices of the American Mathematical Society, 69(10):1758–1769, November 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DiPrisco:2021:MFR

[Di 21] Carlos A. Di Prisco. MAD families and the Ramsey property. Notices of the American Mathematical Society, 68(7):1081– 1086, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dickenstein:2020:AGT

[Dic20] Alicia Dickenstein. Algebraic geometry tools in systems biology. Notices of the American Mathematical Society, 67(11): 1706–1715, December 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dipierro:2020:CBN

[Dip20] Serena Dipierro. A comparison between the nonlocal and the classical worlds: minimal surfaces, phase transitions, and geometric flows. Notices of the American Mathematical Society, 67(9):1324–1335, October 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumitriu:2022:SNR

[DKR22] Ioana Dumitriu, Todd Kemp, and Kavita Ramanan. Stochastic networks and reflecting Brownian motion: the mathematics of Ruth Williams. Notices of the American Mathematical Society, 69(3):363–374, March 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

deLora:2021:DEL

[dL21] Sarah González de Lora. Dr.

Eduardo Luna Ramia. Notices of the American Mathematical Society, 68(8):1322– 1323, September 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Translated by Patrick Scott.

deLima:2023:MAR

[dLLOW23] Pedro Nascimento de Lima, Jonathan Lamb, Osonde Osoba, and Jonathan Welburn. Modeling America’s racial wealth disparities: mathematical models help chart pathways for closing racial wealth gaps. Notices of the American Mathematical Society, 70(7):1136–1141, August 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Derezinski:2021:DPP

[DM21] Michał Dereziński and Michael W.

Mahoney. Determinantal point processes in randomized numerical linear algebra. Notices of the American Mathematical Society, 68(1):34–45, January 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dougherty:2022:LWV

[DMW22] Steven T. Dougherty, Gretchen L. Matthews, and Jay A. Wood. The life and work of Vera Stepen Pless. Notices of the American Mathematical Society, 69(11):1942–1955, December 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions from Janet Beissinger, Richard Brualdi, Nick Crews, Shmuel Friedland, Xiang-dong Hou, W. Cary Huffman, Jon-Lark Kim, Ben Pless, Dan Pless, Naomi Pless, Patrick Solé, Sarah Spence Adams, Vladimir D. Tonchev, Judy Walker and Harold Ward.

Dabo-Niang:2024:UDA

[DNF24] Sophie Dabo-Niang and Camille Frévent. Uncovering data across continua: an introduction to functional data analysis. Notices of the American Mathematical Society, 71(7):869–877, August 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dolgachev:2020:KSY

[Dol20] Igor Dolgachev. Kummer surfaces: 200 years of study. Notices of the American Mathematical Society, 67(10):1527– 1533, November 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2020:TSC

[DR20] Della Dumbaugh and Adrian Rice. A template for success: celebrating the work of Judith Grabiner. Notices of the American Mathematical Society, 67 (3):336–344, March 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Duff:2023:PSH

[DR23] Timothy Duff and Margaret Regan. Polynomial systems, homotopy continuation, and applications. Notices of the American Mathematical Society, 70 (1):151–155, January 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Daskalopoulos:2020:ASG

[DS20a] Panagiota Daskalopoulos and Natasa Sesum. Ancient solutions to geometric flows. Notices of the American Mathematical Society, 67(4):467–474, April 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dundas:2020:IAL

[DS20b] Bjørn Ian Dundas and Christian Skau. Interview with Abel laureate Karen Uhlenbeck. Notices of the American Mathematical Society, 67(3):393–403, March 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Reprinted from the Newsletter of the European Mathematical Society, September 2019.

Dundas:2021:IALa

[DS21a] Bjørn Ian Dundas and Christian Skau. Interview with Abel Laureate 2020 Gregory Margulis. Notices of the American Mathematical Society, 68(6):992–997, June/July 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dundas:2021:IALb

[DS21b] Bjørn Ian Dundas and Christian Skau. Interview with Abel Laureate 2020 Hillel Furstenberg [reprint of 4226847]. Notices of the American Mathematical Society, 68(7):1189–1196, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dundas:2022:AIL

[DS22] Bjørn Ian Dundas and Christian F. Skau. Abel Interview 2021: László Lovász and Avi Wigderson. Notices of the American Mathematical Society, 69(5):828–843, May 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dundas:2023:IAL

[DS23] Bjørn Ian Dundas and Christian F. Skau. Interview with Abel Laureate 2022 Dennis Sullivan. Notices of the American Mathematical Society, 70 (4):632–642, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dundas:2024:AIL

[DS24] Bjørn Ian Dundas and Christian F. Skau. Abel interview 2023: Luis Ángel Caffarelli. Notices of the American Mathematical Society, 71(1):74–78, January 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Reprint of [4655867].

Dundas:2025:AIM

[DS25] Bjørn Ian Dundas and Christian F. Skau. Abel interview 2024: Michel Talagrand. Notices of the American Mathematical Society, 72(5):536–546, May 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2021:PWM

[Dum21] Della Dumbaugh. Prospering within mathematics. Notices of the American Mathematical Society, 68(11):1969–1977, December 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dumbaugh:2022:BGH

[Dum22] Della Dumbaugh. “Brilliance and generosity of heart”: Elizabeth Meckes (1980–2020). Notices of the American Mathematical Society, 69(3):413–416, March 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2020:ON

[Dun20] Edward Dunne. Otto Neugebauer. Notices of the American Mathematical Society, 67(9): 1397–1399, October 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2021:DCI

[Dun21a] Edward Dunne. Don’t count on it. Notices of the American Mathematical Society, 68 (1):114–118, January 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2021:GMM

[Dun21b] Edward Dunne. Geography and MathSciNet(R)/Mathematical Reviews. Notices of the American Mathematical Society, 68 (4):597–603, April 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2021:MTC

[Dun21c] Edward Dunne. Math in the time of coronavirus. Notices of the American Mathematical Society, 68(7):1211–1214, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2021:MDC

[Dun21d] Edward Dunne. MathSciNet for developing countries program. Notices of the American Mathematical Society, 68(11):1985– 1987, December 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dunne:2023:NG

[Dun23] Edward Dunne. The next generation. Notices of the American Mathematical Society, 70 (1):134–139, January 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Draisma:2021:GMG

[DV21] Jan Draisma and Alejandro Vargas. On the gonality of metric graphs. Notices of the American Mathematical Society, 68 (5):687–695, May 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dorff:2022:CLM

[DW22] Michael J. Dorff and Scott A.

Wolpert. A current look at mathematics graduate programs. Notices of the American Mathematical Society, 69 (6):998–1003, June/July 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dyatlov:2023:QET

[Dya23] Semyon Dyatlov. Quantum ergodicity in theorems and pictures. Notices of the American Mathematical Society, 70(10): 1592–1601, November 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

E:2021:DNE

[E21] Weinan E. The dawning of a new era in applied mathematics. Notices of the American Mathematical Society, 68 (4):565–571, April 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Edwards:2024:ITM

[Edw24] Richard A. Edwards. Interdisciplinary teaching of mathematics with primary historical sources. Notices of the American Mathematical Society, 71 (4):467–471, April 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Elkik:2021:LS

[EGMD^{+}21] Renée Elkik, Dorian Goldfeld, Mireille Martin-Deschamps, Christian Peskine, Thomas Tucker, and Shou-Wu Zhang. Lucien Szpiro (1941–2020). Notices of the American Mathematical Society, 68(10):1763– 1777, November 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Erdogan:2021:WLD

[EGT21] M. Burak Erdoğan, William R. Green, and Ebru Toprak. What is … the Dirac equation? Notices of the American Mathematical Society, 68(10):1782– 1785, November 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Escobar:2020:CPA

[EK20] Laura Escobar and Kiumars Kaveh. Convex polytopes, algebraic geometry, and combinatorics. Notices of the American Mathematical Society, 67 (8):1116–1123, September 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eur:2025:CHT

[EL25] Christopher Eur and Matt Larson. Combinatorial Hodge theory. Notices of the American Mathematical Society, 72 (3):261–270, March 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eley:2024:WKS

[Ele24] Peter Eley. What I know for sure. Notices of the American Mathematical Society, 71(2): 200–203, February 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eppstein:2025:PWM

[ELW25] David Eppstein, Joel Brewster Lewis, and Russ Woodroofe. Princ-wiki-a mathematica: Wikipedia editing and mathematics. Notices of the American Mathematical Society, 72(1):65–73, January 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eisenbud:2022:MAW

[EMS22] David Eisenbud, Nicole Mullen, and Cliff Stoll. Mathematics at the airport? Why sure! Notices of the American Mathematical Society, 69(4):612–615, April 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Edmonds:2021:PRM

[EO21] Ranthony A. C. Edmonds and Omayra Ortega. Perseverance and representation: a memorial for Katherine Coleman Goble Johnson (1918–2020). Notices of the American Mathematical Society, 68(3):379–386, March 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eppstein:2025:WDT

[Epp25] David Eppstein. What is … treewidth? Notices of the American Mathematical Society, 72(2):172–175, February 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Esnault:2023:SAP

[Esn23] Hélène Esnault. Some arithmetic properties of complex local systems. Notices of the American Mathematical Society, 70(11):1793–1801, December 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Engquist:2023:MAJ

[ESSV23] Bjorn Engquist, Panagiotis Souganidis, Samuel N. Stechmann, and Vlad Vicol. In memory of Andrew J. Majda. Notices of the American Mathematical Society, 70(10):1648– 1666, November 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eckmann:2024:TDA

[EST24] Jean-Pierre Eckmann, Yaroslav I.

Sobolev, and Tsvi Tlusty. Tumbling downhill along a given curve. Notices of the American Mathematical Society, 71(6): 740–747, June/July 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eisenbud:2022:RDB

[EW22] David Eisenbud and Jerzy Weyman. Remembering David Buchsbaum. Notices of the American Mathematical Society, 69(1):76–87, January 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ellenberg:2025:NB

[EW25] Jordan Ellenberg and Judy L.

Walker. Nigel Boston (1961– 2024). Notices of the American Mathematical Society, 72(6):

614–623, June/July 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Farris:2021:MAD

[Far21] Frank A. Farris. Mathematical art as a discipline. Notices of the American Mathematical Society, 68(6):980–986, June/July 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Farley:2022:WGR

[Far22] Jonathan David Farley. Was Gelfand right? The many loves of lattice theory. Notices of the American Mathematical Society, 69(2):190–197, February 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ferguson:2021:OFI

[Fer21] Samuel J. Ferguson. Obamacare and a fix for the IRS iteration. Notices of the American Mathematical Society, 68(7):1197– 1204, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ferreiros:2025:AAM

[Fer25] José Ferreirós. Axiomatics and abstraction in the mid-20th century [book review of 4605073]. Notices of the American Mathematical Society, 72(2):177–179, February 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Figalli:2025:CMF

[FGKS25] Alessio Figalli, André Guerra, Sunghan Kim, and Henrik Shahgholian. Constraint maps and free boundaries. Notices of the American Mathematical Society, 72(5):494–503, May 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fallat:2020:IEP

[FHLS20] Shaun M. Fallat, Leslie Hogben, Jephian C.-H. Lin, and Bryan L. Shader. The inverse eigenvalue problem of a graph, zero forcing, and related parameters. Notices of the American Mathematical Society, 67(2):257–261, February 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fields:2020:PPL

[Fie20] Evan Fields. Poisson processes and linear programs for allocating shared vehicles. Notices of the American Mathematical Society, 67(11):1804–1805, December 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fisher:2020:RDZ

[Fis20] David Fisher. Recent developments in the Zimmer program. Notices of the American Mathematical Society, 67 (4):492–499, April 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Faranda:2024:CSI

[FLOS24] Davide Faranda, Théo Lacombe, Nina Otter, and Kristian Strommen. Climate science at the interface between topological data analysis and dynamical systems theory. Notices of the American Mathematical Society, 71(2):267–271, February 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Foffano:2023:TGG

[FLU23] Giulia Foffano, Claire Lenz, and Emmanuel Ullmo. Tackling the gender gap from an institutional perspective: the example of IHES. Notices of the American Mathematical Society, 70 (3):470–478, March 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedlander:2020:MAS

[FM20] Eric M. Friedlander and Alexander Merkurjev. In memoriam: Andrei Suslin. Notices of the American Mathematical Society, 67(6):832–841, June/July 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fahroo:2024:SWP

[FMM24] Fariba Fahroo and Reza Malek-Madani. Shock waves and PDEs: a conversation with Barbara Lee Keyfitz. Notices of the American Mathematical Society, 71(3):394–401, March 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fokoue:2020:MSO

[Fok20] Ernest Fokoué. Model selection for optimal prediction in statistical machine learning. Notices of the American Mathematical Society, 67(2):155–168, February 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Foo:2021:CES

[Foo21] Jasmine Foo. Cancer evolution in spatially structured tissues. Notices of the American Mathematical Society, 68(9):1574– 1575, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fouque:2020:PTM

[Fou20] Jean-Pierre Fouque. Probabilistic theory of mean field games, Volumes I & II [book review of 3752669; 3753660]. Notices of the American Mathematical Society, 67(6):855–859, June/ July 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Franceschini:2024:WIB

[Fra24] Alberto Franceschini. What is … a Białynicki–Birula decomposition? Notices of the American Mathematical Society, 71 (4):515–517, April 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Farb:2020:RAM

[FS20] Benson Farb and Peter Shalen. Roger Alperin [Memorial tribute]. Notices of the American Mathematical Society, 67(11): 1768–1769, December 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fuller:2024:UBC

[Ful24]  Edgar Fuller.  Are univer-

sity budget cuts becoming a threat to mathematics? Notices of the American Mathematical Society, 71(5):656–663, May 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gal:2023:APA

[Gal23] Sorin G. Gal. Almost periodic and almost automorphic functions in abstract spaces [book review of 4277844]. Notices of the American Mathematical Society, 70(2):314–316, February 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gamburd:2020:SAB

[Gam20] Alexander Gamburd. Singular adventures of Baron Bourgain in the labyrinth of the continuum. Notices of the American Mathematical Society, 67(11):1716–1733, December 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gamarnik:2025:TSN

[Gam25] David Gamarnik. Turing in the shadows of Nobel and Abel: an algorithmic story behind two recent prizes. Notices of the American Mathematical Society, 72(5):485–493, May 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ganguly:2022:RMG

[Gan22] Shirshendu Ganguly. Random metric geometries on the plane and Kardar–Parisi–Zhang universality. Notices of the American Mathematical Society, 69 (1):26–35, January 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Garcia:2024:WBH

[Gar24a] Stephan Ramon Garcia. What is … the Bateman–Horn conjecture? Notices of the American Mathematical Society, 71 (10):1378–1380, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Garrity:2024:MBR

[Gar24b] Thomas Garrity. The mathematician [book review, Friesen-Press, Altona, MB, 2023]. Notices of the American Mathematical Society, 71(8):1054– 1055, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Golden:2020:MSI

[GBC^{+}20] Kenneth M. Golden, Luke G. Bennetts, Elena Cherkaev, Ian Eisenman, Daniel Feltham, Christopher Horvat, Elizabeth Hunke, Christopher Jones, Donald K. Perovich, Pedro Ponte-Castañeda, Courtenay Strong, Deborah Sulsky, and Andrew J. Wells. Modeling sea ice. Notices of the American Mathematical Society, 67(10):1535–1555, November 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gorban:2023:SST

[GGT23] Alexander Gorban, Bogdan Grechuk, and Ivan Tyukin. Stochastic separation theorems: how geometry may help to correct AI errors. Notices of the American Mathematical Society, 70(1):25–33, January 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Giansiracusa:2022:MM

[Gia22] Noah Giansiracusa. The mathematics of misinformation. Notices of the American Mathematical Society, 69(10):1707– 1715, November 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Giorgi:2023:MSB

[Gio23] Elena Giorgi. The mathematics of stable black holes. Notices of the American Mathematical Society, 70(4):552–563, April 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gallardo:2022:AAC

[GK22] Patricio Gallardo and Matt Kerr. Algebraic and analytic compactifications of moduli spaces. Notices of the American Mathematical Society, 69(9):1476–1485, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Grimmett:2020:HKP

[GL20] Geoffrey R. Grimmett and Gregory F. Lawler. Harry Kesten (1931–2019): a personal and scientific tribute. Notices of the American Mathematical Society, 67(6):822–831, June/July 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gonzalez:2022:FGC

[GL22] José Luis González and Antonio Laface. Finite generation of Cox rings. Notices of the American Mathematical Society, 69(8):1320–1333, September 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Giunti:2023:DCD

[GLR23] Barbara Giunti, Jānis Lazovskis, and Bastian Rieck. DONUT: creation, development, and opportunities of a database. Notices of the American Mathematical Society, 70(10):1640–1644, November 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gu:2020:CCG

[GLY20] Xianfeng Gu, Feng Luo, and Shing-Tung Yau. Computational conformal geometry behind modern technologies. Notices of the American Mathematical Society, 67(10):1509– 1525, November 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Golbeck:2021:IFD

[GM21] Amanda L. Golbeck and Craig A. Molgaard. The inimitable F. N. David: a renaissance statistician. Notices of the American Mathematical Society, 68(3):321–330, March 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Geller:2023:ADS

[GM23a] William Geller and Michał Misiurewicz. Ability and diversity of skills. Notices of the American Mathematical Society, 70(1):60– 66, January 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Golbeck:2023:MWG

[GM23b] Amanda L. Golbeck and Madhu Mazumdar. Mary Wheat Gray: “I never give up”. Notices of the American Mathematical Society, 70(3):390–399, March 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gurvich:2022:SP

[GN22] Vladimir Gurvich and Mariya Naumova. Supercentenarian paradox. Notices of the American Mathematical Society, 69 (9):1472–1475, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gnang:2021:HAG

[Gna21] Edinah K. Gnang. A hypermatrix analog of the general linear group. Notices of the American Mathematical Society, 68(10):1693–1699, November 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goins:2021:UMP

[Goi21] Edray Herber Goins. Updated “MAD pages” website unveiled October 9, 2020. Notices of the American Mathematical Society, 68(2):254–255, February 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goodstein:2020:OTT

[Goo20] Judith R. Goodstein. Olga Taussky-Todd. Notices of the American Mathematical Society, 67(3):345–353, March 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

                                        www.ams.org/notices/202003/
                                        rnoti-p345.pdf.

Granville:2022:MWL

[GP22] Andrew Granville and Carl Pomerance. The man who loved problems: Richard K. Guy. Notices of the American Mathematical Society, 69(4):574–585, April 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goldstine:2022:AMM

[GPS22] Susan Goldstine, Elizabeth Paley, and Henry Segerman. Ars mathemalchemica: from math to art and back again. Notices of the American Mathematical Society, 69(7):1220–1229, August 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Green:2020:WLM

[Gre20] Richard M. Green. What is … a minuscule representation? Notices of the American Mathematical Society, 67(6):850–853, June/July 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Greene:2021:HFH

[Gre21] Joshua Evan Greene. Heegaard Floer homology. Notices of the American Mathematical Society, 68(1):19–33, January 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Greenbaum:2023:OPA

[Gre23] Anne Greenbaum. Open problems in the analysis of Krylov subspace methods. Notices of the American Mathematical Society, 70(5):720–726, May 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gryak:2022:EMS

[Gry22]  Jonathan Gryak.  Explor-

ing mathematical scholarship through the mathematics collaboration graph. Notices of the American Mathematical Society, 69(5):849–855, May 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gorsky:2024:CW

[GS24] Mikhail Gorsky and José Simental. Clusters and weaves. Notices of the American Mathematical Society, 71(8):1004– 1013, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gupta:2023:SAA

[Gup23] Neena Gupta. Some applications of G_{a}-actions on affine varieties. Notices of the American Mathematical Society, 70(9): 1407–1415, October 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gursky:2020:SYA

[GW20] Matthew J. Gursky and Yi Wang.

Sun-Yung Alice Chang and geometric analysis. Notices of the American Mathematical Society, 67(3):318–326, March 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gu:2022:GWW

[GW22] Chong Gu and Yuedong Wang. Grace Wahba and the Wisconsin spline school. Notices of the American Mathematical Society, 69(3):340–350, March 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gwynne:2020:RSL

[Gwy20] Ewain Gwynne. Random surfaces and Liouville quantum gravity. Notices of the American Mathematical Society, 67 (4):484–491, April 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hamblin:2021:MUC

[Ham21] Elizabeth Platt Hamblin. MSRI-UP: using community and connectedness to increase diversity in mathematics. Notices of the American Mathematical Society, 68(5):819–823, May 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hamkins:2025:NPP

[Ham25] Joel David Hamkins. The nearly perfect prediction theorem. Notices of the American Mathematical Society, 72 (3):308–309, March 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harriss:2021:CBB

[Har21] Edmund Harriss. Counting books and beyond: some mathematical books for children. Notices of the American Mathematical Society, 68(5):788–791, May 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harington:2023:APO

[Har23] Robert M. Harington. AMS primer on open access. Notices of the American Mathematical Society, 70(3):492–501, March 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hart:2024:WCW

[Har24] George Hart. What can we say about “math/art”? Notices of the American Mathematical Society, 71(4):520–525, April 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hasenstab:2023:CNN

[Has23] Kyle Hasenstab. Convolutional neural networks and their applications in medical imaging: a primer for mathematicians. Notices of the American Mathematical Society, 70(1):45–59, January 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Haunsperger:2024:AIT

[Hau24] Deanna Haunsperger. Aspiring and inspiring: tenure and leadership in academic mathematics [book review, Amer. Math. Soc., Providence, RI, 2023]. Notices of the American Mathematical Society, 71(2):235–237, February 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hayes:2021:WCB

[Hay21] Brian Hayes. The Weil conjectures [book review of 3931749]. Notices of the American Mathematical Society, 68(2):240–243, February 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Higgins:2022:SDF

[HB22] Raegan Higgins and Heidi Berger. The N_{0} story: discrete fractional calculus. Notices of the American Mathematical Society, 69(2):180–189, February 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hicks:2024:MFZ

[HB24] Illya V. Hicks and Boris Brimkov. The many face(t)s of zero forcing. Notices of the American Mathematical Society, 71(2):167–173, February 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

He:2022:CPL

[He22] Xuhua He. Chevalley Prize in Lie Theory. Notices of the American Mathematical Society, 69(4):641–642, April 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hedden:2020:KTC

[Hed20] Matthew Hedden. Knot theory and complex curves. Notices of the American Mathematical Society, 67(4):538–539, April 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Heilman:2021:DSE

[Hei21] Steven Heilman. Designing stable elections. Notices of the American Mathematical Society, 68(4):516–527, April 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hershberger:2021:ACF

[Her21a] Scott Hershberger. AMS Congressional Fellow AJ Stewart: from baking bread to using mathematics for social justice. Notices of the American Mathematical Society, 68(9):1597– 1598, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hershberger:2021:ITK

[Her21b] Scott Hershberger. An interview with Tyler Kloefkorn. Notices of the American Mathematical Society, 68(10):1814– 1816, November 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Herron:2022:YMD

[Her22a] Isom H. Herron. A year that makes the difference: achievements of Ford Foundation postdocs in the mathematical sciences. Notices of the American Mathematical Society, 69 (8):1358–1367, September 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hershberger:2022:ALN

[Her22b] Scott Hershberger. An arXiv for lecture notes: instructors and authors share how they have benefited from AMS open math notes. Notices of the American Mathematical Society, 69 (4):668–670, April 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hershberger:2022:CCC

[Her22c] Scott Hershberger. Courts, commissions, and consultations: how mathematicians are working to end gerrymandering. Notices of the American Mathematical Society, 69(4):616–623, April 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hershberger:2022:MHM

[Her22d] Scott Hershberger. The mathematicians at the heart of MathSciNet(R): associate editors at Mathematical Reviews share how their work makes possible an invaluable research tool. Notices of the American Mathematical Society, 69 (6):1062–1064, June/July 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hershberger:2022:OTJ

[Her22e] Scott Hershberger. One teen’s journey from local math club to professional publication. Notices of the American Mathematical Society, 69(3):470–472, March 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hershberger:2022:WTF

[Her22f] Scott Hershberger. Working at their full potential: the impact of the AMS Birman Fellowship. Notices of the American Mathematical Society, 69(3):452–457, March 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hershberger:2022:WYM

[Her22g] Scott Hershberger. The world of YouTube math communication. Notices of the American Mathematical Society, 69(10):1789– 1793, November 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hurlbert:2021:GPB

[HK21] Glenn Hurlbert and Franklin Kenter. Graph pebbling: a blend of graph theory, number theory, and optimization. Notices of the American Mathematical Society, 68(11):1900– 1913, December 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hamman:2025:WAG

[HKL25] Kira Hamman, Gizem Karaali, and Lew Ludwig. Who’s afraid of generative AI? Ethical considerations for mathematics instructors. Notices of the American Mathematical Society, 72 (3):292–299, March 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Huang:2022:TST

[HL22] Lan-Hsuan Huang and Dan A. Lee. Trapped surfaces, topology of black holes, and the positive mass theorem. Notices of the American Mathematical Society, 69(4):536–545, April 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Han:2024:IIA

[HL24] Qing Han and Marta Lewicka. Isometric immersions and applications. Notices of the American Mathematical Society, 71 (6):749–755, June/July 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harris:2020:LLH

[HLQ^{+}20] Pamela E. Harris, Alicia Prieto Langarica, Vanessa Rivera Quiñones, Luis Sordo Vieira, Rosaura Uscanga, and Andrés R. Vindas Meléndez. 2020 Lathisms: Latinxs and Hispanics in the mathematical sciences. Notices of the American Mathematical Society, 67(8):1165– 1169, September 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Huerta:2021:WLP

[HM21] César Lozano Huerta and Alex Massarenti. On the weak Lefschetz principle in birational geometry. Notices of the American Mathematical Society, 68 (1):47–55, January 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Horgan:2022:AMP

[HM22] Cornelius O. Horgan and Jeremiah G. Murphy. On an angle with magical properties. Notices of the American Mathematical Society, 69(1):22–25, January 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hauenstein:2025:RNA

[HMR^{+}25] Jonathan D. Hauenstein, Dhagash Mehta, Margaret H. Regan, Samantha N. Sherman, and Charles W. Wampler. Real numerical algebraic geometry.

Notices of the American Mathematical Society, 72(2):217–221, February 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Howell:2022:GAW

[HN22] Karin-Therese Howell and Nancy Ann Neudauer. Grace Alele-Williams: Nigerian mathematician of many firsts — breaking down barriers and opening paths. Notices of the American Mathematical Society, 69(3):439–445, March 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Holtz:2024:CVZ

[Hol24] Olga Holtz. Classical values of zeta, as simple as possible but not simpler. Notices of the American Mathematical Society, 71(6):725–731, June/ July 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Houston:2023:NAM

[Hou23a] Johnny L. Houston. National Association of Mathematicians, Inc. (NAM) passing the torch: a reflection of NAM’s development and growth by NAM’s leaders/contributors — the first five decades. Notices of the American Mathematical Society, 70(2):283–292, February 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Houston:2023:RPA

[Hou23b] Johnny L. Houston. Remembering Professor Aderemi O. Kuku (1941–2022), an internationally recognized mathematician and scholar. Notices of the American Mathematical Society, 70(2):278–282, February 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Houston:2024:LEB

[Hou24]  Johnny L. Houston.  The

legacy of Evelyn Boyd Granville (1924–2023). Notices of the American Mathematical Society, 71(2):204–211, February 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hollings:2025:TAE

[HP25] Christopher D. Hollings and R. B. Parkinson. Triangulating ancient Egyptian mathematics. Notices of the American Mathematical Society, 72 (4):406–421, April 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harris:2021:LLH

[HPLQ^{+}21a] Pamela E. Harris, Alicia Prieto-Langarica, Vanessa Rivera Quiñones, Luis Sordo Vieira, Rosaura Uscanga, and Andrés R. Vindas Meléndez. 2021 Lathisms: Latinxs and Hispanics in the mathematical sciences. Notices of the American Mathematical Society, 68(8):1351– 1355, September 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harris:2021:ITC

[HPLQ^{+}21b] Pamela E. Harris, Alicia Prieto-Langarica, Vanessa Rivera Quiñones, Luis Sordo Vieira, Rosaura Uscanga, and Andrés R. Vindas Meléndez. Introducing Testimonios: a collection of stories of Latinx and Hispanic mathematicians. Notices of the American Mathematical Society, 68(8):1336–1344, September 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Halverson:2022:GWG

[HR22] Tom Halverson and Arun Ram. Gems from the work of Georgia Benkart. Notices of the American Mathematical Society, 69 (3):375–384, March 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hasto:2024:SFU

[HS24] Peter Hästö and Jon R. Star. Strategy flexibility in university mathematics. Notices of the American Mathematical Society, 71(7):917–923, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Huebschmann:2023:CM

[Hue23] Johannes Huebschmann. Crossed modules. Notices of the American Mathematical Society, 70(11):1802–1813, December 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hughes:2024:WLR

[Hug24] Rhonda Hughes. A word from … Rhonda Hughes: the status of women in mathematics. Notices of the American Mathematical Society, 71(3):300–301, March 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hernandez:2023:GW

[HW23] Eugenio Hernández and Edward N. Wilson. Guido L. Weiss (1928–2021). Notices of the American Mathematical Society, 70(1):95–103, January 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions by Ronald Coifman, Mauro Maggioni, Yves Meyer, Fulvio Ricci, Hrvoje Šikić, Fernando Soria, Anita Tabacco and Rodolfo H. Torres.

Hanlon:2025:PHR

[HW25a] Philip J. Hanlon and Mark C.

Wilson. Philip Hanlon reflects on his career in university administration. Notices of the American Mathematical Society, 72(4):386–391, April 2025. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Harington:2025:RHT

[HW25b] Robert Harington and Mark C. Wilson. Robert Harington on trends in AMS publishing. Notices of the American Mathematical Society, 72(4):426–429, April 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hunt:2021:JDB

[HWW21] Fern Y. Hunt, Robin T. Wilson, and Daniel A. Williams. James A. Donaldson — Black mathematician, advocate, gentleman (1941–2019). Notices of the American Mathematical Society, 68(2):216–225, February 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hynd:2021:HJE

[Hyn21] Ryan Hynd. The Hamilton– Jacobi equation, then and now. Notices of the American Mathematical Society, 68(9):1457– 1467, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Inniss:2022:DRJ

[IJS22] Tasha Inniss, Raymond L. Johnson, and Sherry Scott. Dr. Raymond L. Johnson: a mathematical journey and some reflections on African Americans in graduate mathematical sciences programs in the US. Notices of the American Mathematical Society, 69(2):232–238, February 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Inniss:2022:TFI

[ILM^{+}22] Tasha Inniss, Jim Lewis, Irina Mitrea, Kasso Okoudjou, Adriana Salerno, Francis Su, and Dylan Thurston. Towards a fully inclusive mathematics profession — one year later. Notices of the American Mathematical Society, 69(7):1214– 1219, August 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Iyer:2023:MEJ

[IMNR23] Uma N. Iyer, Susan Montgomery, Siu-Hung Ng, and David Radford. In memory of Earl Jay Taft (1931–2021). Notices of the American Mathematical Society, 70(6):939–943, June/July 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Inniss:2024:DSM

[Inn24] Tasha R. Inniss. Dr. Shirley M. McBay and Dr. Dionne L. Price: leaving a legacy of leadership and lifelong inspiration. Notices of the American Mathematical Society, 71(11):1521– 1525, 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jefferson:2023:MHP

[Jef23a] Brett Jefferson. Mathematizing human perception. Notices of the American Mathematical Society, 70(7):1081–1088, August 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jeffries:2023:DPN

[Jef23b] Jack Jeffries. Differentiating by prime numbers. Notices of the American Mathematical Society, 70(11):1772–1779, December 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jensen:2021:CFA

[Jen21] David Jensen. Chip firing and algebraic curves. Notices of the American Mathematical Society, 68(11):1875–1881, December 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jett:2022:REM

[Jet22] Christopher C. Jett. Racial equity in mathematics: reflections and recommendations from a black mathematics educator. Notices of the American Mathematical Society, 69(9):1566– 1569, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jenda:2021:MPU

[JL21] Overtoun Jenda and Edward Lungu. Masamu program, a US–Africa partnership. Notices of the American Mathematical Society, 68(2):262–265, February 2021. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Jiang:2021:CMD

[JLL^{+}21] Yu Hang Jiang, Tong Liu, Zhiya Lou, Jeffrey S. Rosenthal, Shanshan Shangguan, Fei Wang, and Zixuan Wu. The coupling/minorization/drift approach to Markov chain convergence rates. Notices of the American Mathematical Society, 68(4):471–479, April 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jones:2023:DSS

[JMM^{+}23] Quindel Jones, Andrés R. Vindas Meléndez, Ariana Mendible, Manuchehr Aminian, Heather Zinn Brooks, Nathan Alexander, Carrie Diaz Eaton, and Philip Chodrow. Data science and social justice in the mathematics community. Notices of the American Mathematical Society, 70(9):1479–1489, October 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Johnson:2024:PPM

[JMW24] Joseph Johnson, Theo McKenzie, and Tian An Wong. Predictive policing: a mathematical primer. Notices of the American Mathematical Society, 71 (7):929–937, August 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jorgensen:2020:REG

[Jor20] Palle Jorgensen. Remembrances of Edward G. Effros. Notices of the American Mathematical Society, 67(10):1573–1587, November 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Joshi:2020:DPE

[Jos20] Nalini Joshi. Discrete Painlevé equations. Notices of the American Mathematical Society, 67 (6):797–805, June/July 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jorgenson:2023:LHW

[JSS23] Jay Jorgenson, Thomas Shemanske, and Lejla Smajlović. Lynne Heather Walling (1958– 2021). Notices of the American Mathematical Society, 70 (3):435–449, March 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jungic:2024:TNS

[Jun24] Veselin Jungić. Teaching: no strangers at this party. Notices of the American Mathematical Society, 71(10):1331– 1334, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kass:2020:JSE

[Kas20] Jesse Leo Kass. James L. Solomon and the end of segregation at the University of South Carolina. Notices of the American Mathematical Society, 67(2):192–200, February 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kasman:2021:MLM

[Kas21] Alex Kasman. Machines like me [book review, Nan A. Talese, New York, NY, 2019]; and A universe of sufficient size [book review, Picador Australia, Sydney, 2019]. Notices of the American Mathematical Society, 68(10):1786–1789, November 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kashina:2023:SMJ

[Kas23a] Yevgenia Kashina. Susan Montgomery: a journey in noncommutative algebra. Notices of the American Mathematical Society, 70(3):368–379, March 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kass:2023:JCC

[Kas23b] Jesse Leo Kass. Joseph Carter Corbin: Arkansas’s “profound mathematician”. Notices of the American Mathematical Society, 70(2):293–300, February 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kawamuro:2023:TBB

[Kaw23] Keiko Kawamuro. The tiling book [book review of 4567741].

Notices of the American Mathematical Society, 70(9):1470– 1475, October 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kehoe:2022:IAA

[Keh22]  Elaine Kehoe.  2022 IMU

Awards announced. Notices of the American Mathematical Society, 69(10):1823–1827, November 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kenig:2020:IMU

[Ken20a] Carlos E. Kenig. The International Mathematical Union (IMU) at 100. Notices of the American Mathematical Society, 67(3):404–407, March 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kent:2020:AMC

[Ken20b] Deborah Kent. American mathematics 1890–1913: catching up to Europe [book review of MR3675486]. Notices of the American Mathematical Society, 67(5):701–705, May 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kessel:2022:TWB

[Kes22] Cathy Kessel. Tenured women at Berkeley before 1980. Notices of the American Mathematical Society, 69(3):427–438, March 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kahrobaei:2023:GBC

[KFN23] Delaram Kahrobaei, Ramón Flores, and Marialaura Noce. Group-based cryptography in the quantum era. Notices of the American Mathematical Society, 70(5):752–763, May 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Khare:2025:NWM

[Kha25] Chandrashekhar B. Khare.

Notes on writing a mathematical memoir. Notices of the American Mathematical Society, 72(6):643–645, June/July 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kirkpatrick:2021:DMM

[Kir21] Kay L. Kirkpatrick. Dark matter may be a Bose–Einstein condensate of axions. Notices of the American Mathematical Society, 68(9):1576–1577, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Khanin:2021:MF

[KLSS21] Konstantin Khanin, Mikhail Lyubich, Eric D. Siggia, and Yakov Sinai. Mitchell Feigenbaum. Notices of the American Mathematical Society, 68 (5):757–767, May 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Klawe:2025:MKRa

[KLW25a] Maria Klawe, Haydee Lindo, and Mark C. Wilson. Maria Klawe reflects on her career so far, part I. Notices of the American Mathematical Society, 72 (1):26–31, January 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See also [KLW25b].

Klawe:2025:MKRb

[KLW25b] Maria Klawe, Haydee Lindo, and Mark C. Wilson. Maria Klawe reflects on her career so far, Part II. Notices of the American Mathematical Society, 72(2):154–159, February 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See also [KLW25a].

Kim:2023:PP

[KM23] Woojin Kim and Facundo Mémoli. Persistence over posets. Notices of the American Mathematical Society, 70 (8):1214–1225, September 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Khadjavi:2021:NUP

[KMMM21] Lily Khadjavi, Reza Malek-Madani, and Tanya Moore. Navigating an uncharted path: the life and legacy of Dr. Gladys B. West. Notices of the American Mathematical Society, 68 (3):357–364, March 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Khadjavi:2020:SLH

[KMWW20] Lily Khadjavi, Tanya Moore, Kimberly Weems, and Ulrica Wilson. Shining a light on a hidden figure: Dorothy Hoover. Notices of the American Mathematical Society, 67(3):368–372, March 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kikianty:2022:NRR

[KN22] Eder Kikianty and Loyiso Nongxa. The national research rating system in South Africa: the past impacts the future. Notices of the American Mathematical Society, 69(2):239–247, February 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Komarova:2020:MCE

[Kom20] Natalia L. Komarova. Mathematics of cellular evolution and some biomedical applications. Notices of the American Mathematical Society, 67(5):708–709, May 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kirby:2024:MCC

[KR24] Rob Kirby and Marc A. Rieffel.

Memories of Calvin C. Moore. Notices of the American Mathematical Society, 71(9):1186– 1195, September 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kuznetsov:2020:LJM

[Kuz20] Nikolay Kuznetsov. The legacy of Józef Marcinkiewicz: four hallmarks of genius. In memoriam of an extraordinary analyst. Notices of the American Mathematical Society, 67 (5):690–700, May 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lamb:2020:GGC

[Lam20] Evelyn Lamb. GROWing a graduate cohort. Notices of the American Mathematical Society, 67(5):724–729, May 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lamb:2021:IJP

[Lam21a] Evelyn Lamb. An interview with Jill Pipher. Notices of the American Mathematical Society, 68(4):644–647, April 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lamb:2021:IRC

[Lam21b] Evelyn Lamb. An interview with Ruth Charney. Notices of the American Mathematical Society, 68(2):276–279, February 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lam:2025:SRM

[Lam25] Thomas Lam. Selected results from the mathematical conventions survey. Notices of the American Mathematical Society, 72(4):423–425, April 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lange:2022:CCC

[Lan22] Karen Lange. Calibrating computational complexity via definability: the work of Julia F. Knight. Notices of the American Mathematical Society, 69 (3):351–362, March 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

LaRose:2024:TML

[LaR24] P. Gavin LaRose. Teaching and managing large undergraduate mathematics courses. Notices of the American Mathematical Society, 71(10):1334–1337, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lauret:2020:FS

[Lau20a] Jorge Lauret. Finding solitons.

Notices of the American Mathematical Society, 67(5):647–657, May 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lauter:2020:HKY

[Lau20b] Kristin Lauter. How to keep your secrets in a post-quantum world. Notices of the American Mathematical Society, 67(1):22– 29, January 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lavine:2020:APV

[Lav20] Michael Lavine. Assessing procedures vs. assessing evidence and the principles of sufficiency, conditionality, and likelihood. Notices of the American Mathematical Society, 67(7):988–993, August 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lawrence:2020:MSJ

[Law20] Emille Davie Lawrence. Mathematics for social justice [book review of MR3967051]. Notices of the American Mathematical Society, 67(4):532–533, April 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Li:2024:LBP

[LDAST24] Yunyi Li, Maria De-Arteaga, and Maytal Saar-Tsechansky. Label bias: a pervasive and invisibilized problem. Notices of the American Mathematical Society, 71(8):1069–1077, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

LeBris:2020:MMD

[Le 20] Claude Le Bris. Mathematics for the modeling of defects in materials. Notices of the American Mathematical Society, 67(6):788–796, June/ July 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Levine:2020:SRT

[Lev20] Marc Levine. Some recent trends in motivic homotopy theory. Notices of the American Mathematical Society, 67(1):9– 20, January 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Licata:2022:CGM

[Lic22] Joan E. Licata. Contact geometry and the mapping class group. Notices of the American Mathematical Society, 69 (4):515–521, April 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lewis:2022:UCD

[LIJM22] Taryn Butler Lewis, Tasha R.

Inniss, Monica Jackson, and Calandra Tate Moore. An unlikely cohort: disrupting the stereotypes about mathematicians and their careers. Notices of the American Mathematical Society, 69(3):417–426, March 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Liu:2022:LRT

[Liu22] Chun-Hung Liu. Legacy of Robin Thomas. Notices of the American Mathematical Society, 69(6):966–977, June/July 2022. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Liu:2021:MMV

[LLM21] Ju Liu, Ingrid S. Lan, and Alison L. Marsden. Mathematical modeling of the vascular system. Notices of the American Mathematical Society, 68 (5):707–720, May 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lim:2023:WLE

[LN23] Lek-Heng Lim and Bradley J. Nelson. What is … an equivariant neural network? Notices of the American Mathematical Society, 70(4):619–625, April 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lobb:2024:FKH

[Lob24] Andrew Lobb. A feeling for Khovanov homology. Notices of the American Mathematical Society, 71(5):621–631, May 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Long:2023:TNB

[Lon23] D. D. Long. Topology of numbers [book review of 4567712]. Notices of the American Mathematical Society, 70(7):1130– 1133, August 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lopez:2021:MTP

[Lop21]  Hector Lopez.  Mathemat-

ics in the theory and practice of auctions. Notices of the American Mathematical Society, 68(10):1793–1795, November 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lorenat:2020:RPM

[Lor20] Jemma Lorenat. A richer picture of mathematics: the Göttingen tradition and beyond [book review of MR3753659]. Notices of the American Mathematical Society, 67(7):1029– 1031, August 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lamb:2023:RMT

[LOW23] Evelyn Lamb, Omayra Ortega, and Robin Wilson. The role of mathematics in today’s movement for racial justice. Notices of the American Mathematical Society, 70(2):319–324, February 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lewis-Pye:2020:CPC

[LP20] Andrew Lewis-Pye. Cryptocurrencies: protocols for consensus. Notices of the American Mathematical Society, 67(9):1307– 1315, October 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lewis:2024:MCP

[LP24] Albert C. Lewis and Karen Hunger Parshall. Mathematicians confront political tests: the American Mathematical Society and the Red Scare in 1954. Notices of the American Mathematical Society, 71(4):500–513, April 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lauda:2022:IC

[LS22] Aaron D. Lauda and Joshua Sussan. An invitation to categorification. Notices of the American Mathematical Society, 69(1):11–21, January 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lucas:2020:MDS

[LST20] Stephen K. Lucas, Evelyn Sander, and Laura Taalman. Modeling dynamical systems for 3D printing. Notices of the American Mathematical Society, 67(11):1692–1705, December 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lupercio:2021:NCG

[Lup21]  Ernesto Lupercio.  Non-

commutative geometry indomitable. Notices of the American Mathematical Society, 68 (1):7–18, January 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Larson:2023:MAD

[LV23] Eric Larson and Isabel Vogt.

Making accessible documents using LaTEX. Notices of the American Mathematical Society, 70(1):68–71, January 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
journals/notices/202301/rnoti-
p68.pdf.

Makowsky:2020:MFM

[Mak20] Johann A. Makowsky. Modernism, fiction and mathematics [book review, Edinburgh Univ. Press, Edinburgh, 2018]. Notices of the American Mathematical Society, 67(10):1589– 1595, November 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mann:2021:SHD

[Man21] Kathryn Mann. The structure of homeomorphism and diffeomorphism groups. Notices of the American Mathematical Society, 68(4):482–492, April 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Marland:2020:MSB

[Mar20] Eric Marland. Mathematics for sustainability [book review of 3791821]. Notices of the American Mathematical Society, 67(11):1784–1787, December 2020. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Margalit:2023:MSA

[Mar23] Dan Margalit. Mixing surfaces, algebra, and geometry. Notices of the American Mathematical Society, 70(5):737–749, May 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Marciniak:2024:GTA

[Mar24] Małgorzata Marciniak. Graph theory in America: the first hundred years [book review of 4574842]. Notices of the American Mathematical Society, 71 (1):68–71, January 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Massey:2022:CCB

[Mas22] William A. Massey. From CAARMS25 to CAARMS 2021 and beyond: conferences for African–American researchers in the mathematical sciences. Notices of the American Mathematical Society, 69(2):261–268, February 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mavinga:2023:SSE

[Mav23] Nsoki Mamie Mavinga. Steklov spectrum and elliptic problems with nonlinear boundary conditions. Notices of the American Mathematical Society, 70(2): 214–222, February 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mawi:2023:FOO

[Maw23] Henok Mawi. Freeform optics:

optimal transport, Minkowski method, and Monge–Ampèretype equations. Notices of the American Mathematical Society, 70(2):223–230, February 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

McFaddin:2022:VGR

[McF22] Patrick K. McFaddin. A.T.V.’s for (geometric) off-roading: a gentle introduction to arithmetic toric varieties. Notices of the American Mathematical Society, 69(7):1113–1125, August 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Melnick:2021:RTG

[Mel21] Karin Melnick. Rigidity of transformation groups in differential geometry. Notices of the American Mathematical Society, 68(5):721–732, May 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mendiratta:2020:RAT

[Men20] Veena B. Mendiratta. Reliability analysis in telecommunications. Notices of the American Mathematical Society, 67(6): 867–875, June/July 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2020:CMT

[Mer20a] Sophia D. Merow. “Creative mathematician” transforms student errors into “Sartorial celebration”. Notices of the American Mathematical Society, 67 (6):883–885, June/July 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2020:GUF

[Mer20b] Sophia D. Merow. Gift from Uhlenbeck funds Karen EDGE Fellowship. Notices of the American Mathematical Society, 67(2):228–230, February 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2020:ME

[Mer20c] Sophia D. Merow. Mathematicians eponymous. Notices of the American Mathematical Society, 67(8):1173–1175, September 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2020:SMM

[Mer20d] Sophia D. Merow. A standup makes a math joke … and it lands. Notices of the American Mathematical Society, 67(11):1801–1803, December 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2020:YFF

[Mer20e] Sophia D. Merow. Your favorite field a superpower? Notices of the American Mathematical Society, 67(3):415–417, March 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2021:PSL

[Mer21] Sophia D. Merow. Po-Shen Loh makes good. Notices of the American Mathematical Society, 68(9):1565–1568, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2022:ARS

[Mer22a] Sophia D. Merow. Artist residency sparks fruitful collaboration. Notices of the American Mathematical Society, 69(7): 1236–1237, August 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2022:MNL

[Mer22b] Sophia D. Merow. Mathematician (no longer) pseudonymous. Notices of the American Mathematical Society, 69(8):1388– 1391, September 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2022:MTM

[Mer22c] Sophia D. Merow. Mathematician turns melodist. Notices of the American Mathematical Society, 69(11):1984–1985, December 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Merow:2022:TMT

[Mer22d]  Sophia D. Merow.  Tricky

math, but trippy graphics: the quixotic search for the “3D Mandelbrot”. Notices of the American Mathematical Society, 69(4):624–627, April 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mezic:2021:KOG

[Mez21] Igor Mezić. Koopman operator, geometry, and learning of dynamical systems. Notices of the American Mathematical Society, 68(7):1087–1105, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Michalek:2023:EGM

[Mic23] Mateusz Michałek. Enumerative geometry meets statistics, combinatorics, and topology. Notices of the American Mathematical Society, 70 (4):588–597, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Miller:2022:MQC

[Mil22] Carl A. Miller. The mathematics of quantum coin-flipping. Notices of the American Mathematical Society, 69(11):1908– 1917, December 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Milner:2023:ST

[Mil23] Fabio Augusto Milner. SAC-NAS turning 50. Notices of the American Mathematical Society, 70(10):1704–1709, November 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moore:2024:UOD

[ML24] Tanya Moore and Josh Levy.

The unfailing optimism of Dr. Gloria Ford Gilmer (1928– 2021). Notices of the American Mathematical Society, 71 (3):366–373, March 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mazur:2024:JCC

[MLW^{+}24] Barry Mazur, Stephen Lichtenbaum, Andrew Wiles, Leila Schneps, Sujatha, and Mahesh Kakde. John Coates (1945– 2022): celebration of a life. Notices of the American Mathematical Society, 71(1):46–56, January 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Matheus:2021:DAL

[MM21] Carlos Matheus and Carlos Gustavo Moreira. Diophantine approximation, Lagrange and Markov spectra, and dynamical Cantor sets. Notices of the American Mathematical Society, 68(8):1301–1311, September 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mitrea:2020:SDT

[MMM20] Dorina Mitrea, Irina Mitrea, and Marius Mitrea. A sharp divergence theorem with nontangential traces. Notices of the American Mathematical Society, 67(9):1295–1305, October 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Melczer:2020:CAA

[MMP20] Stephen Melczer, Marni Mishna, and Robin Pemantle. Combinatorial adventures in analysis, algebra, and topology. Notices of the American Mathematical Society, 67(2):262–265, February 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Malek-Madani:2025:FFM

[MMS25] Reza Malek-Madani and Karen Saxe. Federal funding for mathematics research — what’s changed in the last six years? Notices of the American Mathematical Society, 72(6):646–649, June/July 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

McKeow:2024:UYP

[MN24] Stephen E. McKeow and Cheikh Birahim Ndiaye. Uniformization and the Yamabe problem. Notices of the American Mathematical Society, 71 (7):878–883, August 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moraga:2024:RMM

[Mor24a] Joaquín Moraga. Running a minimal model program. Notices of the American Mathematical Society, 71(1):17–27, January 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mori:2024:WIP

[Mor24b] J. Carlos Martínez Mori. What is… a parking function? Notices of the American Mathematical Society, 71(8):1062– 1065, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Melhuish:2025:MBS

[MPDL25] Kathleen Melhuish, Cody L.

Patterson, Paul C. Dawkins, and Kristen Lew. Moving beyond show and tell in proofbased courses. Notices of the American Mathematical Society, 72(1):51–59, January 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mumford:2021:RCT

[Mum21] David Mumford. Ruminations on cosmology and time. Notices of the American Mathematical Society, 68(10):1715– 1725, November 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Murray:2021:AWM

[Mur21] Margaret A. M. Murray. American women mathematics PhDs of the 1940s. Notices of the American Mathematical Society, 68(3):399–405, March 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Murray:2023:AWM

[Mur23] Margaret A. M. Murray. American women mathematics PhDs of the 1950s. Notices of the American Mathematical Society, 70(3):450–457, March 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Musso:2022:BBC

[Mus22]  Monica Musso.  Bubbling

blow-up in critical elliptic and parabolic problems. Notices of the American Mathematical Society, 69(10):1700–1706, November 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mutze:2024:HCG

[Müt24] Torsten Mütze. On Hamilton cycles in graphs defined by intersecting set systems. Notices of the American Mathematical Society, 71(5):583–592, May 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moynihan:2023:WMG

[MWW^{+}23] Benjamin Moynihan, Robin T.

Wilson, Joan Wynne, Lee J. McEwan, Mary M. West, Frank E. Davis, Herbert Clemens, Greg Budzban, Edith Aurora Graf, and Aidan Soguero. What’s math got to do with it?: Bob Moses, algebra, and the movement for civil rights, January 23, 1935–July 25, 2021. Notices of the American Mathematical Society, 70(2):258–277, February 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nagloo:2021:MTD

[Nag21] Joel Nagloo. Model theory and differential equations. Notices of the American Mathematical Society, 68(2):177–185, February 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Naibo:2021:BPO

[Nai21] Virginia Naibo. Bilinear pseudodifferential operators and the Hörmander classes. Notices of the American Mathematical Society, 68(7):1119–1130, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nariman:2024:FDG

[Nar24] Sam Nariman. Foliations and diffeomorphism groups. Notices of the American Mathematical Society, 71(11):1471– 1483, 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nathanson:2023:TQA

[Nat23] Melvyn B. Nathanson. Three questions about John von Neumann [book review of The man from the future, Norton, New York, NY, 2022]. Notices of the American Mathematical Society, 70(5):813–817, May 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nelson:2020:DRE

[Nel20] Jelani Nelson. Dimensionality reduction in Euclidean space. Notices of the American Mathematical Society, 67(10):1498– 1507, November 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nielsen:2022:MFI

[Nie22] Frank Nielsen. The many faces of information geometry. Notices of the American Mathematical Society, 69(1):36–45, January 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nielsen:2023:WDI

[Nie23] Frank Nielsen. What is … an inductive mean? Notices of the American Mathematical Society, 70(11):1851–1855, December 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Norton:2020:SIP

[Nor20] Douglas Norton. Simplicity:

ideals of practice in mathematics and the arts [book review of MR3642967]. Notices of the American Mathematical Society, 67(2):209–215, February 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Oh:2022:ETH

[Oh22] Hee Oh. Euclidean traveller in hyperbolic worlds. Notices of the American Mathematical Society, 69(11):1888–1897, December 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Olson:2023:CMC

[Ols23a] Emily J. Olson. Count me in: community and belonging in mathematics [book review, MAA Press, Providence, RI, 2022]. Notices of the American Mathematical Society, 70 (3):464–467, March 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Olsson:2023:CB

[Ols23b] Jørn B. Olsson. Christine Bessenrodt, 1958–2022. Notices of the American Mathematical Society, 70(7):1101–1105, August 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Omar:2021:CPS

[Oma21] Mohamed Omar. Conversational problem solving [book review of 4249565]. Notices of the American Mathematical Society, 68(9):1569–1571, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ortega:2020:PMC

[Ort20] Angela Ortega. The Prym map: from coverings to Abelian varieties. Notices of the American Mathematical Society, 67(9): 1316–1323, October 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Okoudjou:2022:RS

[ORT22] Kasso Okoudjou, Luke Rogers, and Alexander Teplyaev. Robert Strichartz. Notices of the American Mathematical Society, 69 (6):979–981, June/July 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

OShea:2025:MAH

[O’S25] Donal B. O’Shea. In memoriam — Alan H. Durfee (1943– 2024). Notices of the American Mathematical Society, 72 (4):404–405, April 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Osborne:2020:MSA

[OWD20] Jason A. Osborne, Melody Wen, and Neil Dey. MLBDecideR: a SHINY app for baseball. Notices of the American Mathematical Society, 67(9): 1389–1395, October 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Orellana:2023:SGT

[OZ23] Rosa Orellana and Mike Zabrocki.

The symmetric group through a dual perspective. Notices of the American Mathematical Society, 70(6):897–904, June/ July 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Parker:2024:TPM

[PA24] Ross Parker and Alejandro Aceves. Topological photonics: a mathematical perspective. Notices of the American Mathematical Society, 71(8): 994–1003, August 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pak:2022:UCB

[Pak22] Igor Pak. The unity of combinatorics [book review of 4306671]. Notices of the American Mathematical Society, 69(1):108–111, January 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pak:2025:HDP

[Pak25] Igor Pak. How do physicists publish papers? The case of Physical Review E. Notices of the American Mathematical Society, 72(2):198–201, February 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Parshall:2021:GBG

[Par21] Karen Hunger Parshall. Giovanni Battista Guccia: pioneer of international cooperation in mathematics [book review of 3822345]. Notices of the American Mathematical Society, 68 (3):411–413, March 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Park:2023:TPR

[Par23] Jinyoung Park. Threshold phenomena for random discrete structures. Notices of the American Mathematical Society, 70(10):1615–1625, November 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pasten:2023:DA

[Pas23] Hector Pasten. Definability and arithmetic. Notices of the American Mathematical Society, 70(9):1385–1393, October 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pinzon-Caicedo:2022:AIF

[PCR22] Juanita Pinzón-Caicedo and Daniel Ruberman. Applications of instanton Floer homology. Notices of the American Mathematical Society, 69(8):1307– 1319, September 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Peherstorfer:2022:BKB

[Peh22] Benjamin Peherstorfer. Breaking the Kolmogorov barrier with nonlinear model reduction. Notices of the American Mathematical Society, 69(5):725–733, May 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pereira:2024:COF

[Per24] Jorge Vitório Pereira. Codimension one foliations on projective manifolds. Notices of the American Mathematical Society, 71(8):1025–1031, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pierce:2021:EMS

[Pie21] Lillian B. Pierce. Elias M. Stein (1931–2018) [Memorial tribute]. Notices of the American Mathematical Society, 68(4):546–563, April 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions.

Pinsky:2024:NKL

[Pin24] Tali Pinsky. New knots in the Lorenz equations. Notices of the American Mathematical Society, 71(10):1314–1319, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pinto:2025:WBO

[PL25] Alon Pinto and Yvonne Lai. What is … a boundary object? Notices of the American Mathematical Society, 72(5):523–527, May 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Poisson:2022:VDG

[Poi22] Eric Poisson. Visual differential geometry and forms — a mathematical drama in five acts [book review of 4298719]. Notices of the American Mathematical Society, 69(11):1979–1981, December 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pollack:2021:EGT

[Pol21] Aaron Pollack. Exceptional groups and their modular forms. Notices of the American Mathematical Society, 68(2): 194–203, February 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pommersheim:2025:IQA

[Pom25] Jamie Pommersheim. Introduction to quantum algorithms [book review of 4731955]. Notices of the American Mathematical Society, 72(3):284–288, March 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Potrie:2021:PHD

[Pot21] Rafael Potrie. Partially hyperbolic dynamics and 3-manifold topology. Notices of the American Mathematical Society, 68 (6):885–891, June/July 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pilgrim:2025:GPE

[PQY^{+}25] Mary E. Pilgrim, Kelsey Quaisley, Franklin Yu, Mary Beisiegel, Pablo Duran Oliva, Mariah Moschetti, Rebecca Segal, and ELITE PD Research Group. Guiding principles for equity-oriented professional development for mathematics graduate teaching assistants. Notices of the American Mathematical Society, 72(2):164–171, February 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Previato:2020:VE

[Pre20] Emma Previato. Victor Enolski (1945–2019). Notices of the American Mathematical Society, 67(11):1755–1767, December 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Phong:2022:MK

[PS22] Duong H. Phong and Yum-Tong Siu. Masatake Kuranishi (1924– 2021). Notices of the American Mathematical Society, 69 (5):788–805, May 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pach:2024:EGR

[PSAH24] János Pach, Micha Sharir, Noga Alon, and Andreas Holmsen. Eli Goodman (1933–2021) and Ricky Pollack (1935–2018). Notices of the American Mathematical Society, 71(8):1044– 1053, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Premchandar:2023:NFA

[PST23] Sumegha Premchandar, Bhattacharya Shrijita, and Maiti Tapabrata. Normalizing flows aided variational inference: a useful alternative to MCMC? Notices of the American Mathematical Society, 70(7):1059– 1070, August 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Parker:2023:CSH

[PT23] John R. Parker and Ser Peow Tan. Caroline Series and hyperbolic geometry. Notices of the American Mathematical Society, 70(3):380–389, March 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Possamai:2024:PSS

[PT24] Dylan Possamaï and Ludovic Tangpi. On the population size in stochastic differential games. Notices of the American Mathematical Society, 71(4):454–462, April 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pudwell:2020:PPP

[Pud20] Lara Pudwell. From permutation patterns to the periodic table. Notices of the American Mathematical Society, 67(7): 994–1001, August 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Peters:2020:PI

[PV20] Han Peters and Liz Vivas.

Parabolic implosion. Notices of the American Mathematical Society, 67(8):1095–1103, September 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rajapakse:2021:CDS

[Raj21] Indika Rajapakse. Conversation with Dr. Steve Smale and Dr. Lee Hartwell. Notices of the American Mathematical Society, 68(9):1578–1582, October 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ramos:2022:GMT

[Ram22] Eric Ramos. The graph minor theorem meets algebra. Notices of the American Mathematical Society, 69(8):1297– 1305, September 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ranganathan:2023:TGF

[Ran23] Dhruv Ranganathan. Tropical geometry forwards and backwards. Notices of the American Mathematical Society, 70(7): 1048–1058, August 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rathjen:2023:AMS

[Rat23] Michael Rathjen. The art of measuring the strength of theories. Notices of the American Mathematical Society, 70(7): 1071–1079, August 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Reys:2024:MDO

[Rey24] Robert Reys. Mathematics departments that offer a PhD in mathematics education: what courses in mathematics education are required? Notices of the American Mathematical Society, 71(7):925–927, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Reyes:2025:TNS

[Rey25] Manuel Reyes. A tour of noncommutative spectral theories. Notices of the American Mathematical Society, 72(2):145–153, February 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ribet:2021:APB

[Rib21] Kenneth A. Ribet. The Abel Prize 2013–2017 [book review of 3930576]. Notices of the American Mathematical Society, 68(1):99–101, January 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rice:2024:ENR

[Ric24] Adrian Rice. An enchantress of number? Reassessing the mathematical reputation of Ada Lovelace. Notices of the American Mathematical Society, 71 (3):374–385, March 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Riehl:2023:CCT

[Rie23] Emily Riehl. Could ∞-category theory be taught to undergraduates? Notices of the American Mathematical Society, 70 (5):727–736, May 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ringer:2024:PC

[Rin24] Talia Ringer. Proofs and conversations. Notices of the American Mathematical Society, 71(9):1179–1182, September 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ritore:2023:GMS

[Rit23] Manuel Ritoré. On the geometry of metric spaces. Notices of the American Mathematical Society, 70(9):1417–1425, October 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ross-Nazzal:2023:TIS

[RN23] Jennifer Ross-Nazzal. From the till to the International Space Station: Josephine Jue and spaceflight computing. Notices of the American Mathematical Society, 70(3):458–463, March 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Robinson:2023:MDS

[Rob23] Angela Robinson. The mathematics of digital signatures. Notices of the American Mathematical Society, 70(6):998–999, June/July 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rohatgi:2023:DMC

[Roh23] Ranjan Rohatgi. The discrete mathematical charms of Paul Erdős: a simple introduction [book review of 4282416]. Notices of the American Mathematical Society, 70(11):1856– 1859, December 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rojas:2025:IAF

[Roj25] J. Maurice Rojas. An introduction to algorithmic fewnomial theory over R. Notices of the American Mathematical Society, 72(6):594–603, June/ July 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rosenhouse:2020:BUC

[Ros20a] Jason Rosenhouse. Bicycle or unicycle: a collection of intriguing mathematical puzzles [book review, Math. Assoc. America, Washington, D.C., 2020]. Notices of the American Mathematical Society, 67(9):1382– 1385, October 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rossi:2020:NDE

[Ros20b] Julio D. Rossi. Nonlocal diffusion equations with integrable kernels. Notices of the American Mathematical Society, 67 (8):1125–1133, September 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ross:2021:HMS

[Ros21] William T. Ross. The history of mathematics: a source-based approach (Volumes 1 & 2) [book review, Math. Assoc. America, Providence, RI, 2019, 2021]. Notices of the American Mathematical Society, 68(11):1979– 1981, December 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rosenberg:2022:MM

[Ros22] Noah A. Rosenberg. Mendel of mathematics. Notices of the American Mathematical Society, 69(9):1564–1565, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rowe:2022:MUI

[Row22] David E. Rowe. Meeting under the integral sign? The Oslo Congress of Mathematicians on the eve of the Second World War [book review of 4249622]. Notices of the American Mathematical Society, 69 (4):605–609, April 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rapinchuk:2021:RDT

[RR21] Andrei S. Rapinchuk and Igor A. Rapinchuk. Recent developments in the theory of linear algebraic groups: good reduction and finiteness properties. Notices of the American Mathematical Society, 68(6): 899–910, June/July 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Roberts:2022:HM

[RR22a] David P. Roberts and Fernando Rodriguez Villegas. Hypergeometric motives. Notices of the American Mathematical Society, 69(6):914–929, June/July 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rosenthal:2022:OSR

[RR22b] Daniel Rosenthal and Jeffrey S. Rosenthal. Optimal strategies and rules for the game of horse. Notices of the American Mathematical Society, 69(6):1014– 1017, June/July 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rodriguez:2024:FIB

[RR24] Rubí E. Rodríguez and Anita M.

Rojas. A fruitful interaction between algebra, geometry, and topology: varieties through the lens of group actions. Notices of the American Mathematical Society, 71(6):715–723, June/ July 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Reys:2022:SPP

[RRS22] Robert Reys, Barbara Reys, and Jeffrey Shih. Some patterns of PhDs in mathematics awarded annually by institutions of higher education in the United States over the last two decades. Notices of the American Mathematical Society, 69 (1):96–107, January 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rayan:2021:WLH

[RS21] Steven Rayan and Laura P.

Schaposnik. What is … a hyperpolygon? Notices of the American Mathematical Society, 68(1):94–97, January 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rincon:2023:TC

[RTY23] Felipe Rincón, Ngoc Mai Tran, and Josephine Yu. Tropical combinatorics. Notices of the American Mathematical Society, 70(1):34–43, January 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rubio:2020:CME

[Rub20] Ivelisse M. Rubio. The covering method for exponential sums and some applications. Notices of the American Mathematical Society, 67(5):616–624, May 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ruiz:2023:RPF

[Rui23] Patricia Alonso Ruiz. Random phenomena with fractallike features. Notices of the American Mathematical Society, 70(10):1626–1633, November 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ruthotto:2024:DEC

[Rut24] Lars Ruthotto. Differential equations for continuous-time deep learning. Notices of the American Mathematical Society, 71(5):613–620, May 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rolland:2022:SPM

[RW22] Rita Jiménez Rolland and Jennifer C. H. Wilson. Stability properties of moduli spaces. Notices of the American Mathematical Society, 69(4):522–533, April 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ryba:2022:CMA

[Ryb22a] Alex Ryba. Conway’s mathematics after Conway. Notices of the American Mathematical Society, 69(7):1145–1155, August 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions by Philip Ehrlich, Richard Kenyon, Jeffrey Lagarias, James Propp and Louis H. Kauffman.

Ryba:2022:JHC

[Ryb22b]  Alex Ryba.  John Horton

Conway (1937–2020). Notices of the American Mathematical Society, 69(7):1171–1187, August 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions by R. T. Curtis, Richard Borcherds, Manjul Bhargava, Dierk Schleicher, Jane Gilman, Aaron Siegel and Louis H. Kauffman.

Sam:2020:SRS

[Sam20] Steven V. Sam. Structures in representation stability. Notices of the American Mathematical Society, 67(1):38–43, January 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sangwin:2021:HFS

[San21] Chris Sangwin. Hex: the full story [book review of 3931302]. Notices of the American Mathematical Society, 68(8):1345– 1347, September 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saxe:2020:AAW

[Sax20] Karen Saxe. African American women honored with Congressional Gold Medals. Notices of the American Mathematical Society, 67(3):408–409, March 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saxe:2021:EL

[Sax21] Karen Saxe. Eric Lander. Notices of the American Mathematical Society, 68(7):1205– 1210, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saxe:2022:MCI

[Sax22] Karen Saxe. A mathematician in Congress: interview with Representative Jerry McNerney. Notices of the American Mathematical Society, 69(2): 269–273, February 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schaposnik:2020:HBR

[Sch20] Laura P. Schaposnik. Higgs bundles — recent applications.

Notices of the American Mathematical Society, 67(5):625–634, May 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Scharf:2021:SAA

[Sch21a] Henry Scharf. Statistical analysis of animal movement: understanding behavior through hierarchical parametric models. Notices of the American Mathematical Society, 68(6):911–924, June/July 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schlenk:2021:DMS

[Sch21b] Felix Schlenk. Dusa McDuff and symplectic geometry. Notices of the American Mathematical Society, 68(3):346–356, March 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Scherich:2024:MDB

[Sch24] Nancy Scherich. Math dance:

bringing the process to life. Notices of the American Mathematical Society, 71(10):1385– 1389, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schilling:2025:CTF

[Sch25] Anne Schilling. The combinatorial theory flip. Notices of the American Mathematical Society, 72(1):74–77, January 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Scott:2022:MUA

[Sco22] Patrick Scott. In memoriam: Ubiratan D’Ambrosio (1932– 2021). Notices of the American Mathematical Society, 69 (8):1348–1351, September 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Searle:2023:SSL

[Sea23] Catherine Searle. Symmetries of spaces with lower curvature bounds. Notices of the American Mathematical Society, 70 (4):564–575, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Segerman:2021:RAA

[Seg21] Henry Segerman. Rolling acrobatic apparatus. Notices of the American Mathematical Society, 68(7):1106–1118, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shinder:2025:HGR

[Shi25] Evgeny Shinder. How to give a really bad math talk. Notices of the American Mathematical Society, 72(3):271–273, March 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shlapentokh:2021:HTP

[Shl21] Alexandra Shlapentokh. Hilbert’s tenth problem [book review of 3931317]. Notices of the American Mathematical Society, 68 (4):576–579, April 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Silver:2022:EYB

[Sil22] Daniel S. Silver. Erasable you [book review of 4298723]. Notices of the American Mathematical Society, 69(2):252–255, February 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sinapova:2022:ICC

[Sin22] Dima Sinapova. Infinite cardinals and combinatorial principles. Notices of the American Mathematical Society, 69(9): 1510–1517, October 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Skripka:2020:UNO

[Skr20] Anna Skripka. Untangling noncommutativity with operator integrals. Notices of the American Mathematical Society, 67 (1):45–55, January 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schapira:2024:MMS

[SKT^{+}24] Pierre Schapira, Takahiro Kawai, Yoshitsugu Takei, Masaki Kashiwara, Barry M. McCoy, Craig Tracy, Tetsuji Miwa, and Michio Jimbo. Memories of Mikio Sato (1928–2023). Notices of the American Mathematical Society, 71(5):645–653, May 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schmidt:2024:UGA

[SM24] Paul G. Schmidt and Amnon J. Meir. Using generative AI for literature searches and scholarly writing: is the integrity of the scientific discourse in jeopardy? Notices of the American Mathematical Society, 71(1):93–104, January 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Spolaor:2022:MFC

[Spo22]  Luca Spolaor.  Monotonic-

ity formulas in the calculus of variation. Notices of the American Mathematical Society, 69(10):1731–1737, November 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Springer:2025:CNS

[Spr25] Max Springer. The curious nature of science communication: my summer at Scientific American. Notices of the American Mathematical Society, 72(2):222–223, February 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Semenov:2023:BLT

[SSU23] Evgenii Semenov, Fedor Sukochev, and Alexandr Usachev. Banach limits and their applications. Notices of the American Mathematical Society, 70(6):918–925, June/July 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Salter:2020:SBT

[ST20] Nick Salter and Bena Tshishiku.

Surface bundles in topology, algebraic geometry, and group theory. Notices of the American Mathematical Society, 67(2): 146–154, February 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shen:2022:NKM

[ST22] Alexander Shen and Serge Tabachnikov. Nikolay Konstantinov, 01.02.1932–07.03.2021, a mathematical educator par excellence. Notices of the American Mathematical Society, 69(11):1933–1941, December 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Steele:2023:CSH

[Ste23] John Steele. Combining science and history: Lis Brack-Bernsen’s contributions to the history of Babylonian astronomy. Notices of the American Mathematical Society, 70 (3):400–407, March 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saez-Trumper:2022:PWM

[STHRVG22] Mariel Sáez-Trumper, Jeanne Hersant, Francisca Rioseco, and Andrea Vera-Gajardo. Promoting women in mathematics: experiences from Chile. Notices of the American Mathematical Society, 69(8):1376–1381, September 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stinga:2023:FDF

[Sti23] Pablo Raúl Stinga. Fractional derivatives: Fourier, elephants, memory effects, viscoelastic materials, and anomalous diffusions. Notices of the American Mathematical Society, 70 (4):576–587, April 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sun:2022:GCY

[Sun22] Song Sun. Geometry of Calabi– Yau metrics. Notices of the American Mathematical Society, 69(4):546–556, April 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sutton:2021:SGP

[Sut21] Craig Sutton. Spectral geometry in the presence of symmetry. Notices of the American Mathematical Society, 68(6):869–883, June/July 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sulem:2020:RWC

[SW20] Catherine Sulem and C. Eugene Wayne. Remembrances of Walter Craig. Notices of the American Mathematical Society, 67 (4):520–531, April 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sorger:2025:CSD

[SW25] Christoph Sorger and Mark C.

Wilson. Christoph Sorger discusses the IMU. Notices of the American Mathematical Society, 72(5):547–551, May 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tadmor:2021:MSE

[Tad21] Eitan Tadmor. On the mathematics of swarming: emergent behavior in alignment dynamics. Notices of the American Mathematical Society, 68 (4):493–503, April 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tao:2025:MAP

[Tao25] Terence Tao. Machine-assisted proof. Notices of the American Mathematical Society, 72(1):6– 13, January 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taylor:2024:CBF

[Tay24] Krystal Taylor. Connections between fractal geometry and projections. Notices of the American Mathematical Society, 71 (11):1458–1469, 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Teixeira:2020:RTN

[Tei20] Eduardo V. Teixeira. Regularity theory for nonlinear diffusion processes. Notices of the American Mathematical Society, 67(4):475–483, April 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tenka:2021:WLV

[Ten21] Samuel C. Tenka. What is … the VC-dimension? Notices of the American Mathematical Society, 68(4):572–575, April 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tosh:2025:PPL

[TGG^{+}25] Christopher Tosh, Philip Greengard, Ben Goodrich, Andrew Gelman, Aki Vehtari, and Daniel Hsu. The piranha problem: large effects swimming in a small pond. Notices of the American Mathematical Society, 72(1):15–25, January 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Thompson:2020:PBM

[Tho20] Katherine Thompson. The place of blogs in the modern math world. Notices of the American Mathematical Society, 67(11):1780–1783, December 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Trillos:2023:OST

[THSA23] N. García Trillos, B. Hosseini, and D. Sanz-Alonso. From optimization to sampling through gradient flows. Notices of the American Mathematical Society, 70(6):905–917, June/July 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tiede:2022:AST

[Tie22] Hans-Joerg Tiede. The 2022 AAUP survey of tenure practices. Notices of the American Mathematical Society, 69(10): 1794–1801, November 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Trillos:2023:AGP

[TJ23] Nicolás García Trillos and Matt Jacobs. An analytical and geometric perspective on adversarial robustness. Notices of the American Mathematical Society, 70(8):1193–1204, September 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tattersall:2023:CSA

[TM23] James J. Tattersall and Shawn L.

McMurran. Caroline Seely at the AMS (1913–1935): from executive assistant to associate editor. Notices of the American Mathematical Society, 70(7): 1121–1129, August 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Torres:2020:AOF

[Tor20] Rodolfo H. Torres. Almostorthogonality in Fourier analysis: from discrete characterizations of function spaces, to singular integrals, to Leibniz rules for fractional derivatives. Notices of the American Mathematical Society, 67(8):1105– 1115, September 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Trevino:2023:ATO

[Tre23] Rodrigo Treviño. Aperiodic tilings, order, and randomness. Notices of the American Mathematical Society, 70(8):1179– 1191, September 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Trefethen:2025:RA

[Tre25a] Lloyd N. Trefethen. Rational approximation. Notices of the American Mathematical Society, 72(1):78–81, January 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Trefethen:2025:UGB

[Tre25b] Lloyd N. Trefethen. Unbounded growth of band-limited functions. Notices of the American Mathematical Society, 72(6): 666–669, June/July 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Trillos:2022:MFG

[TSAY22] N. García Trillos, D. Sanz-Alonso, and R. Yang. Mathematical foundations of graphbased Bayesian semi-supervised learning. Notices of the American Mathematical Society, 69(10):1717–1729, November 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tsimerman:2024:AOC

[Tsi24] Jacob Tsimerman. The André– Oort conjecture. Notices of the American Mathematical Society, 71(10):1307–1313, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tuffley:2022:IIU

[Tuf22] Christopher Tuffley. IUM1:

Indigenising University Mathematics Symposium, September 20–21, 2021. Notices of the American Mathematical Society, 69(10):1802–1805, November 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Held online.

Turner:2022:ILL

[Tur22] Laura E. Turner. The impact and legacy of The Ladies’ Diary (1704–1840): a women’s declaration [book review of 4306746]. Notices of the American Mathematical Society, 69(3):458–462, March 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Turner:2024:ESS

[Tur24] Laura E. Turner. Early stages of space computing and the involvement of Evelyn Boyd Granville. Notices of the American Mathematical Society, 71 (11):1526–1533, 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Udeigwe:2023:IMM

[Ude23] Lawrence C. Udeigwe. Interfacing music and mathematics: a case for more engagement. Notices of the American Mathematical Society, 70(2):309–313, February 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Udeigwe:2024:RBS

[Ude24] Lawrence Udeigwe. Rate-based synaptic plasticity rules. Notices of the American Mathematical Society, 71(9):1161– 1171, September 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Uhlenbeck:2024:CCL

[Uhl24] Karen Uhlenbeck. The contributions of Chuu-Lian Terng to geometry. Notices of the American Mathematical Society, 71 (3):327–337, March 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ullman:2021:BBR

[Ull21] Daniel H. Ullman. THE BOOK [book review of 3823190]. Notices of the American Mathematical Society, 68(7):1183– 1185, August 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ullman:2022:PBR

[Ull22] Daniel H. Ullman. “The Putnam” [book review of The William Lowell Putnam Mathematical Competition, 2001– 2016: problems, solutions, and commentary, MAA Press, Providence, RI, 2020]. Notices of the American Mathematical Society, 69(9):1574–1577, October 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Urschel:2025:NSG

[Urs25] John Urschel. Numerical stability in Gaussian elimination. Notices of the American Mathematical Society, 72(6):605–613, June/July 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Varilly-Alvarado:2021:GDD

[VA21] Anthony Várilly-Alvarado. The geometric disposition of Diophantine equations. Notices of the American Mathematical Society, 68(8):1291–1300, September 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Valdinoci:2025:TLR

[Val25] Enrico Valdinoci. Toward a long-range theory of capillarity. Notices of the American Mathematical Society, 72(2):134–143, February 2025. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vasquez:2024:MMV

[Vas24] Paula A. Vasquez. Multiscale modeling of viscoelastic fluids. Notices of the American Mathematical Society, 71(8):1015– 1023, August 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

vandenDries:2022:TMT

[vdD22] Lou van den Dries. Transseries, model theory, and Hardy fields. Notices of the American Mathematical Society, 69(4):557–563, April 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Velleman:2022:LPM

[Vel22] Daniel J. Velleman. Lectures on the philosophy of mathematics [book review of4205950]. Notices of the American Mathematical Society, 69(7):1208– 1211, August 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vendramin:2024:WSB

[Ven24] Leandro Vendramin. What is … a skew brace? Notices of the American Mathematical Society, 71(1):65–67, January 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

vanOorschot:2025:CBR

[vOL25] Paul C. van Oorschot and Susan Landau. A conversation with 2024 Bertrand Russell Prize winner Susan Landau. Notices of the American Mathematical Society, 72(2):182–187, February 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

vanWilligenburg:2022:CLP

[vW22] Stephanie van Willigenburg.

Coloring with a limited paintbox. Notices of the American Mathematical Society, 69(6): 930–935, June/July 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vlasenko:2023:VBU

[VZ23] Masha Vlasenko and Efim Zelmanov. Voices from the bombed universities of Ukraine. Notices of the American Mathematical Society, 70(6):987–991, June/July 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wang:2024:RGC

[Wan24] Yilin Wang. From the random geometry of conformally invariant systems to the Kähler geometry of universal Teichmüller space. Notices of the American Mathematical Society, 71(6): 732–739, June/July 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wegert:2020:SMG

[Weg20] Elias Wegert. Seeing the monodromy group of a Blaschke product. Notices of the American Mathematical Society, 67 (7):965–975, August 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weintraub:2020:ECS

[Wei20] Steven H. Weintraub. An expanded classification system for sessions and talks at future JMMs: a letter to the mathematical community. Notices of the American Mathematical Society, 67(10):1602–1605, November 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weiss:2023:JT

[Wei23a] Richard M. Weiss. Jacques Tits (1930–2021). Notices of the American Mathematical Society, 70(1):77–93, January 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With tributes.

Weissman:2023:WIM

[Wei23b] Martin H. Weissman. What is … a metaplectic group? Notices of the American Mathematical Society, 70(5):806–811, May 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Werner:2024:WLO

[Wer24] Marcus C. Werner. What is … optical geometry? Notices of the American Mathematical Society, 71(11):1534– 1537, 2024. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weston:2023:YAB

[Wes23] Anthony Weston. 3000 years of analysis [book review of 4240993]. Notices of the American Mathematical Society, 70 (1):115–121, January 2023. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wiggins:2022:PPP

[WGG22] Stephen Wiggins and Víctor J.

García-Garrido. Painting the phase portrait of a dynamical system with the computational tool of Lagrangian descriptors. Notices of the American Mathematical Society, 69(6):936–949, June/July 2022. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Whitaker:2021:SDH

[Whi21] Nathaniel Whitaker. From segregation to department head. Notices of the American Mathematical Society, 68(2):226–233, February 2021. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wilmer:2020:ORF

[Wil20] Elizabeth Wilmer. Or/oand: a review of Figuring fibers [book review of 3838355]. Notices of the American Mathematical Society, 67(8):1158–1161, September 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Witherspoon:2020:HC

[Wit20] Sarah Witherspoon. Hochschild cohomology. Notices of the American Mathematical Society, 67(6):780–787, June/July 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wolpert:2021:CGT

[Wol21] Scott A. Wolpert. Counting geodesics, Teichmüller space, and random hyperbolic surfaces. Notices of the American Mathematical Society, 68(11): 1890–1899, December 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wolpert:2024:MMW

[Wol24] Scott A. Wolpert. The math major — what size?: Accessing data on your program. Notices of the American Mathematical Society, 71(10):1373– 1377, October 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Woods:2020:DSM

[Woo20] Zerotti Woods. Dr. Shirley McBay. Notices of the American Mathematical Society, 67 (3):382–383, March 2020. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wang:2021:MFC

[WWL^{+}21] Li Wang, Guannan Wang, Xinyi Li, Shan Yu, Myungjin Kim, Yueying Wang, Zhiling Gu, and Lei Gao. Modeling and forecasting COVID-19. Notices of the American Mathematical Society, 68(4):585–595, April 2021. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Webster:2023:DMS

[WY23] Ben Webster and Philsang Yoo.

3-dimensional mirror symmetry. Notices of the American Mathematical Society, 70(9):1395– 1406, October 2023. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Xuan:2022:PCG

[XBH^{+}22] Yao Xuan, Robert Balkin, Jiequn Han, Ruimeng Hu, and Hector D. Ceniceros. Pandemic control, game theory, and machine learning. Notices of the American Mathematical Society, 69(11):1878–1887, December 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Xu:2022:SRI

[Xu22] Chenyang Xu. K-stability: the recent interaction between algebraic geometry and complex geometry. Notices of the American Mathematical Society, 69(7):1126–1136, August 2022. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Yakubu:2020:DPC

[Yak20] Abdul-Aziz Yakubu. Demographic population cycles in infectious salmon anemia models. Notices of the American Mathematical Society, 67(2):170–180, February 2020. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Yang:2024:TLI

[Yan24] Yang Yang. Tracing the light — introduction to the mathematics of bioluminescence tomography. Notices of the American Mathematical Society, 71(10): 1320–1329, October 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Yong:2022:WF

[Yon22] Kamuela E. Yong. A word from . . .. Notices of the American Mathematical Society, 69 (4):513–514, April 2022. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zabell:2025:JSM

[Zab25] Sandy Zabell. Judith Sally (March 23, 1937–January 28, 2024). Notices of the American Mathematical Society, 72(2): 162–163, February 2025. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zelati:2024:MIF

[ZCIM24] Michele Coti Zelati, Gianluca Crippa, Gautam Iyer, and Anna L. Mazzucato. Mixing in incompressible flows: transport, dissipation, and their interplay. Notices of the American Mathematical Society, 71(5):593–604, May 2024. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zhang:2021:BVM

[Zha21] Yan X. Zhang. Blockchain viewed from mathematics. Notices of the American Mathematical Society, 68(10):1740– 1751, November 2021. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zhao:2024:ELE

[Zha24] Yufei Zhao. Equiangular lines and eigenvalue multiplicities. Notices of the American Mathematical Society, 71(9):1151– 1159, September 2024. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zong:2020:CYP

[Zon20] Chuanming Zong. Can you pave the plane with identical tiles? Notices of the American Mathematical Society, 67(5):635–646, May 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zong:2023:JEE

[Zon23] Chuanming Zong. The journey of Euclid’s Elements to China. Notices of the American Mathematical Society, 70(6):953–961, June/July 2023. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).